### v1

# DialectBench: A NLP Benchmark for Dialects, Varieties, and Closely-Related Languages

Abstract
Language technologies should be judged on their usefulness in real-world use cases. An often overlooked aspect in natural language processing (NLP) research and evaluation is language variation in the form of non-standard dialects or language varieties (hereafter, varieties). Most NLP benchmarks are limited to standard language varieties. To fill this gap, we propose DialectBench, the first-ever large-scale benchmark for NLP on varieties, which aggregates an extensive set of task-varied variety datasets (10 text-level tasks covering 281 varieties). This allows for a comprehensive evaluation of NLP system performance on different language varieties. We provide substantial evidence of performance disparities between standard and non-standard language varieties, and we also identify language clusters with larger performance divergence across tasks. We believe DialectBench provides a comprehensive view of the current state of NLP for language varieties and one step towards advancing it further.  [^1] More information can be found at the following:
Code/data: https://github.com/ffaisal93/DialectBench Website: https://fahimfaisal.info/DialectBench.io
1 Introduction
Benchmarking is important for tracking the progress the field of natural language processing (NLP) has made in various tasks. In the past few years, large-scale multilingual benchmarks like XTREME Hu et al. (2020), XTREME-R Ruder et al. (2021), and XGLUE Liang et al. (2020) have played a pivotal role in evaluating the multilingual capabilities of NLP models. These efforts have sought to make model evaluation more accessible to researchers and representative of a variety of languages Song et al. (2023). However, most benchmarks have focused on the standard varieties of languages, largely neglecting non-standard dialects and language varieties Blasi et al. (2022).
Figure 1: DialectBench Evaluation Suite.
We refer to non-standard dialects and language varieties simply as varieties, and sometimes include low-resource related languages, writing system variants, and other kinds of variation. Varieties contain subtle but notable variations in vocabulary, pronunciation, orthography and grammar, reflecting regional, social, and cultural differences Chambers and Trudgill (1998). The non-standard nature of these language varieties oftentimes contributes to the scarcity of substantial datasets that accurately capture these variations Hedderich et al. (2021). As a result, they have often been absent from widely adopted benchmarks, even from admirable efforts like XTREME-up Ruder et al. (2023), GLUECoS Khanuja et al. (2020) and CreoleVal  Lent et al. (2023), which focus on under-resourced, code-switched, and creole languages, respectively.
Figure 2: DialectBench language clusters with their variety counts per task. "Other" encompasses 18 clusters (full cluster list in Appendix Table 8).
It is currently challenging to accurately test the robustness of multilingual models on a large suite of varieties without establishing an NLP evaluation framework covering multiple language clusters (each of which contains standard languages alongside its closely related varieties).
To this end, we create DialectBench, a large-scale benchmark covering 40 language clusters with 281 varieties, spanning 10 NLP tasks. We observe that the performance disparity between different varieties of the same language cluster becomes more pronounced when we shift from zero-shot evaluation to fine-tuning on variety data, because of uneven data availability across varieties. Certain language clusters exhibit varying performance across downstream tasks within the same category, due to low-resource limitations. Additionally, we improve the dialectal task coverage for natural language inference by constructing a translate-test evaluation dataset. Putting these all together, DialectBench serves as a comprehensive suite that attains a multifaceted purpose: identifying broader limitations in dialectal NLP, while reflecting on potential areas for improvement.
2 DialectBench
Category
Task
Metric
Source Dataset
Structured
DEP parsing
UAS
Universal Dependency Zeman et al. (2021), TwitterAAE Blodgett et al. (2018), Singlish Wang et al. (2017)
Prediction
POS tagging
F1
Universal Dependency Zeman et al. (2021), Singlish Wang et al. (2017), Noisy Dialects Blaschke et al. (2023)
NER
F1
Wikiann Pan et al. (2017); Rahimi et al. (2019), Norwegian NER Johansen (2019)
Classification
DId
F1
MADAR Bouamor et al. (2018), DMT Jauhiainen et al. (2019), Greek Sababa and Stassopoulou (2018), DSL-TL Zampieri et al. (2023), Swiss Germans Scherrer et al. (2019)
SA
F1
TSAC Medhaffar et al. (2017), TUNIZI Fourati et al. (2021), DzSentiA Abdelli et al. (2019), SaudiBank Alqahtani et al. (2022), MAC Garouani and Kharroubi (2022), ASTD Nabil et al. (2015), AJGT Alomari et al. (2017), OCLAR Al Omari et al. (2019)
TC
F1
SIB-200 Adelani et al. (2023)
NLI
F1
XNLI Conneau et al. (2018) translate-test
Question
MRC
F1
Belebele Bandarkar et al. (2023)
Answering
EQA
Span F1
SDQA Faisal et al. (2021)
Generation
MT
BLEU
CODET Alam et al. (2023), TIL-MT Mirzakhalov (2021)
Table 1: The tasks and data sources of DialectBench (Detailed discussion: LABEL:{sec:tasks_benchmark}).
DialectBench is a benchmark created to unify dialectal datasets across different tasks to foster research on language varieties and non-standard dialects. Below we describe the design choices we undertook to achieve this goal. This includes our language variety and task selection procedures, data collation methods, and evaluation principles.
Variety Selection
We first looked through papers published in the ACL Anthology [^2] https://aclanthology.org from the last 10 years to find usable language resources, as well as commonly used online data repositories Littauer (n.d.). We selected languages that have well-established, high-resourced varieties. Varieties may vary by location, ethnicity, or other factors. We also found instances where varieties are classified by writing system or even by genre (e.g., Twitter). When varying by location, varieties may be classified by different datasets at different levels of granularity, sometimes country, region, or city. In some cases, we found resources with two or more varieties within one dataset (e.g., the UD_Portuguese-Bosque dependency treebank Rademaker et al. (2017) includes examples from both European and Brazilian Portuguese variants. To incorporate all these cases under one paradigm, we formulate a cluster-variety mapping procedure.
Cluster-Variety Mapping
We construct several language clusters comprising of both high-resourced varieties and their low-resourced counterparts. We use the Glottolog language database Nordhoff (2012) to define clusters and assign varieties as outlined in Figure 1. This design choice enables us to keep varieties that are closely related in terms of either mutual intelligibility, phylogenetic similarity or geographic proximity within the same cluster. Hence, all cluster varieties always root back to the closest common linguistic ancestor and the whole cluster maps to an established phylogenetic subtree. For example, Fiji Hindi and Hindi, with Hindustani [^3] https://glottolog.org/resource/languoid/id/hind1270 as their closest common ancestor, are placed in the Hindustani cluster.
We primarily use the Glottocode language identification scheme Hammarström and Forkel (2022), ensuring a standardized naming scheme across all varieties. For instance, AAE variety from TwitterAAE Blodgett et al. (2018) dependency parsing dataset is renamed as African American Vernacular English with a corresponding Glottocode afri1276. In cases where Glottocodes are unavailable, like for spoken English from South India, we substitute with the immediate ancestor Glottocode (indi1255) and further categorize the varieties using the following metadata identifiers:
  1. 1.
    Area (a): the region where the variety is spoken or where its dataset was collected.
  2. 2.
    Language register (r): frozen, formal, consultative, casual, and intimate.
  3. 3.
    Language mode (m): written, spoken, and signed language.
  4. 4.
    Orthography (o): In DialectBench this is only specific to Sinitic varieties. This could be either traditional or simplified.
  5. 5.
    Identifier (i): Dataset-specific metadata, could be domain (eg. twitter).
We encapsulate all this information, into a naming convention, and use the template: {Glottocode name}-(a:{},r:{},m:{},o:{},i:{}).  [^4] For example, mandarin chinese (a:mainland, o:simplified refers to Mandarin Chinese (mand1415) spoken in Mainland China and written in simplified characters.
Cluster Representative
Each cluster will often have a high-resourced variety usually with the largest speaker population. We choose this high-resourced variety as the cluster representative. This selection might vary across downstream tasks depending on the data availability. We primarily utilize this representative variety to evaluate the performance gap across cluster varieties, and also rely on it for transfer-learning in resource-scarce settings. Sometimes, the members of a cluster are considered closely related languages, and sometimes dialects; to avoid making this distinction, we refer to all the members of a cluster simply as varieties of the cluster representative.
Task and Dataset Selection
In selecting tasks, we maintain a balanced approach, promoting task diversity while also including tasks that require diverse levels of textual understanding. In the end, our complete list of tasks are as follows:
  1. 1.
    Dependency parsing (DEP parsing)
  2. 2.
    Parts-of-speech tagging (POS tagging)
  3. 3.
    Named entity recognition (NER)
  4. 4.
    Dialect identification (DId)
  5. 5.
    Sentiment analysis (SA)
  6. 6.
    Topic classification (TC)
  7. 7.
    Natural language inference (NLI)
  8. 8.
    Multiple-choice machine reading comprehension (MRC)
  9. 9.
    Extractive question answering (EQA)
  10. 10.
    Machine translation (MT)
In Table 1, we present the task and dataset details. We mostly keep the datasets in their originally published form (except for varieties renaming). For NLI, we use the existing English test set of XNLI Conneau et al. (2018) and construct a multilingual dialect-focused translated evaluation dataset. We refer to this as translate-test NLI.
Evaluation Principles
On the ground level, we evaluate existing NLP systems on text-based tasks using standard evaluation metrics (e.g., UAS for parsing, F1 for classification tasks, BLEU for translation). At a global level, we believe a sustainable NLP system should be user-focused while providing substantial (i) linguistic utility and (ii) demographic utility Song et al. (2023); Blasi et al. (2022). Blasi et al. (2022) defined the utility of a task and language, as the corresponding performance normalized by the best possible performance (usually human-level performance). Demographic utility considers the demand for a language technology within a specific language, where the demand is proportional to the number of speakers of that language. Linguistic utility, on the other hand, asserts that “all languages are created equal” regardless of the number of speakers, and hence all languages in the world should receive identical weights.
Overall, we want to capture the performance gap between language clusters (e.g., Anglic vs. Italian Romance) as well as within language clusters (e.g., Norwegian Bokmål vs. Nynorsk). To attain this, we define the performance gap metrics in §3.3 . We also vary the experimental settings in §3.2, including zero-shot and few-shot cross-lingual transfer, as well as fine-tuning with similar high-resource languages. This is essential given that we lack clean annotated data in many varieties.
3 Experiments
Here, we report the entire process involved in creating and evaluating baselines for all of the tasks and varieties in DialectBench. Additionally, we define a dialectal gap metric to analyze performance disparities within and across clusters.
3.1 Models
We evaluate using two multilingual models: mBERT Devlin et al. (2019) and XLM-R Conneau et al. (2020) for all tasks except MT. For MT, we do zero-shot evaluation with NLLB NLLB Team et al. (2022), using both the 600M and 1.3B variants. In addition, we use Mistral 7B Jiang et al. (2023) to evaluate the current capability of LLMs on multilingual and dialectal understanding tasks. Our main goal is collating dialectal data across different languages and tasks under a single platform, hence we do not optimize for the best model performance. Rather, we focus on understanding and reporting the current state of performance on all DialectBench varieties.
3.2 Training and Evaluation
Training and evaluation procedures are largely determined by the availability of training or evaluation data for each task.
For any cluster $C$, let ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{C}}$ be the highest-resourced variety (which is usually the cluster representative) of $C$. In addition, for any variety $v\in C$, we write ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}}$ for the highest-resourced variety, that is, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{C}}$. For any varieties $t$ and $v$, let ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{S}_{t}(v)}$ be the raw evaluation score of a system fine-tuned on $t$ and tested on $v$ (higher is better).
We use six general approaches for task-specific model training:
  1. 1.
    In-variety fine-tuning (${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{S}_{v}(v)}$): In cases where there is available training data for a variety $v$, we fine-tune the base model on $v$. This primarily applies to tasks such as POS tagging and dependency parsing.
  2. 2.
    In-cluster fine-tuning (${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{S}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}}}(v)}$): In-variety fine-tuning is quite resource-intensive when we have a large number of varieties within a cluster $C$. In such cases, we fine-tune the base model on ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{C}}$. Then we evaluate this model on each variety $v\in C$. This is the most common setting in our experiments, as it allows us to evaluate the dialectal performance gap without increasing the computation cost. For the DId task, we typically use a dataset of sentences annotated with variety labels to fine-tune one dialect-identification model for each language cluster.
  3. 3.
[…]
    Task
    In-variety FT
    In-cluster FT
    Combined FT
    Zero-shot
    No ref-erence
    In-context learning
    DEP
    $\checkmark$
    $\checkmark$
    POS
    $\checkmark$
    $\checkmark$
    NER
    $\checkmark$
    $\checkmark$
    EQA
    $\checkmark$
    $\checkmark$
    $\checkmark$
    MRC
    $\checkmark$
    NLI
    $\checkmark$
    TC
    $\checkmark$
    $\checkmark$
    SA
    $\checkmark$
    $\checkmark$
    DI
    $\checkmark$
    MT
    $\checkmark$
    $\checkmark$
    Table 2: Task-specific training and evaluation procedure.
  4. 4.
    Zero-shot evaluation (${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{S}_{\text{eng}}(v)}$): For certain varieties, obtaining training data even for in-cluster fine-tuning can be a challenge. Fortunately, English training data is always available for the datasets we study, so we use English to fill the gaps when we lack in-variety, in-cluster, or combined training data. At the same time, we aim to assess the feasibility of using this zero-shot cross-lingual transfer in reducing any existing performance gap across varieties. So we perform zero-shot cross-lingual transfer from English to each variety for 6 tasks in total. We only leave out those tasks such as dialect identification that explicitly require in-cluster training data.
  5. 5.
    In-context learning (${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{S}_{\text{icl}}(v)}$): When evaluating large language models, we do not fine-tune them but instead rely on prompting and in-context learning. For this, we provide instructions and 5 examples in English as exemplars, followed by a prompt for predicting the test examples. Employing Mistral 7B Jiang et al. (2023), we assess the present effectiveness of a close-to-state-of-the-art LLM on language varieties. The task-specific example prompts are reported in Appendix I.
Table 2 summarizes the task-specific training procedures that we employ based on data availability. Note that, for MT, we perform zero-shot evaluation specifically in the translation direction, (standard variety to English) tested on (dialectal variety to English). But evaluation is a challenge because human-created reference translations into or out of non-standard varieties are usually very limited. Therefore, we adopt an evaluation protocol from previous work (Alam et al., 2023) that uses pseudo-references. Given $x$, a sentence in a variety and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{x}}$, the translation of $x$ into the standard variety, let $y$ be the output of the MT system on input $x$ and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{y}}$ be the output on input ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{x}}$. Then we measure the quality of $y$ (using, e.g., BLEU) compared against ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{y}}$ as a pseudo-reference.
3.3 Quantifying the Dialectal Gap
To quantify the performance disparity across various resource-specific settings, language clusters and varieties, we introduce a dialect performance gap metric ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{t}(u,v)}$: the relative decrease in performance of a system fine-tuned on variety $t$, tested on variety $v$ compared to a baseline variety $u$:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{t}(u,v)}=\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{S}_{t}(u)}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{S}_{t}(v)}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{S}_{t}(u)}}$
with a special case for in-variety fine-tuning:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{in-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}variety}}}(u,v)}=\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{S}_{u}(u)}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{S}_{v}(v)}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{S}_{u}(u)}}.$
For the baseline score, we use either the score on the standard variety for each cluster ($u={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}}$), or, in the zero-shot setting, the score on the language used for fine-tuning, namely English ($u=t=\text{eng}$). Rather than computing an absolute gap, we opt for a relative gap (i.e., dividing by the baseline score). We also indicate whether the training setting is zero-shot ($t=\text{eng}$) or fine-tuned on in-variety, in-cluster, or assembled data. Putting all these together, we compute the following three variations of dialectal inequality.
  1. 1.
    ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}(\text{eng},v)}$: We calculate this metric to get a comprehensive measurement of global disparity across all varieties in a resource-limited environment (zero-shot transfer from English).
  2. 2.
    ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},v)}$: Using this variation, we keep the setting fixed as zero-shot and calculate the gap between the representative variety and any other variety.
  3. 3.
    ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{t}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},v)}$: The two aforementioned metrics shed light on the extent of the variety performance gap in a resource-limited setting. To gain a more comprehensive perspective, we additionally compute another metric, this time utilizing the availability of resources. The computation approach remains as straightforward as before. We just use fine-tuning on a variety $t$ instead of zero-shot transfer from Standard English: $t=\text{in-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}variety}}$ for in-variety fine-tuning, $t={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}}$ for in-cluster fine-tuning, or $t$ is some set of varieties for combined fine-tuning.
For all three ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}}$ metrics, we compute them at the variety level and then average them at the cluster level:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{t}(u,C)}$ $=\frac{1}{|C|}\sum_{v\in C}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{t}(u,v)}$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{in-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}variety}}}(u,C)}$ $=\frac{1}{|C|}\sum_{v\in C}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{in-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}variety}}}(u,v)}.$
4 Results
num. num. avg.
Category Task cl. var. score Max-score cluster/variety Min-score cluster/variety
Structured prediction DEP parsing 16 40 64.3 southwestern shifted romance/brazilian portuguese 94.4 tupi-guarani subgroup i.a/mbyá guaraní (a:brazil) 9.0
POS tagging 17 51 72.1 norwegian/norwegian bokmål (m:written) 98.7 tupi-guarani subgroup i.a/mbyá guaraní (a:brazil) 1.9
NER 27 85 70.1 eastern romance/romanian 94.2 anglic/jamaican creole english 0.0
Sequence classification NLI 15 38 64.2 anglic/english 83.4 sotho-tswana (s.30)/southern sotho 34.6
TC 15 38 77.7 sinitic/classical-middle-modern sinitic (o:traditional) 89.8 kurdish/central kurdish 19.4
DId 6 49 67.0 sinitic/mandarin chinese (a:taiwan, o:simplified) 98.6 southwestern shifted romance/portuguese (m:written) 17.4
SA 1 9 80.3 arabic/tunisian arabic 94.6 arabic/south levantine arabic 58.9
Qusetion Answering MRC 4 11 40.9 anglic/english 53.4 sotho-tswana (s.30)/southern sotho 29.0
EQA 5 24 74.2 arabic/arabic (a:saudi-arabia) 77.9 swahili/swahili (a:tanzania) 63.5
Generation MT-dialect 12 73 25.2 arabic/gulf arabic (a:riy) 43.1 common turkic/sakha 2.5
MT-region 2 41 33.0 high german/central alemannic (a:ur) 44.1 italian romance/italian (a:sardegna) 13.0
Table 3: Task specific result summary using Maximum Obtainable Score. The varieties with the minimum scores exhibit a noticeable lag in performance across various tasks when compared to the average task performance.
We, first of all, discuss results by highlighting the highest possible score per variety, aka the maximum obtainable evaluation scores regardless of evaluation method or training data. Next, we extend our discussion further by reporting the existing dialectal disparity across clusters and varieties.
4.1 Maximum Obtainable Scores
Here we provide key findings from our evaluation on each task. A task-specific summary is reported in Table 3. Detailed results comprising all tasks, models, language clusters and varieties are reported in Tables 10 to 20 in Appendix E.
Figure 3: Maximum scores (max. UAS) in Dep. Parsing task. Yellow-shaded region: Komi is the only cluster having no varieties seen during mBERT pertaining. Colored bars with diagonal stripes: the cluster representative variety. Low-resourced cluster varieties score lower compared to high-resource Germanic clusters.
Structured prediction
We present visualizations for the task-specific maximum scores. We show the one for Dependency Parsing in Fig. 3, where we observe that low-resource varieties from language clusters such as Tupi-Guarani (indigenous South American cluster), Saami and Komi (low-resource Uralic language clusters) have the lowest performance compared to Standard English and other closely related Germanic clusters. These low-resource varieties are also not included in the pretraining stage of our base language models (eg. mBERT). Furthermore, this trend is evident across all three structured prediction tasks. On the other hand, high-resource Indo-European languages such as Portuguese, French, and Norwegian usually perform better.
Sequence classification
For DId and SA, we generally collate different datasets for each language cluster and therefore, report the comparative classification results together. As a result, the locality level (e.g. city/region/country) also varies across cluster. For example, we report city-level DId results for Arabic and High German but country-level results for Portuguese, Spanish and English. In the case of SA, we have region/country-level results for Arabic varieties. For TC and NLI, we have the same set of clusters and varieties. However, we only report the zero-shot transfer performance from Standard English for NLI using the newly created translate-test NLI dataset.
For TC and NLI, we observe the largest in-cluster disparity in the Kurdish cluster, with Northern Kurdish outperforming all others. The Sotho varieties consistently perform significantly lower compared to other clusters. For all three sequence classification tasks, we generally find the Chinese cluster performing on par with high-resource Latin counterparts.
Question answering
We generally do not see large gaps in performance within varieties in each language cluster. In EQA zero-shot experiments, English and its varieties have the highest performance overall and Korean varieties score the lowest. Combined fine-tuning boosts performance on all language clusters except in English. For MRC the performance threshold peaks at 53.4 for Standard English and the minimum we get is 29 for Southern Sotho. More detailed results are presented in Tables 19 and 15.
Machine translation
The performance gap here varies widely across and within language varieties. Performance is similar within the Swiss-German cluster, with higher performance (see Figure 5) across regions in Northern Switzerland, which is geographically closer to Germany. The performance gap for Norwegian dialects (Figure 9(b)) is surprising as we perform zero-shot transfer from Norwegian Nynorsk (a Western dialect) but obtain better performance on the Eastern dialect. Within Arabic, Riyadh is the highest performer while Sfax performs the worst. For the Bengali cluster, Jessore has the highest performance –this is not surprising since it is one of the dialects from which standard Bengali originated Alam et al. (2023). The Ethiopian variety of Tigrinya exhibits a higher performance than the Eritrean one, even though Tigrinya is more commonly spoken in Eritrea  [^5] https://en.wikipedia.org/wiki/Tigrinya_language. Amongst the clusters within the Basque cluster, Barkoxe and Maule have the lowest score while Azkaine scores the highest.
4.2 Dialectal Gap Across Language Clusters
(a) Dependency parsing
(b) Topic classification
(c) Extractive question answering
Figure 4: Dialectal gap visualization for language clusters utilizing zero-shot cross-lingual transfer from Standard English. In the x-axis, values far from zero have a larger performance gap from English whereas, in the y-axis, values far from zero have a larger within cluster gap. Ideally, we want both of them to be close to zero.
In Fig. 4, we plot the zero-shot dialectal gap for three tasks. In the $x$-axis, we report the aggregated cluster-level gap ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}(\text{eng},v)}$, compared against the fine-tuning variety (standard English) while in $y$-axis we report ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},v)}$, the gap compared against the representative variety of a cluster. In an ideal scenario, we would want both of these gap values to be close to zero. However, this is certainly not the case. The general observed trend is that the low-resource clusters have higher gaps of both ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}(\text{eng},C)}$ and $\mathcal{G}_{\text{eng}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},C)$, whereas high-resource Germanic and Sinitic language clusters consistently exhibit low dialectal gaps. That said, certain specific high-resource varieties, such as Standard German and its dialectal counterparts like Swiss German, showcase significant within-cluster dialectal gaps (Fig. 3(a)).
Task
Gap metric Avg. val cluster (max) cluster (min)
DEP parsing ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}(\text{eng},C)}$ 34.0 arabic, 50.8 sw. shift. romance, 19.4
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},C)}$ 15.7 anglic, 34.2 sw. shift. romance, $-0.9$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{in-variety}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},C)}$ 26.4 arabic, 93.8 italian romance, 0.6
POS tagging ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}(\text{eng},C)}$ 27.4 arabic, 58.5 norwegian, 14.7
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},C)}$ 6.7 anglic, 20.9 ew. armenian, $-2.1$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{in-variety}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},C)}$ 6.2 arabic, 29.1 neva, $-0.5$
NER ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}(\text{eng},C)}$ 31.7 kurdish, 77.1 modern dutch, 12.3
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},C)}$ 22.3 kurdish, 78.2 hindustani, $-23.6$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},C)}$ $-28.6$ kurdish, 91.5 sorbian, -1162.7
TC ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}(\text{eng},C)}$ 22.4 kurdish, 74.2 sinitic, 0.4
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},C)}$ 12.5 kurdish, 60.6 norwegian, $-2.2$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},C)}$ $-1.9$ latvian, 21.0 kurdish, $-61.3$
Table 4: Comparative cluster-level dialectal gap across tasks. In general, the average disparity is larger for zero-shot transfer ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}(\text{eng},C)}$. However, when we move from zeroshot to finetune (i.e. $\mathcal{G}_{\text{eng}}\rightarrow\mathcal{G}_{\bar{v}/\text{in-variety}}$) and compute the distance from a cluster representative ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}}$, we observe increased dialectal disparity $|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},C)}|$.
We primarily report dialectal gaps using zero-shot transfer because the finetuning data available across task and cluster is very disproportionate. Often the in-cluster/variety data is not good enough in terms of data quality and quantity. For example, we have 37 varieties in 13 clusters for dependency parsing but out of these, only 20 varieties have data available for in-variety fine-tuning. This lacking becomes more apparent when we compare the statistics of two types of within-cluster dialectal gaps: zero-shot $\mathcal{G}_{\text{eng}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},C)$ against fine-tuning $\mathcal{G}_{\bar{v}/\text{in-variety}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},C)$ in Table 4. In general, the within-cluster dialectal disparity is smaller for zero-shot transfer (i.e. ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},C)}\leq|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\bar{v}/\text{in-variety}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},C)}|$). Here, the in-cluster/variety fine-tuning results in a higher performance deviation primarily due to the inconsistent variety-specific finetuning data quality.
5 Discussion
High resource vs. low resource varieties
The highest-performing varieties are mostly standard high-resource languages and a few high-resource dialects (Norwegian dialects) whereas, the majority of the lowest-performing language variants are low-resourced varieties. This clear distinction of language varieties points towards the large existence of in-cluster dialectal gaps. Furthermore, this finding correlates with language script differences. We observe that 77.2% top-10 varieties in terms of maximum obtainable score are written with Latin script. Another finding is the performance instability of low-resource varieties across tasks. For instance, Old Guarani performs better in DEP parsing whereas, Mbyá Guarani (Paraguay) surpasses it in POS tagging even though the dataset remains the same (i.e. UD). For more detailed comparisons, in Appendix Table 21, we report the top-10 highest and lowest-scoring varieties across different tasks.
zero-shot few-shot / FT
Task mBERT XLM-R mBERT XLM-R
DEP. Parsing 61.6 61.3 76.2 64.3
POS Tagging 69.5 69.7 89.8 89.1
NER 59.7 57.8 65.8 61.4
NLI 56.9 62.5
TC 72.3 71.4 73.1 68.9
MRC 39.4 40.3
EQA 53.9 51.9 69.2 67.2
SA 78.8 80.1
DId 65.8 59.3
win 4/6 2/6 6/8 2/8
Table 5: Base model comparison. We found mBERT was easier to fine-tune using the default hyperparameter setting thus, resulting in a higher winning rate.
Model hyperparameter tuning
In Table 5, we compare the baseline multilingual models. mBERT was comparatively easier to train than XLM-R using the default learning rates reported in the earlier task experiments. Often for tasks such as MRC, we needed to tune hyperparameters (e.g. learning rate, max. sequence length) in case of XLM-R. However, once we identify a hyperparameter configuration that converges in a zero-shot setting, we use the same to train all the language-variant training for that specific task. As a result, for some low-resource languages, XLM-R does not converge compared to mBERT. For example, XLM-R dependency parsing UAS score for Norwegian-NynorskLIA is 56.08 (zeroshot) and 8.25 (in-variety FT) whereas, we get 78.39 for in-variety mBERT fine-tuning. We suspect this hyperparameter tuning issue is one of the contributing factors toward a lower winning rate of XLM-R in few-shot / fine-tuning settings. However, this could be improved further with an extensive parameter grid search and settings specifically tailored for each language cluster.
Positive zero-shot transfer for Latin varieties
Low-resource varieties written in Latin script receive greater benefit in zero-shot because of effective transfer from high-resource Standard English. With the presence of in-cluster/variety finetuning data, we effectively diminish this script effect to some extent. For example in the Hindi cluster, Fiji performs better than its Latin non-standard counterparts with in-cluster finetuning for NER (Table 12). In summary, if the standard variety of a cluster is non-latin but high-resource, then the success rate of in-cluster/variety fine-tuning tends to be higher. However, where all dialects are low-resource, Latin script varieties utilizing zero-shot transfer, eventually surpass others in the performance hierarchy. As an example, we report the zero-shot NER instances where the low-resource varieties perform better than the representative ones in Appendix Table 22 (most of these use Latin script).
LLM evaluation via In-context learning
For SA and EQA tasks, we have in-context learning results using the Mistral7B LLM. Comparing the performance against zero-shot and fine-tuning using our encoder-based models, we find dialect performance of LLM is better than zero-shot transfer but falls behind the finetuned results. On top of that, data contamination Ahuja et al. (2023) during LLM evaluation is another existing issue while considering these few available dialectal resources. Creating translation-based comparable data might be a solution to perform a fair benchmarking of LLMs on low-resource varieties.
Misinterpreting evaluation metrics
We also report the cluster-level population-weighted average (i.e. demographic utility) which rewards a system more when it provides increased linguistic utility (eg. raw F1 score) for varieties, spoken by a larger population compared to varieties spoken by a smaller population. Alone, this metric could be misleading if we consider the fact that the performance gap among all varieties from a particular cluster should be minimal. On the other hand, solely looking into the linguistic utility average does not give a clear picture either (e.g. often overshadows the larger performance deviation of certain varieties having extreme scores). So for all clusters and tasks, we report the linguistic utility average as well as the demographic utility average, the minimum score of a cluster, and the standard deviation in Tables 23 to 30.
6 Conclusion
We propose DialectBench, the first-ever inclusive Dialectal NLP benchmark reporting performance evaluation and disparity across standard and non-standard varieties. This is one step towards the effort of bringing more and more language under the paradigm of NLP technology. We would like to further improve the benchmark, constructing high-quality comparable data and expanding the task coverage to speech-based NLP technologies.
Limitations
The data quality and quantity, variety coverage vary significantly across tasks because of data scarcity issues. We avoid full-scale llm-evaluation consciously because of the uncertain data-contamination issue and their well-known lower performance threshold compared to smaller masked-language-modeling-based fine-tuned models. In addition, we focus on text-based NLP tasks for this current iteration. Moreover, we do not claim the representative varieties of each language clusters to be any kind of superior or standardized forms over the other varieties. These varieties are chosen to perform a well-informed comparison among the perceived well-resourced linguistic variety and its counterparts having lesser data availability. At the same time, the mutual intelligibility and phylogenetic similarity of the similar cluster varieties also vary across cluster and this was not selected in a numerically quantifiable manner.
Evaluation Limitations
To further improve the evaluation fairness of the current version of DialectBench, we need (i) Parallel corpus utilization to prepare task-specific data (ii) Translation-based task data generation to perform comparable analysis (iii) Quantifying the resource-supply and demand as well as population-coverage Song et al. (2023) to identify where a variety stands in the global landscape of linguistic utility. Here, we have accumulated data for diverse varieties across tasks that vary significantly in terms of quality, example count, and domain. However, to perform a perfectly fair comparison of dialectal inequality, we should consider high-quality comparable data (e.g. parallel corpus, similar varieties across tasks) which is not available at this point.
Continuity of DialectBench
Despite our best effort, this benchmark does not include every one of the already published task-specific dialectal datasets. So, our next steps on this project involve hosting the benchmark on the website that displays the current statistics of the datasets in DialectBench. We will also encourage researchers to add new and existing datasets for tasks, language clusters that might be currently missing alongside the respective baselines.
Space Limitations
Our study encompasses a large set of evaluation result tables, their corresponding visualizations and findings analysis. Due to space limitations, we have to move the detailed reports (Appendix E) and the rest of the visualizations (Appendix D) in the Appendix. To better assist, we include an Appendix Table of Content (Table 6) at the introductory section of Appendix (Appendix).
Ethics Statement
This work is a compilation of existent dialectal datasets across different tasks, including structured prediction and generative tasks. Our experiments do not particularly optimize for the best model performance of these tasks. Therefore we acknowledge that for some of the tasks, the baseline models might not be robust enough to handle dialectal text hence resulting in wrong predictions and generations. We believe that this underscores the need for building models robust to different language variations and future work should focus on this.
Appendix
In this supplementary material, we provide the following: (i) Relevant literature review (ii) Overall results and dataset details that we could not fit into the main body of the paper.
Section Desctiption
Appendix A Related Work
Appendix B Tasks of DialectBench
    Translate-Test NLI Dataset Statistics (LABEL:{tab:nli_data})
Appendix C Varieties and Clusters of DialectBench
    DialectBench Variety list (Table 8)
    Language Clusters and Representative Varieties (Table 9)
Appendix D Result Visualizations
Regional maps with aggregated Machine Translation scores (Figs. 5 to 6)
    Map of Switzerland with aggregated BLEU scores of Swiss-German variety per region (Fig. 5)
    Map of Italy with aggregated BLEU scores of Italian variety per region (Fig. 6)
Task Specific Plot for Maximum Scores (Figs. 7 to 10)
    Parts-of-Speech Tagging (Fig. 6(a))
    Named entity recognition (Fig. 6(b))
    Topic classification (Fig. 6(c))
    Natural language inference (Fig. 7(a))
    Extractive question answering (Fig. 7(b))
    Multiple-choice machine reading comprehension (Fig. 7(c))
    Sentiment analysis (Fig. 8(a))
    Dialect identification (Fig. 8(b))
    Machine translation (MT-region) (Fig. 9(a))
    Machine translation (MT-dialect) (Fig. 9(b))
Dialectal Gap visualization for language clusters utilizing zero-shot cross-lingual transfer from Standard English.
    Parts-of-Speech Tagging (Fig. 10(a))
    Named entity recognition (Fig. 10(b))
    Dialect identification (Fig. 10(c))
Appendix E Task Specific Evaluation Result Tables
    Dependency parsing (Table 10)
    Parts-of-Speech tagging (Table 11)
    Named entity recognition (Table 12)
    Natural language inference (Table 13)
    Extractive question answering (Tables 14 to 15)
    Dialect identification (Table 16)
    Topic classification (Table 17)
    Sentiment analysis (Table 18)
    Multiple-choice machine reading comprehension (Table 19)
    Machine translation (Table 20)
Appendix F Highest performing and lowest performing varieties (LABEL:{tab:high-lang})
Appendix G Low-resource Variety performing better in zero-shot NER (Table 22)
Appendix H Cluster level Result Summaries with Demographic Utility and Standard Deviation report (Tables 23 to 31)
Appendix I In-Context Learning Details
Table 6: Table of contents for supplementary material.
Appendix A Related Work
The majority of multilingual benchmarks Hu et al. (2020); Ruder et al. (2023, 2021); Liang et al. (2020); Wilie et al. (2020); Park et al. (2021) have heavily focused on dominant language varieties. However, the performance disparity between standard languages and their dialectal counterparts has been studied Kantharuban et al. (2023); Ziems et al. (2022); Rios (2020); Ziems et al. (2023) and shown to be significantly large across different tasks. Still, the large-scale generalization of this finding comprising numerous task categories is yet to be done. There have been previous efforts to bridge this gap though. Some work has focused on curating dialectal datasets across several tasks within one language cluster, while others have focused on single tasks across many language clusters. For instance, ORCA Elmadany et al. (2023), ARLUE Abdul-Mageed et al. (2021) and ALUE Seelawi et al. (2021) are dedicated natural language understanding benchmarks focusing on Arabic varieties alone. Multi-VALUE Ziems et al. (2023) was developed for benchmarking NLP tasks in English varieties. Mirzakhalov et al. (2021a) is a suite of resources for benchmarking MT in several Turkic languages. There has also been a large body of work on dialect identification across several languages Jauhiainen et al. (2022); Hämäläinen et al. (2021). Recently CODET Alam et al. (2023) was released as a contrastive dialectal MT benchmark covering 882 different variations from nine different languages. To the best of our knowledge, we are the first to do a large scale aggregation of dialectal data across several language clusters and tasks.
Appendix B Tasks of DialectBench
DialectBench includes 10 NLP tasks falling into four broader categories: structured prediction, sentence classification, question answering, and text generation. In Table 1, we present statistics for the datasets for each task and briefly discuss each task below.
Dependency parsing
For the dependency parsing task, we include only those Universal Dependencies (UD) Zeman et al. (2021) languages that have dialectal data. Beyond the data available in UD 2.12, we incorporate two additional datasets for African-American English (AAVE) Twitter data Blodgett et al. (2018) and Singlish Wang et al. (2017). To make these two datasets compatible with the UD processing pipeline, we replace the original dependency labels with the labels corresponding to the official UD formalism.
Part of speech (POS) tagging
We use the same UD languages for POS tagging that we used for dependency parsing. At the same time, we use the POS data instances from Singlish Wang et al. (2017). Moreover, we include six Finnish dialects, four Arabic dialects and Occitan through the UPOS label standardized pipeline proposed by Blaschke et al. (2023).
Named entity recognition (NER)
We use data from the 176-language version of the Wikiann dataset processed by Rahimi et al. (2019). All these languages provide both training and test data. In addition, we include dialectal data from the original Wikiann dataset (282 languages) Pan et al. (2017) for evaluation. Moreover, we include three Norwegian dialects Johansen (2019) with train, test and validation datasets that use a slightly different set of NER tags (GEO, ORG, OTH, PER) compared to the one we use in Wikiann (LOC, ORG, PER). We leave these levels as it is and do not convert to the Wikiann tags.
Dialect identification
We include dialect identification experiments on Arabic, Greek, Portuguese, English, Spanish, and Swiss German dialectal datasets. In these datasets, we find large variations in the level of granularity with which dialects are classified. For instance, the MADAR corpus differentiates Arabic varieties at the city level Bouamor et al. (2018), whereas our English and Spanish datasets are labeled with country names Zampieri et al. (2023).
Sentiment classification
Here we include several different Arabic varieties. Like other dialectal datasets, these datasets do not follow one standard labeling process. However, all datasets contain two main sentiment types: positive and negative. A number of datasets contain additional labels such as “objective” or “neutral.” In our setting, we perform a further split of data to provide validation data for each dialect. However, we do not remove these extra labeled data for information preservation.
Topic classification
We use the SIB-200 dataset Adelani et al. (2023) for topic classification task SIB-200 was constructed from the FLORES-200 translation datasets. The authors annotated the English dataset of FLORES-200 with 6 topic labels and then further propagated these labels to the translated instances for all other languages. For our case of benchmarking dialectal segments, we use the dialectal and regional varieties from SIB-200.
Variety # Sentences
cluster Language code
anglic eng_Latn english 5010
arabic acm_Arab north mesopotamian arabic 5010
acq_Arab ta’izzi-adeni arabic 5010
aeb_Arab tunisian arabic 5010
ajp_Arab south levantine arabic 5010
apc_Arab levantine arabic (a:north) 5010
arb_Arab standard arabic 5010
ars_Arab najdi arabic 5010
ary_Arab moroccan arabic 5010
arz_Arab egyptian arabic 5010
common turkic azb_Arab south azerbaijani 5010
azj_Latn north azerbaijani 5010
tur_Latn central oghuz (m:spoken) 5010
gallo-italian lij_Latn ligurian 5010
lmo_Latn lombard 5010
vec_Latn venetian 5010
gallo-rhaetian fur_Latn friulian 5010
high german lim_Latn limburgan 5010
ltz_Latn luxemburgish 5010
italian romance ita_Latn italian 5010
scn_Latn sicilian 5010
kurdish ckb_Arab central kurdish 5010
kmr_Latn northern kurdish 5010
latvian ltg_Latn east latvian 5010
lvs_Latn latvian 5010
modern dutch nld_Latn dutch 5010
norwegian nno_Latn norwegian nynorsk (m:written) 5010
nob_Latn norwegian bokmål (m:written) 5010
sardo-corsican srd_Latn sardinian 5010
sinitic yue_Hant cantonese 5010
zho_Hans classical-middle-modern sinitic (o:simplified) 5010
zho_Hant classical-middle-modern sinitic (o:traditional) 5010
sotho-tswana (s.30) nso_Latn northern sotho 5010
sot_Latn southern sotho 5010
southwestern shifted romance glg_Latn galician 5010
oci_Latn occitan 5010
por_Latn portuguese (a:european) 5010
spa_Latn spanish 5010
Table 7: Data statistics for newly created translate-test natural language inference (NLI) dataset. We prepare this translate-test NLI dataset by translating XNLI Conneau et al. (2018) english evaluation dataset.
Natural language inference
For the natural language inference task, there is no existing dataset with varieties. So we use the existing English test set of XNLI Conneau et al. (2018) and construct a multilingual dialect-focused translated evaluation dataset. We use a state-of-the-art machine translation model (NLLB-200 3B) to translate the English test set to 12 language clusters encompassing 40 varieties (Complete data statistics are reported in Table 7). After that, we perform zero-shot cross-lingual transfer from the English finetuned NLI model. We refer to this setting as translate-test.
Multiple-choice machine reading comprehension
This task aims to evaluate the capability of multiple-choice question answering given a context passage. The question could be answered by understanding the context passage while the right answer is given at one of the multiple choices. We use the Chinese, Sotho and Arabic clusters data from the recently released Belebele MRC dataset Bandarkar et al. (2023). This is an evaluation-only dataset.
Extractive question answering
This task predicts the answer span given a question and context passage. We use the SD-QA dialectal question-answering dataset Faisal et al. (2021). SD-QA is an evaluation dataset built on top of TyDiQA Clark et al. (2020), another well-known typologically diverse question-answering dataset. SD-QA contains the spoken utterances and transcription of the original TyDiQA question from speakers of English, Bengali, Arabic, Korean, and Swahili. We only use the textual part that contains the transcription of the dialectal spoken question matching the original TyDiQA question text. Note that since the transcriptions of the questions are obtained through automatic speech recognition, they may include both dialectal variations and noise due to ASR transcription errors.
Machine translation
We evaluate variety translation using the CODET benchmark Alam et al. (2023) and the TIL MT corpus Mirzakhalov et al. (2021b). CODET contains a contrastive dataset of 882 different varieties from nine different languages. We evaluate dialects here at city level for all languages except Italian and Swiss-German and aggregate dialects at the region level for them. The TIL corpus contains parallel translations across 22 Turkic languages, but in our evaluations we only include 8 turkic languages (Turkic, Sakha, Kazakh, Karakalpak, Bashkir, Azerbaijani, Kyrgyz) that have parallel English translations.
Appendix C Varieties and Clusters of DialectBench
C.1 DialectBench Variety list
Table 8: Varieties represented in DialectBench.
Language cluster
Variety name
Glottocode mBERT seen UDP POS NER SDQA RCMC NLI TC SC DI MT
albanian
albanian
alba1267 - - - - - - - -
gheg albanian
gheg1238 - - - - - - - - -
anglic
african american vernacular english
afri1276 - - - - - - - - - -
australian english
aust1314 - - - - - - - - - -
english
stan1293 - - -
english (a:scotland)
stan1293 - - - - - - - - - -
english (a:uk)
stan1293 - - - - - - - - - -
english (o:controlled)
stan1293 - - - - - - - - - -
indian english (a:north)
indi1255 - - - - - - - - - -
indian english (a:south)
indi1255 - - - - - - - - - -
irish english
iris1255 - - - - - - - - - -
jamaican creole english
jama1262 - - - - - - - - - -
kenyan english
keny1281 - - - - - - - - - -
new zealand english
newz1240 - - - - - - - - - -
nigerian english
nige1260 - - - - - - - - - -
north american english
nort3314 - - - - - - - - - -
old english (ca. 450-1100)
olde1238 - - - - - - - - - -
philippine english
phil1246 - - - - - - - - - -
singlish
sing1272 - - - - - - - - -
southeast american english
sout3300 - - - - - - - - - -
southern african english
sout3331 - - - - - - - - - -
arabic
aleppo
alep1241 - - - - - - - - -
algerian arabic
alge1239 - - - - - - -
arabian peninsula arabic (a:yemen)
arab1393 - - - - - - - - -
arabic (a:bahrain)
arabic (a:jordan)
arabic (a:saudi-arabia)
egyptian arabic
egyp1253 - - - -
egyptian arabic (a:alx)
egyp1253 - - - - - - - - -
egyptian arabic (a:asw)
egyp1253 - - - - - - - - -
egyptian arabic (a:cai)
egyp1253 - - - - - - - - -
egyptian arabic (a:kha)
egyp1253 - - - - - - - - -
fez. meknes
fezm1238 - - - - - - - - -
gilit mesopotamian arabic
meso1252 - - - - - - - - -
gulf arabic
gulf1241 - - - - - - - - - -
gulf arabic (a:doh)
gulf1241 - - - - - - - - -
gulf arabic (a:jed)
gulf1241 - - - - - - - - -
gulf arabic (a:mus)
gulf1241 - - - - - - - - -
gulf arabic (a:riy)
gulf1241 - - - - - - - - -
levantine arabic
nort3139 - - - - - - - - - -
levantine arabic (a:north)
nort3139 - - - - - - - -
levantine arabic (a:north-dam)
nort3139 - - - - - - - - -
libyan arabic
liby1240 - - - - - - - - - - -
libyan arabic (a:ben)
liby1240 - - - - - - - - -
moroccan arabic
moro1292 - - - - - -
najdi arabic
najd1235 - - - - - - - -
north african arabic
nort3191 - - - - - - - - -
north mesopotamian arabic
nort3142 - - - - - - - -
north mesopotamian arabic (a:bas)
nort3142 - - - - - - - - -
north mesopotamian arabic (a:mos)
nort3142 - - - - - - - - -
rabat-casablanca arabic
raba1252 - - - - - - - - -
sfax
sfax1238 - - - - - - - - -
south levantine arabic
sout3123 - - - - - -
south levantine arabic (a:south-amm)
sout3123 - - - - - - - - -
south levantine arabic (a:south-jer)
sout3123 - - - - - - - - -
south levantine arabic (a:south-sal)
sout3123 - - - - - - - - -
standard arabic
stan1318 - -
sunni beiruti arabic
sunn1238 - - - - - - - - -
ta’izzi-adeni arabic
taiz1242 - - - - - - - - -
tripolitanian arabic
trip1239 - - - - - - - - -
tunisian arabic
tuni1259 - - - - - - -
tunisian arabic (a:tun)
tuni1259 - - - - - - - - -
tunisian arabic (r:casual)
tuni1259 - - - - - - - - - -
basque
basque (a:amenduze)
basq1248 - - - - - - - - -
basque (a:azkaine)
basq1248 - - - - - - - - - -
basque (a:baigorri)
basq1248 - - - - - - - - - -
basque (a:barkoxe)
basq1248 - - - - - - - - - -
basque (a:donibane)
basq1248 - - - - - - - - - -
basque (a:garruze)
basq1248 - - - - - - - - - -
basque (a:iholdi)
basq1248 - - - - - - - - - -
basque (a:jatsu)
basq1248 - - - - - - - - - -
basque (a:jutsi)
basq1248 - - - - - - - - - -
basque (a:larzabale)
basq1248 - - - - - - - - - -
basque (a:luhuso)
basq1248 - - - - - - - - - -
basque (a:sara)
basq1248 - - - - - - - - - -
basque (a:senpere)
basq1248 - - - - - - - - - -
basque (a:suhuskune)
basq1248 - - - - - - - - - -
basque (a:uharte)
basq1248 - - - - - - - - - -
navarro-labourdin basque (a:behorlegi)
basq1249 - - - - - - - - - -
navarro-labourdin basque (a:bidarrai)
basq1249 - - - - - - - - - -
navarro-labourdin basque (a:helete)
basq1249 - - - - - - - - - -
navarro-labourdin basque (a:mugerre)
basq1249 - - - - - - - - - -
navarro-labourdin basque (a:urruna)
basq1249 - - - - - - - - - -
souletin (a:maule)
basq1250 - - - - - - - - - -
bengali
vanga (a:barisal)
vang1242 - - - - - - - - - -
vanga (a:dhaka)
vang1242 - - - - - - - -
vanga (a:jessore)
vang1242 - - - - - - - - - -
vanga (a:khulna)
vang1242 - - - - - - - - - -
vanga (a:kushtia)
vang1242 - - - - - - - - - -
vanga (a:west bengal)
vang1242 - - - - - - - - -
circassian
adyghe
adyg1241 - - - - - - - - - -
kabardian
kaba1278 - - - - - - - - - -
common turkic
bashkir
bash1264 - - - - - - - - -
central oghuz
azer1255 - - - - - - - - - -
central oghuz (m:spoken)
azer1255 - - - - - - - - -
crimean tatar
crim1257 - - - - - - - - - -
kara-kalpak
kara1467 - - - - - - - - - -
kazakh
kaza1248 - - - - - - - - -
kirghiz
kirg1245 - - - - - - - - -
north azerbaijani
nort2697 - - - - - - -
sakha
yaku1245 - - - - - - - - - -
south azerbaijani
sout2697 - - - - - - -
turkish
nucl1301 - - - - - - - -
uzbek
uzbe1247 - - - - - - - - -
eastern romance
aromanian
arom1237 - - - - - - - - - -
moldavian
mold1248 - - - - - - - - - -
romanian
roma1327 - - - - - - - - -
eastern-western armenian
eastern armenian
nucl1235 - - - - - - - -
western armenian
homs1234 - - - - - - - - -
farsic
dari
dari1249 - - - - - - - - - -
frisian
ems-weser frisian
sate1242 - - - - - - - - - -
northern frisian
nort2626 - - - - - - - - - -
western frisian
west2354 - - - - - - - - -
gallo-italian
emiliano-romagnolo
emil1243 - - - - - - - - - -
ligurian
ligu1248 - - - - - -
lombard
lomb1257 - - - - - - -
piemontese
piem1238 - - - - - - - - -
venetian
vene1258 - - - - - - - -
gallo-rhaetian
anglo-norman
angl1258 - - - - - - - - - -
arpitan
fran1260 - - - - - - - - - -
french
stan1290 - - - - - - -
french (a:paris)
stan1290 - - - - - - - - -
friulian
friu1240 - - - - - - - -
old french (842-ca. 1400)
oldf1239 - - - - - - - - -
romansh
roma1326 - - - - - - - - - -
walloon
wall1255 - - - - - - - - - -
greater panjabic
eastern panjabi
panj1256 - - - - - - - - -
western panjabi
west2386 - - - - - - - - -
greek
apulian greek
apul1237 - - - - - - - - - -
cretan
cret1244 - - - - - - - - - -
cypriot greek
cypr1249 - - - - - - - - - - -
cypriot greek (r:casual, m:written, i:fb)
cypr1249 - - - - - - - - - -
cypriot greek (r:casual, m:written, i:other)
cypr1249 - - - - - - - - - -
cypriot greek (r:casual, m:written, i:twitter)
cypr1249 - - - - - - - - - -
modern greek
mode1248 - - - - - - - - -
modern greek (r:casual, m:written, i:fb)
mode1248 - - - - - - - - - -
modern greek (r:casual, m:written, i:other)
mode1248 - - - - - - - - - -
modern greek (r:casual, m:written, i:twitter)
mode1248 - - - - - - - - - -
pontic
pont1253 - - - - - - - - - -
high german
bavarian
bava1246 - - - - - - - - -
central alemannic
swis1247 - - - - - - - - - -
central alemannic (a:ag)
swis1247 - - - - - - - - - -
central alemannic (a:ai)
swis1247 - - - - - - - - - -
central alemannic (a:ar)
swis1247 - - - - - - - - - -
central alemannic (a:be)
swis1247 - - - - - - - - -
central alemannic (a:bl)
swis1247 - - - - - - - - - -
central alemannic (a:bs)
swis1247 - - - - - - - - -
central alemannic (a:fr)
swis1247 - - - - - - - - - -
central alemannic (a:gl)
swis1247 - - - - - - - - - -
central alemannic (a:gr)
swis1247 - - - - - - - - - -
central alemannic (a:lu)
swis1247 - - - - - - - - -
central alemannic (a:nw)
swis1247 - - - - - - - - - -
central alemannic (a:ow)
swis1247 - - - - - - - - - -
central alemannic (a:sg)
swis1247 - - - - - - - - - -
central alemannic (a:sh)
swis1247 - - - - - - - - - -
central alemannic (a:so)
swis1247 - - - - - - - - - -
central alemannic (a:sz)
swis1247 - - - - - - - - - -
central alemannic (a:tg)
swis1247 - - - - - - - - - -
central alemannic (a:ur)
swis1247 - - - - - - - - - -
central alemannic (a:vs)
swis1247 - - - - - - - - - -
central alemannic (a:zg)
swis1247 - - - - - - - - - -
central alemannic (a:zh)
swis1247 - - - - - - -
central bavarian
cent1967 - - - - - - - - - -
german
stan1295 - - - - - - -
kölsch
kols1241 - - - - - - - - - -
limburgan
limb1263 - - - - - - - -
luxemburgish
luxe1243 - - - - - - -
pennsylvania german
penn1240 - - - - - - - - - -
pfaelzisch-lothringisch
pala1330 - - - - - - - - - -
upper saxon
uppe1465 - - - - - - - - - -
hindustani
fiji hindi
fiji1242 - - - - - - - - - -
hindi
hind1269 - - - - - - - - -
inuit
alaskan inupiaq
inup1234 - - - - - - - - - -
kalaallisut
kala1399 - - - - - - - - - -
italian romance
continental southern italian
neap1235 - - - - - - - -
italian
ital1282 - - - - -
italian (a:abruzzo)
ital1282 - - - - - - - - - -
italian (a:basilicata)
ital1282 - - - - - - - - - -
italian (a:calabria)
ital1282 - - - - - - - - - -
italian (a:campania)
ital1282 - - - - - - - - - -
italian (a:emilia-romagna)
ital1282 - - - - - - - - - -
italian (a:friuli-venezia giulia)
ital1282 - - - - - - - - - -
italian (a:lazio)
ital1282 - - - - - - - - - -
italian (a:liguria)
ital1282 - - - - - - - - - -
italian (a:lombardia)
ital1282 - - - - - - - - - -
italian (a:marche)
ital1282 - - - - - - - - - -
italian (a:molise)
ital1282 - - - - - - - - - -
italian (a:piemonte)
ital1282 - - - - - - - - - -
italian (a:puglia)
ital1282 - - - - - - - - - -
italian (a:sardegna)
ital1282 - - - - - - - - - -
italian (a:sicilia)
ital1282 - - - - - - - - - -
italian (a:toscana)
ital1282 - - - - - - - - - -
italian (a:trentino-alto adige/südtirol)
ital1282 - - - - - - - - - -
italian (a:umbria)
ital1282 - - - - - - - - - -
italian (a:unknown)
ital1282 - - - - - - - - - -
italian (a:veneto)
ital1282 - - - - - - - - - -
italian (r:casual, m:written, i:tweet)
ital1282 - - - - - - - -
italian (r:formal, m:written, i:essay)
ital1282 - - - - - - - -
italian romance (a:apulia, m:spoken, i:tarantino)
sicilian
sici1248 - - - - - - -
komi
komi
komi1267 - - - - - - - - - -
komi-permyak
komi1269 - - - - - - - -
komi-zyrian (m:spoken)
komi1268 - - - - - - - - -
komi-zyrian (m:written)
komi1268 - - - - - - - - -
korean
korean (a:south-eastern, m:spoken)
kore1280 - - - - - - - - - -
seoul (m:spoken)
seou1239 - - - - - - - - -
kurdish
central kurdish
cent1972 - - - - - - - -
kurdish
kurd1259 - - - - - - - - - -
northern kurdish
nort2641 - - - - - - - - -
sine’i
sine1239 - - - - - - - - - -
sorani
sora1257 - - - - - - - - - -
latvian
east latvian
east2282 - - - - - - - -
latvian
latv1249 - - - - - - -
mari
eastern mari
east2328 - - - - - - - - - -
western mari
west2392 - - - - - - - - - -
modern dutch
dutch
dutc1256 - - - - - - -
western flemish
vlaa1240 - - - - - - - - - -
zeeuws
zeeu1238 - - - - - - - - - -
neva
central and north pohjanmaa (a:ostrobothnian)
cent1985 - - - - - - - - - -
estonian
esto1258 - - - - - - - - -
finnish
finn1318 - - - - - - - - -
häme (a:tavastian)
hame1240 - - - - - - - - - -
neva (a:south-west trans)
savo (a:savonian)
savo1254 - - - - - - - - - -
southeastern finnish (a:south-east)
sout1743 - - - - - - - - - -
southwestern finnish (a:south-west)
sout2677 - - - - - - - - - -
norwegian
norwegian
norw1258 - - - - - - - - - -
norwegian (a:eastern)
norw1258 - - - - - - - - - -
norwegian (a:setesdal)
norw1258 - - - - - - - - - -
norwegian (a:southwestern)
norw1258 - - - - - - - - - -
norwegian (m:written, i:samnorsk)
norwegian bokmål (m:written)
norw1259 - - - - -
norwegian nynorsk (m:written)
norw1262 - - - - -
norwegian nynorsk (m:written, i:old)
norw1262 - - - - - - - - -
saami
north saami
nort2671 - - - - - - - - -
skolt saami
skol1241 - - - - - - - - -
sabellic
umbrian
umbr1253 - - - - - - - - -
sardo-corsican
corsican
cors1241 - - - - - - - - - -
sardinian
sard1257 - - - - - - - -
serbian-croatian-bosnian
bosnian standard
bosn1245 - - - - - - - - -
croatian standard
croa1245 - - - - - - - - -
serbian standard
serb1264 - - - - - - - - -
serbian-croatian-bosnian
sout1528 - - - - - - - - -
sinitic
cantonese
cant1236 - - - - - - - -
classical chinese
lite1248 - - - - - - - -
classical-middle-modern sinitic (a:hongkong, o:traditional)
clas1255 - - - - - - - - -
classical-middle-modern sinitic (o:simplified)
clas1255 - - - - -
classical-middle-modern sinitic (o:traditional)
clas1255 - - - - - - -
hakka chinese
hakk1236 - - - - - - - - - -
mandarin chinese
mand1415 - - - - - - - - - -
mandarin chinese (a:mainland, o:simplified)
mand1415 - - - - - - - - - -
mandarin chinese (a:mainland, o:traditional, i:synthetic)
mand1415 - - - - - - - - - -
mandarin chinese (a:taiwan, o:simplified)
mand1415 - - - - - - - - - -
mandarin chinese (a:taiwan, o:traditional, i:synthetic)
mand1415 - - - - - - - - - -
min nan chinese
minn1241 - - - - - - - - - -
wu chinese
wuch1236 - - - - - - - - - -
sorbian
lower sorbian
lowe1385 - - - - - - - - - -
upper sorbian
uppe1395 - - - - - - - - - -
sotho-tswana (s.30)
northern sotho
nort3233 - - - - - - -
southern sotho
sout2807 - - - - - - -
southwestern shifted romance
brazilian portuguese
braz1246 - - - - - - - -
extremaduran
extr1243 - - - - - - - - - -
galician
gali1258 - - - - - - -
latin american spanish
amer1254 - - - - - - - - - -
mirandese
mira1251 - - - - - - - - - -
occitan
occi1239 - - - - -
portuguese (a:european)
port1283 - - - - - -
portuguese (i:mix)
port1283 - - - - - - - - -
portuguese (m:written)
port1283 - - - - - - -
spanish
stan1288 - - - - - -
spanish (a:europe)
stan1288 - - - - - - - - - -
swahili
swahili
swah1253 - - - - - - - - - -
swahili (a:kenya)
swah1253 - - - - - - - - - -
swahili (a:tanzania)
swah1253 - - - - - - - - - -
tigrinya
tigrinya (a:eritrea)
tigr1271 - - - - - - - - - -
tigrinya (a:ethiopia)
tigr1271 - - - - - - - - - -
tupi-guarani subgroup i.a
mbyá guaraní (a:brazil)
mbya1239 - - - - - - - - -
mbyá guaraní (a:paraguay)
mbya1239 - - - - - - - - -
old guarani
oldp1258 - - - - - - - - -
west low german
west low german
west2357 - - - - - - -
yoruba
yoruba (a:central nigeria)
yoru1245 - - - - - - - - -
C.2 Language Clusters and Representative Varieties
Table 9: Language clusters and their clusters representatives.
Task Cluster Name Cluster Representative
DEP. Parsing albanian albanian
arabic standard arabic
eastern-western armenian western armenian
sinitic classical-middle-modern sinitic (o:simplified)
anglic english
gallo-rhaetian french
high german german
tupi-guarani subgroup i.a old guarani
italian romance italian (r:formal, m:written, i:essay)
komi komi-zyrian (m:spoken)
norwegian norwegian bokmål (m:written)
southwestern shifted romance brazilian portuguese
saami north saami
POS Tagging albanian albanian
anglic english
arabic standard arabic
eastern-western armenian western armenian
gallo-rhaetian french
high german german
italian romance italian (r:formal, m:written, i:essay)
komi komi-zyrian (m:spoken)
neva finnish
norwegian norwegian bokmål (m:written)
saami north saami
sinitic classical-middle-modern sinitic (o:simplified)
southwestern shifted romance brazilian portuguese
tupi-guarani subgroup i.a mbyá guaraní (a:paraguay)
NER anglic english
arabic standard arabic
circassian adyghe
common turkic turkish
eastern romance romanian
frisian western frisian
gallo-italian piemontese
gallo-rhaetian french
greater panjabic western panjabi
greek modern greek
high german german
hindustani hindi
inuit alaskan inupiaq
italian romance italian
komi komi
kurdish kurdish
latvian latvian
mari eastern mari
modern dutch dutch
norwegian norwegian bokmål (m:written)
sardo-corsican sardinian
serbian-croatian-bosnian croatian standard
sinitic mandarin chinese
sorbian lower sorbian
sotho-tswana (s.30) southern sotho
southwestern shifted romance spanish
EQA anglic southeast american english
arabic arabic (a:saudi-arabia)
bengali vanga (a:dhaka)
korean seoul (m:spoken)
swahili swahili (a:kenya)
MRC anglic english
arabic standard arabic
sinitic classical-middle-modern sinitic (o:simplified)
sotho-tswana (s.30) northern sotho
TC anglic english
arabic standard arabic
common turkic north azerbaijani
gallo-italian venetian
high german luxemburgish
italian romance italian
kurdish northern kurdish
latvian latvian
norwegian norwegian bokmål (m:written)
sinitic classical-middle-modern sinitic (o:simplified)
sotho-tswana (s.30) northern sotho
southwestern shifted romance spanish
SC arabic standard arabic
MT-Dialect arabic gulf arabic (a:riy)
basque basque (a:azkaine)
bengali vanga (a:dhaka)
high german central bavarian
kurdish sorani
norwegian norwegian (a:eastern)
tigrinya tigrinya (a:ethiopia)
common turkic turkish
greek cretan
MT-Region italian romance italian (a:umbria)
high german central alemannic (a:zh)
DId anglic english (a:uk)
arabic standard arabic
greek modern greek (r:casual, m:written, i:twitter)
high german central alemannic (a:zh)
sinitic mandarin chinese (a:mainland, o:simplified)
southwestern shifted romance brazilian portuguese
NLI anglic english
arabic standard arabic
common turkic north azerbaijani
gallo-italian venetian
high german luxemburgish
italian romance italian
kurdish northern kurdish
latvian latvian
norwegian norwegian bokmål (m:written)
sinitic classical-middle-modern sinitic (o:simplified)
sotho-tswana (s.30) northern sotho
southwestern shifted romance spanish
Appendix D Result Visualizations
D.1 Regional maps with aggregated Machine Translation scores
Figure 5: Map of Switzerland with aggregated BLEU scores of Swiss-German variety per region
Figure 6: Map of Italy with aggregated BLEU scores of Italian variety per region
D.2 Task Specific Plot for Maximum Scores
(a) Parts of Speech Tagging
(b) Named Entity Recognition
(c) Topic Classification
Figure 7: Task specific plot of maximum obtainable score for all varieties. The yellow-shaded region represents language clusters having no varieties seen during mBERT pertaining. The bars with colored stripes represent the standard variety of a cluster
(a) Natural Language Inference
(b) Extractive Question Answering
(c) Multiple Choice Machine Reading Comprehension
Figure 8: Task specific plot of maximum obtainable Linguistic Utility for all varieties. The yellow-shaded region represents language clusters having no varieties seen during mBERT pertaining. The bars with colored stripes represent the standard variety of a cluster. The dialect with the Rawlsian score in each cluster is that with the leftmost bar.
(a) Sentiment Analysis
(b) Dialect Identification
Figure 9: Task specific plot of maximum obtainable Linguistic Utility for all varieties. The yellow-shaded region represents language clusters having no varieties seen during mBERT pertaining. The bars with colored stripes represent the standard variety of a cluster. The dialect with the Rawlsian score in each cluster is that with the leftmost bar.
(a) Machine Translation (Region level aggregation)
(b) Machine Translation (Variety level)
Figure 10: Task specific plot of maximum obtainable Linguistic Utility for all varieties. The yellow-shaded region represents language clusters having no varieties seen during mBERT pertaining. The bars with colored stripes represent the standard variety of a cluster. The dialect with the Rawlsian score in each cluster is that with the leftmost bar.
D.3 Dialectal Gap visualizations (zeroshot)
(a) Parts-of-speech (POS) tagging
(b) Named entity recognition (NER)
(c) Natural language inference (NLI)
Figure 11: Dialectal Gap visualization for language clusters utilizing zero-shot cross-lingual transfer from Standard English. The x-axis is for cluster gap while comparing against Standard English variety. Values far from zero have a larger performance gap from English. The y-axis is for aggregated cluster gap while comparing against standard cluster variety and values far from zero have larger within cluster gap. Ideally, we want both of them to be close to zero.
Appendix E Evaluation results
cluster variety
UD-code
Zeroshot (mBERT)
Zeroshot (XLM-R)
Finetune (mBERT)
Finetune (XLM-R)
albanian
albanian
UD_Albanian-TSA
81.78
83.08
gheg albanian
UD_Gheg-GPS
38.14
43.50
anglic
english
UD_English-EWT
91.55
91.10
91.55
91.1
singlish
singlish
69.88
68.55
82.36
11.0
african american vernacular english
TwitterAAE
50.53
52.53
arabic
standard arabic
UD_Arabic-PADT
50.87
55.91
88.47
2.48
south levantine arabic
UD_South_Levantine_Arabic-MADAR
49.94
45.75
north african arabic
UD_Maghrebi_Arabic_French-Arabizi
34.33
28.04
5.46
54.7
eastern-western armenian
western armenian
UD_Western_Armenian-ArmTDP
54.63
59.71
88.84
89.29
eastern armenian
UD_Armenian-ArmTDP
53.27
62.63
85.55
87.01
gallo-italian
ligurian
UD_Ligurian-GLT
50.22
43.78
13.06
9.1
gallo-rhaetian
french
UD_French-ParTUT
80.87
82.17
93.93
92.09
french (a:paris)
UD_French-ParisStories
57.64
61.36
77.54
77.54
old french (842-ca. 1400)
UD_Old_French-SRCMF
56.12
46.52
91.51
89.78
high german
german
UD_German-LIT
72.65
76.50
central alemannic (a:zh)
UD_Swiss_German-UZH
36.77
34.70
italian romance
italian
UD_Italian-PUD
78.99
78.58
italian (r:formal, m:written, i:essay)
UD_Italian-MarkIT
76.83
79.38
86.34
83.81
italian (r:casual, m:written, i:tweet)
UD_Italian-PoSTWITA
61.88
62.13
85.82
86.32
continental southern italian
UD_Neapolitan-RB
30.00
50.00
komi
komi-zyrian (m:spoken)
UD_Komi_Zyrian-IKDP
26.89
32.14
komi-permyak
UD_Komi_Permyak-UH
26.12
30.91
komi-zyrian (m:written)
UD_Komi_Zyrian-Lattice
21.01
27.55
norwegian
norwegian bokmål (m:written)
UD_Norwegian-Bokmaal
79.55
82.95
93.57
93.49
norwegian nynorsk (m:written)
UD_Norwegian-Nynorsk
76.45
76.76
93.15
93.14
norwegian nynorsk (m:written, i:old)
UD_Norwegian-NynorskLIA
56.58
56.08
78.39
8.25
saami
north saami
UD_North_Sami-Giella
23.58
16.87
67.56
5.96
skolt saami
UD_Skolt_Sami-Giellagas
19.41
28.28
sabellic
umbrian
UD_Umbrian-IKUVINA
33.21
28.75
sinitic
classical-middle-modern sinitic (a:hongkong, o:traditional)
UD_Chinese-HK
58.97
61.78
classical-middle-modern sinitic (o:simplified)
UD_Chinese-GSDSimp
53.35
52.21
87.85
87.37
classical chinese
UD_Classical_Chinese-Kyoto
46.72
35.14
19.65
18.11
southwestern shifted romance
portuguese (a:european)
UD_Portuguese-PUD
76.67
77.28
portuguese (i:mix)
UD_Portuguese-Bosque
75.89
77.96
92.64
92.15
brazilian portuguese
UD_Portuguese-GSD
73.30
74.21
94.37
93.75
portuguese (m:written)
UD_Portuguese-CINTIL
69.34
72.94
83.35
84.31
tupi-guarani subgroup i.a
old guarani
UD_Guarani-OldTuDeT
22.58
29.03
mbyá guaraní (a:paraguay)
UD_Mbya_Guarani-Thomas
13.51
11.15
mbyá guaraní (a:brazil)
UD_Mbya_Guarani-Dooley
8.95
4.23
west low german
west low german
UD_Low_Saxon-LSDC
40.75
37.38
Table 10: Dependency parsing evaluation report comprising zeroshot score and in-language finetuning. We report UAS as evaluation score. Zeroshot scores are evaluated using model finetuned on Standerd English. If training data is not available, we skip those languages (mentioned as ’-’) for in-language finetuning.
cluster variety Dataset-code dataset
Zeroshot (mBERT)
Zeroshot (XLM-R)
Finetune (mBERT)
Finetune (XLM-R)
albanian
albanian
UD_Albanian-TSA ud
75.80
84.41
gheg albanian
UD_Gheg-GPS ud
48.96
55.84
anglic
english
UD_English-EWT ud
96.41
97.16
96.41
97.16
singlish
singlish ud
76.27
77.55
87.38
87.96
arabic
south levantine arabic
UD_South_Levantine_Arabic-MADAR ud
51.99
61.84
standard arabic
UD_Arabic-PADT ud
39.74
56.67
95.72
96.11
gulf arabic
dar-glf noisy
38.84
50.12
egyptian arabic
dar-egy noisy
36.14
51.39
levantine arabic
dar-lev noisy
32.66
43.76
north african arabic
UD_Maghrebi_Arabic_French-Arabizi ud
28.30
26.01
67.89
59.29
eastern-western armenian
eastern armenian
UD_Armenian-ArmTDP ud
71.78
82.63
91.31
92.36
western armenian
UD_Western_Armenian-ArmTDP ud
70.27
75.31
94.86
95.14
gallo-italian
ligurian
UD_Ligurian-GLT ud
58.90
52.78
13.16
5.09
gallo-rhaetian
french
UD_French-ParTUT ud
84.36
85.47
96.88
96.42
french (a:paris)
UD_French-ParisStories ud
81.37
82.77
96.70
96.76
old french (842-ca. 1400)
UD_Old_French-SRCMF ud
64.70
59.41
95.64
95.64
high german
german
UD_German-LIT ud
87.08
88.36
central alemannic (a:zh)
UD_Swiss_German-UZH ud
62.56
47.18
italian romance
italian
UD_Italian-PUD ud
81.09
83.12
italian (r:formal, m:written, i:essay)
UD_Italian-MarkIT ud
80.00
81.87
93.35
92.70
italian (r:casual, m:written, i:tweet)
UD_Italian-PoSTWITA ud
73.71
76.45
95.80
96.83
continental southern italian
UD_Neapolitan-RB ud
30.00
57.14
komi
komi-zyrian (m:spoken)
UD_Komi_Zyrian-IKDP ud
41.25
46.66
komi-permyak
UD_Komi_Permyak-UH ud
29.52
43.67
komi-zyrian (m:written)
UD_Komi_Zyrian-Lattice ud
20.40
35.12
neva
finnish
UD_Finnish-TDT ud
81.29
86.21
96.05
97.76
estonian
UD_Estonian-EDT ud
80.34
85.17
96.49
97.20
hame (a:tavastian)
murre-HAAM noisy
55.63
70.08
central and north pohjanmaa (a:ostrobothnian)
murre-POH noisy
55.09
69.68
savo (a:savonian)
murre-SAV noisy
54.70
69.42
southeastern finnish (a:south-east)
murre-KAA noisy
51.68
68.71
southwestern finnish (a:south-west)
murre-LVA noisy
49.80
68.54
neva (a:south-west trans)
murre-LOU noisy
42.57
61.72
norwegian
norwegian bokmål (m:written)
UD_Norwegian-Bokmaal ud
88.53
89.55
98.19
98.67
norwegian nynorsk (m:written)
UD_Norwegian-Nynorsk ud
85.06
85.81
97.83
98.32
norwegian nynorsk (m:written, i:old)
UD_Norwegian-NynorskLIA ud
73.25
79.29
95.47
95.72
saami
north saami
UD_North_Sami-Giella ud
35.92
32.13
78.89
71.50
skolt saami
UD_Skolt_Sami-Giellagas ud
20.26
34.15
sabellic
umbrian
UD_Umbrian-IKUVINA ud
11.90
5.44
sinitic
classical-middle-modern sinitic (a:hongkong, o:traditional)
UD_Chinese-HK ud
68.99
35.49
classical-middle-modern sinitic (o:simplified)
UD_Chinese-GSDSimp ud
58.26
30.92
94.72
95.02
classical chinese
UD_Classical_Chinese-Kyoto ud
35.80
20.85
89.62
89.49
southwestern shifted romance
portuguese (a:european)
UD_Portuguese-PUD ud
80.08
81.38
brazilian portuguese
UD_Portuguese-GSD ud
78.63
80.12
98.19
98.50
portuguese (i:mix)
UD_Portuguese-Bosque ud
78.48
79.85
97.71
97.81
occitan
ROci noisy
76.84
65.80
portuguese (m:written)
UD_Portuguese-CINTIL ud
76.19
78.76
97.67
97.87
tupi-guarani subgroup i.a
mbyá guaraní (a:paraguay)
UD_Mbya_Guarani-Thomas ud
27.89
28.77
old guarani
UD_Guarani-OldTuDeT ud
8.96
10.30
mbyá guaraní (a:brazil)
UD_Mbya_Guarani-Dooley ud
1.94
0.59
west low german
west low german
UD_Low_Saxon-LSDC ud
69.65
54.93
Table 11: Parts-of-speech evaluation report comprising zeroshot score and in-language finetuning. We report F1 as evaluation score. Zeroshot scores are evaluated using model finetuned on Standerd English. If training data is not available, we skip those languages (mentioned as ‘-’) for in-language finetuning.
Table 12: Named entity recognition (NER) evaluation report comprising zeroshot score and in-group finetuning. We report F1 as evaluation score. Zeroshot scores are evaluated using model finetuned on Standerd English. If training data is not available, we skip those languages (mentioned as ‘-’) for in-language finetuning.
cluster variety
target-code
source
dataset
support
Zeroshot (mBERT)
Zeroshot (XLM-R)
Finetune (mBERT)
Finetune (XLM-R)
anglic
english (o:controlled)
simple
en
wikiann
1000
89.07
86.03
89.07
86.03
english
en
en
wikiann
10000
84.15
82.11
84.15
82.11
old english (ca. 450-1100)
ang
en
wikiann
100
54.41
55.94
54.41
55.94
jamaican creole english
jam
en
wikiann
0
0.00
0.00
0.00
0.00
arabic
egyptian arabic
arz
ar
wikiann
100
43.82
50.58
67.77
73.33
standard arabic
ar
ar
wikiann
10000
41.12
41.76
89.10
87.85
circassian
adyghe
ady
en
wikiann
693
67.33
54.03
kabardian
kbd
en
wikiann
1482
47.51
34.79
common turkic
turkish
tr
tr
wikiann
10000
73.56
75.71
92.90
91.80
north azerbaijani
az
az
wikiann
1000
67.31
61.01
89.17
88.26
crimean tatar
crh
tr
wikiann
100
47.81
40.57
57.99
52.67
south azerbaijani
azb
az
wikiann
2567
31.67
11.16
30.22
22.14
eastern romance
romanian
ro
ro
wikiann
10000
74.62
70.82
94.17
93.47
aromanian
roa-rup
el
wikiann
732
64.78
62.66
68.92
63.33
moldavian
mo
ro
wikiann
345
63.31
56.53
60.60
64.56
frisian
western frisian
fy
nl
wikiann
1000
80.17
71.96
81.06
78.21
ems-weser frisian
stq
nl
wikiann
1085
59.45
57.47
64.55
55.15
northern frisian
frr
nl
wikiann
100
46.67
46.22
54.78
52.63
gallo-italian
piemontese
pms
it
wikiann
100
79.53
71.10
87.88
78.72
lombard
lmo
it
wikiann
100
72.49
68.17
79.37
78.66
venetian
vec
it
wikiann
100
62.71
55.20
73.73
67.51
ligurian
lij
it
wikiann
100
45.36
34.75
56.51
47.79
emiliano-romagnolo
eml
it
wikiann
100
33.55
33.23
42.80
45.45
gallo-rhaetian
french
fr
fr
wikiann
10000
79.16
76.52
90.96
89.00
anglo-norman
nrm
fr
wikiann
1281
66.78
88.56
71.98
71.92
arpitan
frp
it
wikiann
2358
63.30
63.67
68.38
71.13
romansh
rm
it
wikiann
100
56.88
55.19
69.69
67.58
friulian
fur
it
wikiann
100
51.41
50.75
64.12
56.30
walloon
wa
fr
wikiann
100
46.33
41.27
45.19
42.50
greater panjabic
western panjabi
pnb
pa
wikiann
100
64.46
53.78
17.82
0.00
eastern panjabi
pa
pa
wikiann
100
32.90
45.25
30.63
0.00
greek
modern greek
el
el
wikiann
10000
71.76
72.96
91.18
90.68
pontic
pnt
el
wikiann
291
66.37
69.79
68.45
72.58
high german
german
de
de
wikiann
10000
79.08
75.67
89.97
87.80
central alemannic
als
de
wikiann
100
75.36
65.15
84.84
79.02
luxemburgish
lb
nl
wikiann
1000
71.61
49.22
79.22
58.71
limburgan
li
nl
wikiann
100
63.03
63.72
78.03
73.15
bavarian
bar
de
wikiann
100
56.62
55.96
76.36
68.84
kölsch
ksh
nl
wikiann
100
54.80
39.42
62.50
48.51
pfaelzisch-lothringisch
pfl
de
wikiann
1092
49.40
47.19
59.12
50.14
pennsylvania german
pdc
de
wikiann
100
41.76
41.79
39.39
38.76
hindustani
fiji hindi
hif
hi
wikiann
715
81.29
79.67
74.28
85.15
hindi
hi
hi
wikiann
1000
65.74
65.77
88.11
85.83
inuit
alaskan inupiaq
ik
en
wikiann
431
76.27
70.56
kalaallisut
kl
en
wikiann
1403
63.20
60.20
italian romance
italian
it
it
wikiann
10000
81.44
78.35
92.21
90.22
italian romance (a:apulia, m:spoken, i:tarantino)
roa-tara
it
wikiann
3811
62.97
66.02
60.34
64.45
sicilian
scn
it
wikiann
100
60.26
54.46
72.25
60.87
continental southern italian
nap
it
wikiann
100
57.35
54.81
61.48
58.33
komi
komi
kv
en
wikiann
2464
56.78
41.78
komi-permyak
koi
en
wikiann
2798
53.62
47.80
kurdish
kurdish
ku
ku
wikiann
100
31.67
59.48
13.70
0.00
central kurdish
ckb
ku
wikiann
1000
6.90
35.49
1.17
0.00
latvian
latvian
lv
lv
wikiann
10000
69.36
71.07
93.24
92.30
east latvian
ltg
lv
wikiann
1036
48.27
48.86
50.10
49.95
mari
eastern mari
mhr
mhr
wikiann
100
46.67
40.94
32.70
0.00
western mari
mrj
mhr
wikiann
6036
38.29
58.46
5.51
0.00
modern dutch
dutch
nl
nl
wikiann
10000
82.03
80.42
91.96
90.51
western flemish
vls
nl
wikiann
100
73.36
72.08
77.66
79.85
zeeuws
zea
nl
wikiann
100
65.98
66.20
79.55
76.09
norwegian
norwegian nynorsk (m:written)
nynorsk
nynorsk
norwegian_ner
1511
87.58
87.86
87.58
87.86
norwegian (m:written, i:samnorsk)
samnorsk
samnorsk
norwegian_ner
3450
86.96
90.55
86.96
90.55
norwegian bokmål (m:written)
bokmaal
bokmaal
norwegian_ner
1939
85.82
90.54
85.82
90.54
sardo-corsican
sardinian
sc
it
wikiann
917
67.32
65.26
80.87
81.17
corsican
co
it
wikiann
100
56.65
56.41
70.59
66.06
serbian-croatian-bosnian
croatian standard
hr
hr
wikiann
10000
77.59
78.12
92.40
91.02
bosnian standard
bs
hr
wikiann
1000
69.93
74.88
87.50
88.86
serbian standard
sr
hr
wikiann
10000
64.38
60.71
63.68
65.86
serbian-croatian-bosnian
sh
hr
wikiann
10000
38.92
69.14
85.05
85.43
sinitic
wu chinese
wuu
zh
wikiann
100
71.89
35.80
74.15
63.73
min nan chinese
zh-min-nan
zh
wikiann
100
44.68
47.30
21.40
15.08
cantonese
zh-yue
zh
wikiann
10000
43.73
26.55
76.19
71.97
mandarin chinese
zh
zh
wikiann
10000
42.86
24.71
81.12
77.22
classical chinese
zh-classical
zh
wikiann
100
28.03
16.78
67.28
62.39
hakka chinese
hak
zh
wikiann
100
27.43
31.84
36.55
40.00
sorbian
lower sorbian
dsb
hsb
wikiann
862
71.04
68.80
3.03
0.00
upper sorbian
hsb
hsb
wikiann
100
65.44
65.48
38.20
0.00
sotho-tswana (s.30)
southern sotho
st
en
wikiann
339
64.36
69.26
northern sotho
nso
en
wikiann
720
19.08
29.66
southwestern shifted romance
galician
gl
es
wikiann
10000
81.98
80.27
87.99
86.14
spanish
es
es
wikiann
10000
72.80
70.73
92.17
90.32
occitan
oc
it
wikiann
100
72.00
67.58
78.50
75.35
mirandese
mwl
es
wikiann
100
46.20
44.07
49.29
42.81
extremaduran
ext
es
wikiann
100
44.83
38.33
61.82
40.00
west low german
west low german
nds
de
wikiann
100
80.29
66.44
79.40
70.99
cluster variety target-code src mBERT (acc) XLM-R (acc) mBERT (F1) XLM-R (F1)
anglic
english
eng_Latn eng_Latn 81.90 83.35 81.95 83.43
arabic
standard arabic
arb_Arab eng_Latn 65.57 73.83 65.57 73.85
najdi arabic
ars_Arab eng_Latn 59.42 69.02 59.14 68.94
ta’izzi-adeni arabic
acq_Arab eng_Latn 58.84 68.72 58.64 68.62
moroccan arabic
ary_Arab eng_Latn 54.65 58.30 54.61 58.14
egyptian arabic
arz_Arab eng_Latn 54.53 65.87 53.86 65.70
south levantine arabic
ajp_Arab eng_Latn 54.09 64.13 53.42 63.81
north mesopotamian arabic
acm_Arab eng_Latn 52.95 58.94 52.84 58.75
levantine arabic (a:north)
apc_Arab eng_Latn 52.14 61.74 51.40 61.31
tunisian arabic
aeb_Arab eng_Latn 47.54 50.18 47.42 50.20
common turkic
north azerbaijani
azj_Latn eng_Latn 59.76 73.17 59.20 73.17
central oghuz (m:spoken)
tur_Latn eng_Latn 59.14 74.47 58.37 74.52
south azerbaijani
azb_Arab eng_Latn 46.05 41.82 44.58 39.24
gallo-italian
venetian
vec_Latn eng_Latn 65.15 68.52 64.99 68.55
lombard
lmo_Latn eng_Latn 59.38 56.39 59.34 56.16
ligurian
lij_Latn eng_Latn 57.82 57.70 56.70 57.16
high german
luxemburgish
ltz_Latn eng_Latn 60.34 47.33 60.01 46.21
limburgan
lim_Latn eng_Latn 50.80 59.88 50.31 59.75
italian romance
italian
ita_Latn eng_Latn 73.71 78.08 73.71 78.19
sicilian
scn_Latn eng_Latn 62.69 56.17 62.66 55.82
kurdish
central kurdish
ckb_Arab eng_Latn 40.98 44.93 37.40 39.59
northern kurdish
kmr_Latn eng_Latn 39.10 63.45 33.93 63.26
latvian
latvian
lvs_Latn eng_Latn 60.14 73.57 59.95 73.63
east latvian
ltg_Latn eng_Latn 48.62 54.19 47.02 53.54
norwegian
norwegian bokmål (m:written)
nob_Latn eng_Latn 72.44 79.44 72.45 79.51
norwegian nynorsk (m:written)
nno_Latn eng_Latn 68.08 71.16 68.10 71.06
sinitic
classical-middle-modern sinitic (o:simplified)
zho_Hans eng_Latn 68.52 72.61 68.54 72.57
classical-middle-modern sinitic (o:traditional)
zho_Hant eng_Latn 61.72 64.89 61.48 64.49
cantonese
yue_Hant eng_Latn 60.44 68.02 60.27 67.41
sotho-tswana (s.30)
northern sotho
nso_Latn eng_Latn 39.38 40.18 35.06 35.98
southern sotho
sot_Latn eng_Latn 39.36 39.30 34.62 34.16
southwestern shifted romance
spanish
spa_Latn eng_Latn 75.13 79.00 75.15 79.09
portuguese (a:european)
por_Latn eng_Latn 73.73 79.18 73.73 79.22
galician
glg_Latn eng_Latn 73.39 78.48 73.39 78.55
occitan
oci_Latn eng_Latn 68.48 63.09 68.47 62.96
Table 13: Natural language inference (NLI) evaluation report using zeroshot cross-lingual transfer from Standard English. We report F1 as evaluation score. NLI uses F1 as evaluation matric. We prepare a translate-train dataset to perform this evaluation.
cluster variety target-code count Finetune (mBERT) Finetune (XLM-R) Zeroshot (mBERT) Zeroshot (XLM-R)
anglic irish english english–irl 494 73.00 67.98 68.62 62.45
southeast american english english–usa 494 73.51 67.95 68.56 62.97
new zealand english english–nzl 494 73.62 68.50 68.21 63.05
southern african english english–zaf 494 73.50 68.06 68.13 63.19
english (a:scotland) english–gbr 494 73.30 67.57 67.90 62.49
australian english english–aus 494 73.23 68.04 67.71 62.33
nigerian english english–nga 494 72.53 66.56 67.39 62.37
philippine english english–phl 494 73.75 67.29 67.35 61.54
indian english (a:south) english–ind_s 494 70.96 66.04 65.43 61.72
kenyan english english–kenya 494 70.63 65.64 65.30 60.28
indian english (a:north) english–ind_n 494 70.28 65.32 64.46 60.84
arabic moroccan arabic arabic–mar 324 70.94 65.14 50.94 49.56
tunisian arabic arabic–tun 324 71.36 65.31 50.86 50.01
arabic (a:jordan) arabic–jor 324 70.60 65.77 50.81 49.56
arabic (a:bahrain) arabic–bhr 324 70.96 65.86 49.88 50.36
arabic (a:saudi-arabia) arabic–sau 324 70.96 65.58 49.29 49.63
egyptian arabic arabic–egy 324 70.06 65.34 48.72 48.71
algerian arabic arabic–dza 324 68.63 64.71 48.71 48.65
bengali bengali (a:west bengal) bengali–ind 107 67.64 70.89 28.06 38.24
bengali (a:dhaka) bengali–dhaka 107 68.98 66.84 27.17 37.32
korean seoul (m:spoken) korean–korn 60 9.60 28.08 6.89 20.24
korean (a:south-eastern, m:spoken) korean–kors 60 9.27 26.96 6.89 19.66
swahili swahili (a:kenya) swahili–kenya 1000 72.06 71.90 46.20 46.22
swahili (a:tanzania) swahili–tanzania 1000 71.08 70.08 45.11 46.00
Table 14: Extractive dialectal question answering evaluation on SD-QA development set. We report span F1 as evaluation score. Zeroshot scores are evaluated using model finetuned on Standerd English whereas, we use combined training set for supervised finetuning
cluster variety target-code count
Finetune (mBERT)
Finetune (XLM-R)
Zeroshot (mBERT)
Zeroshot (XLM-R)
ICL Mistral
anglic
english (a:scotland)
english–gbr 440
76.38
70.34
71.82
63.15
70.18
southern african english
english–zaf 440
76.66
71.18
71.49
63.87
71.14
new zealand english
english–nzl 440
76.71
71.39
71.22
63.69
70.95
australian english
english–aus 440
75.66
70.89
71.20
62.28
69.23
southeast american english
english–usa 440
77.26
71.50
71.17
63.71
71.76
irish english
english–irl 440
75.52
70.73
70.92
62.15
70.64
philippine english
english–phl 440
76.37
70.64
70.47
62.22
70.11
nigerian english
english–nga 440
73.61
68.33
69.10
61.27
68.10
indian english (a:north)
english–ind_n 440
74.62
68.03
68.84
61.25
68.99
kenyan english
english–kenya 440
72.59
66.68
68.72
58.64
64.91
indian english (a:south)
english–ind_s 440
71.93
66.88
66.49
60.36
65.13
arabic
arabic (a:bahrain)
arabic–bhr 921
77.52
72.11
53.25
53.28
55.33
arabic (a:jordan)
arabic–jor 921
77.35
71.29
52.72
53.72
54.93
arabic (a:saudi-arabia)
arabic–sau 921
77.88
72.11
52.72
53.24
55.66
algerian arabic
arabic–dza 921
77.85
72.34
52.56
53.52
55.18
tunisian arabic
arabic–tun 921
76.72
71.64
52.28
52.94
55.03
moroccan arabic
arabic–mar 921
76.73
71.57
51.86
52.17
53.92
egyptian arabic
arabic–egy 921
76.53
70.75
51.80
51.99
54.66
bengali
bengali (a:west bengal)
bengali–ind 113
68.62
73.27
32.30
36.39
50.06
bengali (a:dhaka)
bengali–dhaka 113
67.37
74.24
31.79
35.52
51.73
korean
seoul (m:spoken)
korean–korn 276
10.15
31.91
7.26
19.62
65.79
korean (a:south-eastern, m:spoken)
korean–kors 276
9.92
31.01
7.22
20.08
64.63
swahili
swahili (a:tanzania)
swahili–tanzania 472
63.54
62.30
38.24
39.38
55.70
swahili (a:kenya)
swahili–kenya 472
72.25
70.53
37.97
41.59
58.71
Table 15: Extractive dialectal question answering evaluation on SD-QA test set. We report span F1 as evaluation score. Zeroshot scores are evaluated using model finetuned on Standerd English whereas, we use combined training set for supervised finetuning
cluster variety
dialect-code
support
precision (mBERT)
precision (XLM-R)
recall (mBERT)
recall (XLM-R)
F1 (mBERT)
F1 (XLM-R)
anglic
english (a:uk)
english:en-gb
249
98.10
89.57
83.13
71.59
90.00
79.58
north american english
english:en-us
349
93.27
88.14
83.38
82.09
88.05
85.01
arabic
aleppo
arabic:ale
200
59.50
58.50
59.50
48.35
59.50
52.94
algerian arabic
arabic:alg
272
79.00
65.50
58.09
62.68
66.95
64.06
arabian peninsula arabic (a:yemen)
arabic:san
177
60.50
55.50
68.36
56.63
64.19
56.06
egyptian arabic (a:alx)
arabic:alx
192
70.50
74.50
73.44
66.82
71.94
70.45
egyptian arabic (a:asw)
arabic:asw
221
56.00
52.00
50.68
45.02
53.21
48.26
egyptian arabic (a:cai)
arabic:cai
130
35.50
36.50
54.62
72.28
43.03
48.50
egyptian arabic (a:kha)
arabic:kha
244
63.50
55.50
52.05
44.05
57.21
49.12
fez. meknes
arabic:fes
196
60.00
59.50
61.22
56.40
60.61
57.91
gilit mesopotamian arabic
arabic:bag
203
57.50
47.50
56.65
49.48
57.07
48.47
gulf arabic (a:doh)
arabic:doh
205
50.00
43.50
48.78
45.55
49.38
44.50
gulf arabic (a:jed)
arabic:jed
196
58.00
46.00
59.18
40.89
58.59
43.29
gulf arabic (a:mus)
arabic:mus
178
38.50
49.50
43.26
42.67
40.74
45.83
gulf arabic (a:riy)
arabic:riy
311
62.00
56.50
39.87
37.92
48.53
45.38
levantine arabic (a:north-dam)
arabic:dam
148
37.50
24.50
50.68
42.98
43.10
31.21
libyan arabic (a:ben)
arabic:ben
238
56.50
47.50
47.48
52.78
51.60
50.00
north mesopotamian arabic (a:bas)
arabic:bas
186
49.50
39.00
53.23
49.68
51.30
43.70
north mesopotamian arabic (a:mos)
arabic:mos
188
71.50
70.00
76.06
69.31
73.71
69.65
rabat-casablanca arabic
arabic:rab
153
50.00
40.00
65.36
60.61
56.66
48.19
sfax
arabic:sfx
215
62.50
64.50
58.14
48.13
60.24
55.13
south levantine arabic (a:south-amm)
arabic:amm
177
40.50
35.00
45.76
35.53
42.97
35.26
south levantine arabic (a:south-jer)
arabic:jer
202
48.50
47.00
48.02
40.34
48.26
43.42
south levantine arabic (a:south-sal)
arabic:sal
167
46.00
66.50
55.09
59.11
50.14
62.59
standard arabic
arabic:msa
244
75.00
98.00
61.48
95.61
67.57
96.79
sunni beiruti arabic
arabic:bei
192
58.00
52.50
60.42
68.18
59.18
59.32
tripolitanian arabic
arabic:tri
201
66.00
60.00
65.67
60.30
65.84
60.15
tunisian arabic (a:tun)
arabic:tun
164
52.50
37.00
64.02
56.49
57.69
44.71
greek
cypriot greek (r:casual, m:written, i:other)
greek:cg_other
81
74.14
84.48
53.09
56.32
61.87
67.59
cypriot greek (r:casual, m:written, i:twitter)
greek:cg_twitter
36
51.11
44.44
63.89
68.97
56.79
54.05
modern greek (r:casual, m:written, i:twitter)
greek:smg_twitter
94
89.83
100.00
56.38
53.15
69.28
69.41
high german
central alemannic (a:be)
swiss-dialects:be
389
72.89
51.58
71.21
62.42
72.04
56.48
central alemannic (a:bs)
swiss-dialects:bs
340
73.50
57.83
75.88
61.14
74.67
59.44
central alemannic (a:lu)
swiss-dialects:lu
335
72.91
65.13
75.52
59.47
74.19
62.17
central alemannic (a:zh)
swiss-dialects:zh
359
78.84
73.33
75.77
63.73
77.27
68.19
sinitic
mandarin chinese (a:mainland, o:simplified)
mandarin_simplified:m
986
97.90
96.10
99.29
90.66
98.59
93.30
mandarin chinese (a:mainland, o:traditional, i:synthetic)
mandarin_traditional:m
977
96.80
92.10
99.08
95.74
97.93
93.88
mandarin chinese (a:taiwan, o:simplified)
mandarin_simplified:t
1014
99.30
90.10
97.93
95.85
98.61
92.89
mandarin chinese (a:taiwan, o:traditional, i:synthetic)
mandarin_traditional:t
1023
99.10
95.90
96.87
92.39
97.97
94.11
southwestern shifted romance
brazilian portuguese
portuguese:pt-br
627
96.94
92.35
90.91
84.98
93.83
88.51
latin american spanish
spanish:es-ar
207
81.06
9.25
88.89
91.30
84.79
16.80
portuguese (a:european)
portuguese:pt-pt
349
91.45
83.64
70.49
63.92
79.61
72.46
portuguese (m:written)
portuguese:pt
15
9.70
0.00
86.67
0.00
17.45
0.00
spanish
spanish:es
290
74.21
74.53
81.38
47.69
77.63
58.16
spanish (a:europe)
spanish:es-es
492
90.99
83.33
82.11
78.89
86.32
81.05
Table 16: Dialect Identification evaluation using language cluster specific datasets. We finetune a classificaiton model using either mBERT or XLM-R and then evaluate on the test data.
cluster variety target-code src mBERT (acc) XLM-R (acc) mBERT (F1) XLM-R (F1)
arabic
standard arabic
arb_Arab arb_Arab 85.25 83.96 86.71 82.27
ta’izzi-adeni arabic
acq_Arab arb_Arab 84.96 82.05 86.44 81.98
najdi arabic
ars_Arab arb_Arab 84.80 84.39 87.41 83.33
north mesopotamian arabic
acm_Arab arb_Arab 82.97 80.95 84.77 80.36
south levantine arabic
ajp_Arab arb_Arab 81.82 80.16 84.16 79.05
levantine arabic (a:north)
apc_Arab arb_Arab 81.59 80.15 83.76 79.88
egyptian arabic
arz_Arab arb_Arab 81.02 76.38 84.43 81.03
tunisian arabic
aeb_Arab arb_Arab 79.45 72.88 83.97 77.33
moroccan arabic
ary_Arab arb_Arab 73.87 79.14 78.76 78.55
common turkic
north azerbaijani
azj_Latn azj_Latn 80.46 79.87 82.00 79.55
central oghuz (m:spoken)
tur_Latn azj_Latn 79.10 84.41 80.61 79.51
south azerbaijani
azb_Arab azj_Latn 65.90 67.08 69.71 68.37
gallo-italian
venetian
vec_Latn ita_Latn 76.72 70.68 75.07 74.28
lombard
lmo_Latn ita_Latn 69.92 59.90 70.65 64.56
ligurian
lij_Latn lij_Latn 66.81 63.42 74.03 57.78
high german
luxemburgish
ltz_Latn nld_Latn 74.74 58.50 77.86 64.83
limburgan
lim_Latn nld_Latn 71.09 65.83 71.12 65.73
italian romance
italian
ita_Latn ita_Latn 87.67 84.92 86.68 85.83
sicilian
scn_Latn ita_Latn 75.22 59.71 72.70 59.47
kurdish
northern kurdish
kmr_Latn ckb_Arab 33.23 68.21 10.45 5.71
central kurdish
ckb_Arab ckb_Arab 13.10 19.37 16.86 12.38
latvian
latvian
lvs_Latn lvs_Latn 76.35 83.75 80.63 82.80
east latvian
ltg_Latn lvs_Latn 55.67 65.02 63.69 67.42
norwegian
norwegian nynorsk (m:written)
nno_Latn nob_Latn 85.66 79.94 89.20 79.06
norwegian bokmål (m:written)
nob_Latn nob_Latn 83.81 82.90 83.82 84.14
sinitic
classical-middle-modern sinitic (o:traditional)
zho_Hant zho_Hans 89.82 86.80 89.02 86.39
cantonese
yue_Hant zho_Hans 89.45 86.46 88.71 87.64
classical-middle-modern sinitic (o:simplified)
zho_Hans zho_Hans 88.74 86.38 88.86 89.15
sotho-tswana (s.30)
northern sotho
nso_Latn nso_Latn 35.62 28.16 34.86 13.55
southern sotho
sot_Latn nso_Latn 32.55 32.31 39.93 19.08
southwestern shifted romance
portuguese (a:european)
por_Latn spa_Latn 88.13 89.10 88.10 87.74
galician
glg_Latn spa_Latn 86.99 89.00 86.93 87.83
spanish
spa_Latn spa_Latn 86.74 85.93 84.87 86.55
occitan
oci_Latn lij_Latn 84.12 74.80 78.53 62.56
Table 17: Topic classification evaluation using SIB-200 language data with dialectal presence. We report span F1 as evaluation score. Zeroshot scores are evaluated using model finetuned on Standerd English whereas, we use in-group training for supervised finetuning
dialect MBERT_Acc MBERT_F1 XLMR_Acc XLMR_F1 mistral7b
lang-group variety
arabic tunisian arabic aeb_arab 94.55 94.56 94.61 94.62 73.3
algerian arabic arq_arab 84.98 85.00 84.70 84.69 76.0
arabic (a:jordan) jor_arab 82.96 82.90 89.07 89.00 82.2
arabic (a:saudi-arabia) sau_arab 81.38 65.97 83.40 67.66 79.8
tunisian arabic (r:casual) aeb_latn 80.95 65.65 79.80 63.70 62.3
standard arabic arb_arab 80.63 70.01 83.96 72.91 65.7
moroccan arabic ary_arab 78.08 61.50 77.41 55.55 58.4
egyptian arabic arz_arab 67.03 40.00 69.03 47.89 50.0
south levantine arabic ar_lb 58.38 34.63 58.90 32.72 0.0
Table 18: Sentiment Analysis results. In addition to, using mBERT and XLM-R as the base models, we also perform in-context learning to evaluate the performance of large language models (Mistral-7B).
dialect acc (mBERT) acc (XLM-R) mBERT (F1) XLM-R (F1)
lang-group language
anglic english eng_Latn 52.22 53.56 51.97 53.44
arabic standard arabic arb_Arab 39.00 43.78 39.01 43.78
levantine arabic (a:north) apc_Arab 38.89 40.78 38.64 40.71
north mesopotamian arabic acm_Arab 38.11 41.33 37.99 41.35
moroccan arabic ary_Arab 37.00 37.67 36.94 37.61
egyptian arabic arz_Arab 36.22 38.00 36.21 37.98
najdi arabic ars_Arab 36.00 38.11 36.05 38.16
sinitic classical-middle-modern sinitic (o:simplified) zho_Hans 50.11 47.22 49.79 47.10
classical-middle-modern sinitic (o:traditional) zho_Hant 47.00 45.11 46.88 44.76
sotho-tswana (s.30) northern sotho nso_Latn 31.11 29.78 31.18 29.72
southern sotho sot_Latn 28.56 29.11 28.52 29.00
Table 19: Multiple-choice machine reading comprehension evaluation using Belebele dataset languages with dialectal presence. We report span F1 as evaluation score. We use combined finetuning using the aggregated training data provided with Belebele evaluationd data.
Table 20: Machine Translation
language_group variety NLLB_600m-bleu NLLB_1_3bn-bleu
arabic gulf arabic (a:riy) 43.07 43.07
gulf arabic (a:mus) 40.32 40.32
egyptian arabic (a:alx) 38.90 38.90
egyptian arabic (a:cai) 38.71 38.71
egyptian arabic (a:kha) 38.10 38.10
south levantine arabic (a:south-sal) 37.79 37.79
fez. meknes 37.19 37.19
north mesopotamian arabic (a:mos) 36.47 36.47
arabian peninsula arabic (a:yemen) 35.93 35.93
gilit mesopotamian arabic 35.78 35.78
egyptian arabic (a:asw) 35.02 35.02
south levantine arabic (a:south-amm) 34.88 34.88
north mesopotamian arabic (a:bas) 34.41 34.41
gulf arabic (a:doh) 34.20 34.20
gulf arabic (a:jed) 34.10 34.10
levantine arabic (a:north-dam) 33.36 33.36
south levantine arabic (a:south-jer) 33.27 33.27
libyan arabic (a:ben) 32.45 32.45
aleppo 31.64 31.64
tripolitanian arabic 31.21 31.21
rabat-casablanca arabic 31.17 31.17
sunni beiruti arabic 31.07 31.07
algerian arabic 27.25 27.25
tunisian arabic (a:tun) 25.10 25.10
sfax 23.24 23.24
basque basque (a:azkaine) 25.61 22.58
navarro-labourdin basque (a:helete) 22.53 22.03
basque (a:garruze) 25.40 22.00
navarro-labourdin basque (a:behorlegi) 24.19 21.35
basque (a:luhuso) 22.58 20.95
basque (a:amenduze) 24.21 20.86
basque (a:senpere) 23.47 20.35
basque (a:sara) 23.04 20.25
basque (a:jutsi) 21.46 19.99
navarro-labourdin basque (a:mugerre) 22.00 19.71
basque (a:baigorri) 20.21 18.80
basque (a:uharte) 21.10 18.57
basque (a:suhuskune) 21.10 18.57
basque (a:donibane) 20.28 18.03
basque (a:larzabale) 20.37 17.88
navarro-labourdin basque (a:bidarrai) 17.94 16.90
navarro-labourdin basque (a:urruna) 18.50 16.25
basque (a:iholdi) 16.68 15.01
basque (a:jatsu) 17.01 14.56
basque (a:barkoxe) 10.93 10.82
souletin (a:maule) 11.55 10.36
bengali vanga (a:jessore) 20.71 21.44
vanga (a:khulna) 18.96 19.73
vanga (a:kushtia) 17.36 19.12
vanga (a:dhaka) 17.18 17.85
vanga (a:barisal) 11.33 12.68
common turkic uzbek 24.68 28.82
turkish 22.24 24.22
bashkir 20.88 23.78
central oghuz 18.82 22.65
kirghiz 18.68 22.53
kazakh 19.56 20.96
kara-kalpak 14.31 18.23
sakha 2.53 2.43
farsic dari 37.02 40.49
greek cretan 25.03 21.34
apulian greek 3.92 3.97
high german central bavarian 9.49 12.01
upper saxon 11.99 8.63
kurdish sorani 35.74 34.88
sine’i 22.56 22.60
norwegian norwegian (a:eastern) 24.00 27.49
norwegian (a:southwestern) 21.88 15.81
norwegian (a:setesdal) 2.24 4.23
southwestern shifted romance occitan 20.73 25.74
tigrinya tigrinya (a:ethiopia) 22.40 25.36
tigrinya (a:eritrea) 18.64 21.02
yoruba yoruba (a:central nigeria) 21.46 18.10
high german central alemannic (a:zh) 43.71 44.06
central alemannic (a:gl) 43.90 43.96
central alemannic (a:tg) 43.94 43.62
central alemannic (a:bs) 42.94 43.48
central alemannic (a:gr) 43.67 43.31
central alemannic (a:ag) 43.62 43.26
central alemannic (a:sh) 43.62 43.01
central alemannic (a:zg) 43.05 42.95
central alemannic (a:ai) 42.27 42.72
central alemannic (a:ar) 43.90 42.67
central alemannic (a:sz) 43.24 42.57
central alemannic (a:vs) 41.69 42.52
central alemannic (a:ur) 44.08 42.41
central alemannic (a:be) 42.51 42.36
central alemannic (a:lu) 42.86 42.22
central alemannic (a:nw) 43.04 41.89
central alemannic (a:bl) 43.11 41.68
central alemannic (a:fr) 42.13 41.05
central alemannic (a:ow) 42.29 40.89
central alemannic (a:sg) 41.11 40.55
central alemannic (a:so) 40.27 39.39
italian romance italian (a:umbria) 40.73 41.62
italian (a:lazio) 39.03 32.23
italian (a:veneto) 29.38 31.27
italian (a:toscana) 28.90 31.24
italian (a:marche) 26.29 30.22
italian (a:trentino-alto adige/südtirol) 23.40 25.56
italian (a:unknown) 22.25 24.54
italian (a:friuli-venezia giulia) 18.62 23.51
italian (a:sicilia) 22.91 23.00
italian (a:calabria) 20.11 20.48
italian (a:liguria) 17.98 20.22
italian (a:lombardia) 16.23 17.32
italian (a:campania) 17.03 17.03
italian (a:molise) 17.73 15.98
italian (a:puglia) 16.50 15.64
italian (a:piemonte) 14.48 15.54
italian (a:basilicata) 13.76 14.36
italian (a:emilia-romagna) 13.86 14.35
italian (a:sardegna) 11.67 12.98
italian (a:abruzzo) 13.64 11.02
Appendix F Highest performing and lowest performing varieties
Varieties with Highest Performance Varieties with Lowest Performance
Task (Dataset)
DEP parsing (UD)
anglic/english*
italian romance/italian (r:formal, m:written, i:essay)*
tupi-guarani subgroup i.a/old guarani*
arabic/north african arabic
albanian/albanian*
southwestern shifted romance/portuguese (a:european)
komi/komi-zyrian (m:written)*
italian romance/continental southern italian
gallo-rhaetian/french*
norwegian/norwegian nynorsk (m:written)*
saami/skolt saami*
komi/komi-zyrian (m:spoken)$\dagger$
norwegian/norwegian bokmål (m:written)*
southwestern shifted romance/portuguese (i:mix)*
tupi-guarani subgroup i.a/mbyá guaraní (a:paraguay)
komi/$\dagger$
italian romance/italian*
southwestern shifted romance/brazilian portuguese*
tupi-guarani subgroup i.a/mbyá guaraní (a:brazil)*
saami/north saami*$\dagger$
EQA (SD-QA-test)
anglic/english (a:scotland)
anglic/irish english
swahili/swahili (a:kenya)*
arabic/algerian $\dagger$
anglic/southern african english
anglic/philippine english
bengali/vanga (a:west bengal)$\dagger$
arabic/tunisian $\dagger$
anglic/new zealand english
anglic/nigerian english
bengali/vanga (a:dhaka)$\dagger$
arabic/moroccan $\dagger$
anglic/australian english
anglic/indian english (a:north)
korean/seoul (m:spoken)$\dagger$
arabic/egyptian arabic*$\dagger$
anglic/southeast american english*
anglic/kenyan english
korean/korean (a:south-eastern, m:spoken)$\dagger$
swahili/swahili (a:tanzania)*
TC (SIB-200)
sinitic/classical-middle-modern sinitic (o:traditional)$\dagger$
italian romance/italian*
latvian/east latvian*
arabic/moroccan $\dagger$
anglic/english*
southwestern shifted romance/galician*
sotho-tswana (s.30)/northern sotho*
high german/limburgan
sinitic/cantonese*$\dagger$
southwestern shifted romance/spanish*
kurdish/northern $\dagger$
gallo-italian/lombard
sinitic/classical-middle-modern sinitic (o:simplified)$\dagger$
norwegian/norwegian nynorsk (m:written)*
sotho-tswana (s.30)/southern sotho*
gallo-italian/ligurian
southwestern shifted romance/portuguese (a:european)
arabic/standard arabic*$\dagger$
kurdish/central $\dagger$
common turkic/south $\dagger$
MRC (Belebele)
anglic/english*
arabic/north mesopotamian $\dagger$
arabic/moroccan $\dagger$
sinitic/classical-middle-modern sinitic (o:simplified)$\dagger$
sinitic/classical-middle-modern sinitic (o:simplified)$\dagger$
arabic/moroccan $\dagger$
arabic/egyptian arabic*$\dagger$
sinitic/classical-middle-modern sinitic (o:traditional)$\dagger$
sinitic/classical-middle-modern sinitic (o:traditional)$\dagger$
arabic/egyptian arabic*$\dagger$
arabic/najdi $\dagger$
arabic/standard arabic*$\dagger$
arabic/standard arabic*$\dagger$
arabic/najdi $\dagger$
sotho-tswana (s.30)/northern sotho*
arabic/levantine arabic (a:north)$\dagger$
arabic/levantine arabic (a:north)$\dagger$
sotho-tswana (s.30)/northern sotho*
sotho-tswana (s.30)/southern sotho*
arabic/north mesopotamian $\dagger$
NER (Wikiann)
anglic/english (o:controlled)*
modern dutch/dutch*
sinitic/classical $\dagger$
mari/western $\dagger$
norwegian/norwegian nynorsk (m:written)*
southwestern shifted romance/galician*
sinitic/hakka chinese*$\dagger$
gallo-italian/emiliano-romagnolo
norwegian/norwegian (m:written, i:samnorsk)
italian romance/italian*
sotho-tswana (s.30)/northern sotho*
greater panjabic/eastern panjabi*$\dagger$
norwegian/norwegian bokmål (m:written)*
hindustani/fiji $\dagger$
kurdish/central $\dagger$
common turkic/south $\dagger$
anglic/english*
frisian/western frisian*$\dagger$
anglic/jamaican creole english
kurdish/kurdish*$\dagger$
NLI (XNLI-translate-test)
anglic/english*
norwegian/norwegian bokmål (m:written)*
common turkic/south $\dagger$
arabic/north mesopotamian $\dagger$
southwestern shifted romance/spanish*
sinitic/classical-middle-modern sinitic (o:simplified)$\dagger$
kurdish/central $\dagger$
arabic/levantine arabic (a:north)$\dagger$
southwestern shifted romance/portuguese (a:european)
southwestern shifted romance/occitan
sotho-tswana (s.30)/northern sotho*
high german/limburgan
italian romance/italian*
norwegian/norwegian nynorsk (m:written)*
sotho-tswana (s.30)/southern sotho*
arabic/tunisian $\dagger$
southwestern shifted romance/galician*
arabic/standard arabic*$\dagger$
kurdish/northern $\dagger$
latvian/east latvian*
POS tagging (UD)
anglic/english*
gallo-rhaetian/french (a:paris)
tupi-guarani subgroup i.a/mbyá guaraní (a:paraguay)
sinitic/classical $\dagger$
norwegian/norwegian bokmål (m:written)*
neva/finnish*
komi/komi-zyrian (m:written)*
arabic/levantine $\dagger$
high german/german*
italian romance/italian*
saami/skolt saami*
italian romance/continental southern italian
norwegian/norwegian nynorsk (m:written)*
neva/estonian*
tupi-guarani subgroup i.a/old guarani*
komi/$\dagger$
gallo-rhaetian/french*
southwestern shifted romance/portuguese (a:european)
tupi-guarani subgroup i.a/mbyá guaraní (a:brazil)*
arabic/north african arabic
writing system (non-MT) Latin (77.2%) Latin (44.2%)
dialect (non-MT) Standard (7%) Standard (4.5%)
Table 21: Varieties with the highest and lowest performance (in terms of raw evaluation score) on various tasks in the zero-shot setting. On the left are the top 10. We find that most of these varieties are high-resource standard varieties, and only a few high-resource non-standard varieties. At right are the bottom 10, which are mainly low-resource, non-standard varieties. We use the following notations for variety type and writing system quantification: * marks standard varieties (for some clusters, here we consider multiple varieties as standard because of the substantial resource presence of both of the varieties; e.g. portuguese/european and portuguese/mix) and $\dagger$ notes mix and non-Latin writing system.
Appendix G Low-resource variety performing better in zero-shot NER
Cluster High-resource  variety Script Low-resource variety Script
Hindustani Hindi Devanagari Fiji hindi Latin
Greater panjabic Eastern panjabi Devanagari Western panjabi Arabic
Sotho-tswana (s.30) Northern sotho Latin Southern sotho Latin
Table 22: NER cases where low-resource varieties perform better in zero-shot. The script plays a significant role here as most of the cases, high-resource scripts such as Latin, and Arabic could uplift the performance of a low-resource variety over a high-resource non-Latin script.
Appendix H Cluster level Result Summaries with Demographic Utility and Standard Deviation report
Table 23: Language clusters with their linguistic utility, demographic utility, variety with minimum linguistic utility and standard deviation. Task: DEP Parsing
clusters linguistic utility demographic utility variety (minimum ling. u.) standard deviation
albanian
63.3
69.1
(gheg albanian, 43.5) 28.0
anglic
75.5
90.7
(african american vernacular english, 52.5) 20.4
arabic
64.4
83.3
(south levantine arabic, 49.9) 21.0
eastern-western armenian
88.2
87.5
(eastern armenian, 87.0) 1.6
gallo-italian
50.2
50.2
(ligurian, 50.2) -
gallo-rhaetian
87.7
93.4
(french (a:paris), 77.5) 8.8
high german
56.6
76.5
(central alemannic (a:zh), 36.8) 28.1
italian romance
75.4
76.5
(continental southern italian, 50.0) 17.3
komi
30.2
30.1
(komi-zyrian (m:written), 27.6) 2.4
norwegian
88.4
(norwegian nynorsk (m:written, i:old), 78.4) 8.6
saami
47.9
67.0
(skolt saami, 28.3) 27.8
sabellic
33.2
(umbrian, 33.2) -
sinitic
65.4
(classical chinese, 46.7) 20.8
southwestern shifted romance
87.2
91.6
(portuguese (a:european), 77.3) 7.9
tupi-guarani subgroup i.a
17.2
22.0
(mbyá guaraní (a:brazil), 9.0) 10.5
west low german
40.8
40.8
(west low german, 40.8) -
Table 24: Language clusters with their linguistic utility, demographic utility, variety with minimum linguistic utility and standard deviation. Task: POS Tagging
clusters linguistic utility demographic utility
variety (minimum ling. u.)
standard deviation
albanian
70.1 74.3
(gheg albanian, 55.8)
20.2
anglic
92.6 97.1
(singlish, 88.0)
6.5
arabic
61.9 81.4
(levantine arabic, 43.8)
18.9
eastern-western armenian
93.8 93.0
(eastern armenian, 92.4)
2.0
gallo-italian
58.9 58.9
(ligurian, 58.9)
gallo-rhaetian
96.4 96.9
(old french (842-ca. 1400), 95.6)
0.7
high german
75.5 88.4
(central alemannic (a:zh), 62.6)
18.2
italian romance
82.6 80.9
(continental southern italian, 57.1)
17.9
komi
41.8 41.6
(komi-zyrian (m:written), 35.1)
6.0
neva
75.4 97.7
(neva (a:south-west trans), 61.7)
13.9
norwegian
97.6 -
(norwegian nynorsk (m:written, i:old), 95.7)
1.6
saami
56.5 78.3
(skolt saami, 34.1)
31.6
sabellic
11.9 -
(umbrian, 11.9)
sinitic
84.5 -
(classical-middle-modern sinitic (a:hongkong, o:traditional), 69.0)
13.7
southwestern shifted romance
90.5 95.7
(occitan, 76.8)
10.5
tupi-guarani subgroup i.a
13.7 18.5
(mbyá guaraní (a:brazil), 1.9)
13.7
west low german
69.7 69.7
(west low german, 69.7)
Table 25: Language clusters with their linguistic utility, demographic utility, variety with minimum linguistic utility and standard deviation. Task: NER
clusters linguistic utility demographic utility variety (minimum ling. u.) standard deviation
anglic
57.3 83.9 (jamaican creole english, 0.0) 40.9
arabic
81.2 86.0 (egyptian arabic, 73.3) 11.1
circassian
57.4 54.1 (kabardian, 47.5) 14.0
common turkic
67.9 85.1 (south azerbaijani, 31.7) 28.8
eastern romance
75.9 79.3 (moldavian, 64.6) 16.0
frisian
66.8 80.4 (northern frisian, 54.8) 13.3
gallo-italian
68.6 68.0 (emiliano-romagnolo, 45.5) 17.3
gallo-rhaetian
71.8 90.3 (walloon, 46.3) 16.5
greater panjabic
54.9 53.2 (eastern panjabi, 45.2) 13.6
greek
81.9 90.1 (pontic, 72.6) 13.2
high german
71.5 88.0 (pennsylvania german, 41.8) 15.9
hindustani
86.6 88.1 (fiji hindi, 85.2) 2.1
inuit
69.7 63.5 (kalaallisut, 63.2) 9.2
italian romance
73.0 88.4 (continental southern italian, 61.5) 13.6
komi
55.2 56.6 (komi-permyak, 53.6) 2.2
kurdish
47.5 54.8 (central kurdish, 35.5) 17.0
latvian
71.7 88.7 (east latvian, 50.1) 30.5
mari
52.6 47.4 (eastern mari, 46.7) 8.3
modern dutch
83.8 91.2 (zeeuws, 79.5) 7.1
norwegian
89.7 - (norwegian nynorsk (m:written), 87.9) 1.5
sardo-corsican
75.9 79.8 (corsican, 70.6) 7.5
serbian-croatian-bosnian
83.1 81.0 (serbian standard, 65.9) 11.9
sinitic
64.3 78.3 (hakka chinese, 40.0) 16.8
sorbian
68.3 67.4 (upper sorbian, 65.5) 3.9
sotho-tswana (s.30)
49.5 41.1 (northern sotho, 29.7) 28.0
southwestern shifted romance
74.0 92.1 (mirandese, 49.3) 18.1
west low german
80.3 80.3 (west low german, 80.3) -
Table 26: Language clusters with their linguistic utility, demographic utility, variety with minimum linguistic utility and standard deviation. Task: NLI
clusters linguistic utility demographic utility
variety (minimum ling. u.)
standard deviation
anglic
83.4 83.4
(english, 83.4)
arabic
63.3 70.0
(tunisian arabic, 50.2)
7.1
common turkic
64.1 62.3
(south azerbaijani, 44.6)
16.9
gallo-italian
61.7 63.6
(ligurian, 57.2)
6.0
gallo-rhaetian
54.6 54.6
(friulian, 54.6)
high german
59.9 59.8
(limburgan, 59.7)
0.2
italian romance
70.4 77.1
(sicilian, 62.7)
11.0
kurdish
51.4 55.5
(central kurdish, 39.6)
16.7
latvian
63.6 71.5
(east latvian, 53.5)
14.2
modern dutch
76.5 76.5
(dutch, 76.5)
norwegian
75.3 -
(norwegian nynorsk (m:written), 71.1)
6.0
sardo-corsican
58.3 58.3
(sardinian, 58.3)
sinitic
68.2 67.4
(classical-middle-modern sinitic (o:traditional), 64.5)
4.1
sotho-tswana (s.30)
35.3 35.6
(southern sotho, 34.6)
1.0
southwestern shifted romance
76.3 79.1
(occitan, 68.5)
5.2
Table 27: Language clusters with their linguistic utility, demographic utility, variety with minimum linguistic utility and standard deviation. Task: TC
clusters linguistic utility demographic utility
variety (minimum ling. u.)
standard deviation
anglic 89.7 89.7
(english, 89.7)
arabic 84.5 85.7
(moroccan arabic, 79.1)
2.4
common turkic 78.7 77.3
(south azerbaijani, 69.7)
7.9
gallo-italian 73.8 73.7
(lombard, 70.6)
3.0
gallo-rhaetian 68.8 68.8
(friulian, 68.8)
high german 74.5 72.7
(limburgan, 71.1)
4.8
italian romance 81.4 86.8
(sicilian, 75.2)
8.8
kurdish 43.8 52.3
(central kurdish, 19.4)
34.5
latvian 75.6 82.0
(east latvian, 67.4)
11.5
modern dutch 89.6 89.6
(dutch, 89.6)
norwegian 86.7 -
(norwegian bokmål (m:written), 84.1)
3.6
sardo-corsican 71.0 71.0
(sardinian, 71.0)
sinitic 89.5 89.5
(classical-middle-modern sinitic (o:simplified), 89.2)
0.3
sotho-tswana (s.30) 37.8 36.9
(northern sotho, 35.6)
3.1
southwestern shifted romance 87.2 86.9
(occitan, 84.1)
2.3
Table 28: Language clusters with their linguistic utility, demographic utility, variety with minimum linguistic utility and standard deviation. Task: DId
clusters linguistic utility demographic utility
variety (minimum ling. u.)
standard deviation
anglic 89.0 88.4
(north american english, 88.0)
1.4
arabic 58.1 89.0
(south levantine arabic (a:south-amm), 43.0)
11.3
greek 64.6 -
(cypriot greek (r:casual, m:written, i:twitter), 56.8)
6.8
high german 74.5 -
(central alemannic (a:be), 72.0)
2.1
sinitic 98.3 98.3
(mandarin chinese (a:mainland, o:traditional, i:synthetic), 97.9)
0.4
southwestern shifted romance 73.3 82.7
(portuguese (m:written), 17.4)
27.9
Table 29: Language clusters with their linguistic utility, demographic utility, variety with minimum linguistic utility and standard deviation. Task: SA
clusters linguistic utility demographic utility variety (minimum ling. u.) standard deviation
arabic 80.3 81.4 (levantine/south, 58.9) 10.7
Table 30: Language clusters with their linguistic utility, demographic utility, variety with minimum linguistic utility and standard deviation. Task: MRC
clusters linguistic utility demographic utility
variety (minimum ling. u.)
standard deviation
anglic 53.4 53.4
(english, 53.4)
arabic 39.9 42.1
(moroccan arabic, 37.6)
2.4
sinitic 48.3 -
(classical-middle-modern sinitic (o:traditional), 46.9)
2.1
sotho-tswana (s.30) 30.1 30.5
(southern sotho, 29.0)
1.5
Table 31: Language clusters with their linguistic utility, demographic utility, variety with minimum linguistic utility and standard deviation. Task: EQA
clusters linguistic utility demographic utility variety (minimum ling. u.) standard deviation
anglic 75.2 75.4 (indian english (a:south), 71.9) 1.8
arabic 77.2 77.0 (egyptian arabic, 76.5) 0.6
bengali 73.8 - (vanga (a:west bengal), 73.3) 0.7
korean 65.2 64.6 (korean (a:south-eastern, m:spoken), 64.6) 0.8
swahili 67.9 64.1 (swahili (a:tanzania), 63.5) 6.2
Appendix I In-Context Learning Details
I.1 Prompts
We adapt the prompts from Super-NaturalInstructions Wang et al. (2022) for our in-context learning experiments.
Sentiment Analysis.
For sentiment analysis we provide 4 few-shot examples in the prompt. The prompt template is given below:
In this task, you are given a piece of text. Your task is to classify the sentiment of the text based on its content.
Sentence: <Sentence Example 1>
Label: <Label for Example 1, Positive, negative, neutral>
$\cdots$
Sentence: <Sentence Example $k$>
Label: <Label for Example $k$, Positive, negative, neutral>
Sentence: <Test Example Input>
Label:
Extractive Question Answering.
We provide 2 few-shot examples i.e. $k=2$ due to the long-form nature of text for this task.
This task is about writing a correct answer for the reading comprehension task. Based on the information provided in a given passage, you should identify the shortest continuous text span from the passage that serves as an answer to the given question. Avoid answers that are incorrect or provides incomplete justification for the question.
Passage: <Passage for Example 1>
Question: <Question for Example 1>
Answer: <Answer for Example 1>
$\cdots$
Passage: <Passage for Example $k$>
Question: <Question for Example $k$>
Answer: <Answer for Example $k$>
Passage: <Passage for Test Example>
Question: <Question for Test Example>
Answer: