### v2

# DialectBench: A NLP Benchmark for Dialects, Varieties, and Closely-Related Languages

Abstract
Language technologies should be judged on their usefulness in real-world use cases. An often overlooked aspect in natural language processing (NLP) research and evaluation is language variation in the form of non-standard dialects or language varieties (hereafter, varieties). Most NLP benchmarks are limited to standard language varieties. To fill this gap, we propose DialectBench, the first-ever large-scale benchmark for NLP on varieties, which aggregates an extensive set of task-varied variety datasets (10 text-level tasks covering 281 varieties). This allows for a comprehensive evaluation of NLP system performance on different language varieties. We provide substantial evidence of performance disparities between standard and non-standard language varieties, and we also identify language clusters with larger performance divergence across tasks. We believe DialectBench provides a comprehensive view of the current state of NLP for language varieties and one step towards advancing it further.  [^1] More information can be found at the following:
Code/data: https://github.com/ffaisal93/DialectBench Website: https://fahimfaisal.info/DialectBench.io
1 Introduction
Benchmarking is important for tracking the progress the field of natural language processing (NLP) has made in various tasks. In the past few years, large-scale multilingual benchmarks like XTREME Hu et al. (2020), XTREME-R Ruder et al. (2021), and XGLUE Liang et al. (2020) have played a pivotal role in evaluating the multilingual capabilities of NLP models. These efforts have sought to make model evaluation more accessible to researchers and representative of a variety of languages Song et al. (2023). However, most benchmarks have focused on the standard varieties of languages, largely neglecting non-standard dialects and language varieties Blasi et al. (2022).
Figure 1: DialectBench Evaluation Suite.
We refer to non-standard dialects and language varieties simply as varieties, and sometimes include low-resource related languages, writing system variants, and other kinds of variation. Varieties contain subtle but notable variations in vocabulary, pronunciation, orthography and grammar, reflecting regional, social, and cultural differences Chambers and Trudgill (1998). The non-standard nature of these language varieties oftentimes contributes to the scarcity of substantial datasets that accurately capture these variations Hedderich et al. (2021). As a result, they have often been absent from widely adopted benchmarks, even from admirable efforts like XTREME-up Ruder et al. (2023), GLUECoS Khanuja et al. (2020) and CreoleVal  Lent et al. (2023), which focus on under-resourced, code-switched, and creole languages, respectively.
Figure 2: DialectBench language clusters with their variety counts per task. "Other" encompasses 18 clusters (full cluster list in Appendix LABEL:lang_list).
It is currently challenging to accurately test the robustness of multilingual models on a large suite of varieties without establishing an NLP evaluation framework covering multiple language clusters (each of which contains standard languages alongside its closely related varieties).
To this end, we create DialectBench, a large-scale benchmark covering 40 language clusters with 281 varieties, spanning 10 NLP tasks. We observe that the performance disparity between different varieties of the same language cluster becomes more pronounced when we shift from zero-shot evaluation to fine-tuning on variety data, because of uneven data availability across varieties. Certain language clusters exhibit varying performance across downstream tasks within the same category, due to low-resource limitations. Additionally, we improve the dialectal task coverage for natural language inference by constructing a translate-test evaluation dataset. Putting these all together, DialectBench serves as a comprehensive suite that attains a multifaceted purpose: identifying broader limitations in dialectal NLP, while reflecting on potential areas for improvement.
2 DialectBench
Category Task Metric Source Dataset
Structured DEP parsing UAS Universal Dependency Zeman et al. (2021), TwitterAAE Blodgett et al. (2018), Singlish Wang et al. (2017)
Prediction POS tagging F1 Universal Dependency Zeman et al. (2021), Singlish Wang et al. (2017), Noisy Dialects Blaschke et al. (2023)
NER F1 Wikiann Pan et al. (2017); Rahimi et al. (2019), Norwegian NER Johansen (2019)
Classification DId F1 MADAR Bouamor et al. (2018), DMT Jauhiainen et al. (2019), Greek Sababa and Stassopoulou (2018), DSL-TL Zampieri et al. (2023), Swiss Germans Scherrer et al. (2019)
SA F1 TSAC Medhaffar et al. (2017), TUNIZI Fourati et al. (2021), DzSentiA Abdelli et al. (2019), SaudiBank Alqahtani et al. (2022), MAC Garouani and Kharroubi (2022), ASTD Nabil et al. (2015), AJGT Alomari et al. (2017), OCLAR Al Omari et al. (2019)
TC F1 SIB-200 Adelani et al. (2023)
NLI F1 XNLI Conneau et al. (2018) translate-test
Question MRC F1 Belebele Bandarkar et al. (2023)
Answering EQA Span F1 SDQA Faisal et al. (2021)
Generation MT BLEU CODET Alam et al. (2023), TIL-MT Mirzakhalov (2021)
Table 1: The tasks and data sources of DialectBench (Detailed discussion: LABEL:{sec:tasks_benchmark}).
DialectBench is a benchmark created to unify dialectal datasets across different tasks to foster research on language varieties and non-standard dialects. Below we describe the design choices we undertook to achieve this goal. This includes our language variety and task selection procedures, data collation methods, and evaluation principles.
Variety Selection
We first looked through papers published in the ACL Anthology [^2] https://aclanthology.org from the last 10 years to find usable language resources, as well as commonly used online data repositories Littauer (n.d.). We selected languages that have well-established, high-resourced varieties. Varieties may vary by location, ethnicity, or other factors. We also found instances where varieties are classified by writing system or even by genre (e.g., Twitter). When varying by location, varieties may be classified by different datasets at different levels of granularity, sometimes country, region, or city. In some cases, we found resources with two or more varieties within one dataset (e.g., the UD_Portuguese-Bosque dependency treebank Rademaker et al. (2017) includes examples from both European and Brazilian Portuguese variants. To incorporate all these cases under one paradigm, we formulate a cluster-variety mapping procedure.
Cluster-Variety Mapping
We construct several language clusters comprising of both high-resourced varieties and their low-resourced counterparts. We use the Glottolog language database Nordhoff (2012) to define clusters and assign varieties as outlined in Figure 1. This design choice enables us to keep varieties that are closely related in terms of either mutual intelligibility, phylogenetic similarity or geographic proximity within the same cluster. Hence, all cluster varieties always root back to the closest common linguistic ancestor and the whole cluster maps to an established phylogenetic subtree. For example, Fiji Hindi and Hindi, with Hindustani [^3] https://glottolog.org/resource/languoid/id/hind1270 as their closest common ancestor, are placed in the Hindustani cluster.
We primarily use the Glottocode language identification scheme Hammarström and Forkel (2022), ensuring a standardized naming scheme across all varieties. For instance, AAE variety from TwitterAAE Blodgett et al. (2018) dependency parsing dataset is renamed as African American Vernacular English with a corresponding Glottocode afri1276. In cases where Glottocodes are unavailable, like for spoken English from South India, we substitute with the immediate ancestor Glottocode (indi1255) and further categorize the varieties using the following metadata identifiers:
  1. 1.
    Area (a): the region where the variety is spoken or where its dataset was collected.
  2. 2.
    Language register (r): frozen, formal, consultative, casual, and intimate.
  3. 3.
    Language mode (m): written, spoken, and signed language.
  4. 4.
    Orthography (o): In DialectBench this is only specific to Sinitic varieties. This could be either traditional or simplified.
  5. 5.
    Identifier (i): Dataset-specific metadata, could be domain (eg. twitter).
We encapsulate all this information, into a naming convention, and use the template: {Glottocode name}-(a:{},r:{},m:{},o:{},i:{}).  [^4] For example, mandarin chinese (a:mainland, o:simplified refers to Mandarin Chinese (mand1415) spoken in Mainland China and written in simplified characters.
Cluster Representative
Each cluster will often have a high-resourced variety usually with the largest speaker population. We choose this high-resourced variety as the cluster representative. This selection might vary across downstream tasks depending on the data availability. We primarily utilize this representative variety to evaluate the performance gap across cluster varieties, and also rely on it for transfer-learning in resource-scarce settings. Sometimes, the members of a cluster are considered closely related languages, and sometimes dialects; to avoid making this distinction, we refer to all the members of a cluster simply as varieties of the cluster representative.
Task and Dataset Selection
In selecting tasks, we maintain a balanced approach, promoting task diversity while also including tasks that require diverse levels of textual understanding. In the end, our complete list of tasks are as follows:
  1. 1.
    Dependency parsing (DEP parsing)
  2. 2.
    Parts-of-speech tagging (POS tagging)
  3. 3.
    Named entity recognition (NER)
  4. 4.
    Dialect identification (DId)
  5. 5.
    Sentiment analysis (SA)
  6. 6.
    Topic classification (TC)
  7. 7.
    Natural language inference (NLI)
  8. 8.
    Multiple-choice machine reading comprehension (MRC)
  9. 9.
    Extractive question answering (EQA)
  10. 10.
    Machine translation (MT)
In Table 1, we present the task and dataset details. We mostly keep the datasets in their originally published form (except for varieties renaming). For NLI, we use the existing English test set of XNLI Conneau et al. (2018) and construct a multilingual dialect-focused translated evaluation dataset. We refer to this as translate-test NLI.
Evaluation Principles
On the ground level, we evaluate existing NLP systems on text-based tasks using standard evaluation metrics (e.g., UAS for parsing, F1 for classification tasks, BLEU for translation). At a global level, we believe a sustainable NLP system should be user-focused while providing substantial (i) linguistic utility and (ii) demographic utility Song et al. (2023); Blasi et al. (2022). Blasi et al. (2022) defined the utility of a task and language, as the corresponding performance normalized by the best possible performance (usually human-level performance). Demographic utility considers the demand for a language technology within a specific language, where the demand is proportional to the number of speakers of that language. Linguistic utility, on the other hand, asserts that “all languages are created equal” regardless of the number of speakers, and hence all languages in the world should receive identical weights.
Overall, we want to capture the performance gap between language clusters (e.g., Anglic vs. Italian Romance) as well as within language clusters (e.g., Norwegian Bokmål vs. Nynorsk). To attain this, we define the performance gap metrics in §3.3 . We also vary the experimental settings in §3.2, including zero-shot and few-shot cross-lingual transfer, as well as fine-tuning with similar high-resource languages. This is essential given that we lack clean annotated data in many varieties.
3 Experiments
Here, we report the entire process involved in creating and evaluating baselines for all of the tasks and varieties in DialectBench. Additionally, we define a dialectal gap metric to analyze performance disparities within and across clusters.
3.1 Models
We evaluate using two multilingual models: mBERT Devlin et al. (2019) and XLM-R Conneau et al. (2020) for all tasks except MT. For MT, we do zero-shot evaluation with NLLB NLLB Team et al. (2022), using both the 600M and 1.3B variants. In addition, we use Mistral 7B Jiang et al. (2023) to evaluate the current capability of LLMs on multilingual and dialectal understanding tasks. Our main goal is collating dialectal data across different languages and tasks under a single platform, hence we do not optimize for the best model performance. Rather, we focus on understanding and reporting the current state of performance on all DialectBench varieties.
3.2 Training and Evaluation
Training and evaluation procedures are largely determined by the availability of training or evaluation data for each task.
For any cluster $C$, let ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{C}}$ be the highest-resourced variety (which is usually the cluster representative) of $C$. In addition, for any variety $v\in C$, we write ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}}$ for the highest-resourced variety, that is, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{C}}$. For any varieties $t$ and $v$, let ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{S}_{t}(v)}$ be the raw evaluation score of a system fine-tuned on $t$ and tested on $v$ (higher is better).
We use five general approaches for task-specific model training:
  1. 1.
    In-variety fine-tuning (${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{S}_{v}(v)}$): In cases where there is available training data for a variety $v$, we fine-tune the base model on $v$. This primarily applies to tasks such as POS tagging and dependency parsing.
  2. 2.
    In-cluster fine-tuning (${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{S}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}}}(v)}$): In-variety fine-tuning is quite resource-intensive when we have a large number of varieties within a cluster $C$. In such cases, we fine-tune the base model on ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{C}}$. Then we evaluate this model on each variety $v\in C$. This is the most common setting in our experiments, as it allows us to evaluate the dialectal performance gap without increasing the computation cost. For the DId task, we typically use a dataset of sentences annotated with variety labels to fine-tune one dialect-identification model for each language cluster.
  3. 3.
[…]
    Task In-variety FT In-cluster FT Combined FT Zero-shot No ref-erence In-context learning
    DEP $\checkmark$ $\checkmark$
    POS $\checkmark$ $\checkmark$
    NER $\checkmark$ $\checkmark$
    EQA $\checkmark$ $\checkmark$ $\checkmark$
    MRC $\checkmark$
    NLI $\checkmark$
    TC $\checkmark$ $\checkmark$
    SA $\checkmark$ $\checkmark$
    DId $\checkmark$
    MT $\checkmark$ $\checkmark$
    Table 2: Task-specific training and evaluation procedure.
  4. 4.
    Zero-shot evaluation (${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{S}_{\text{eng}}(v)}$): For certain varieties, obtaining training data even for in-cluster fine-tuning can be a challenge. Fortunately, English training data is always available for the datasets we study, so we use English to fill the gaps when we lack in-variety, in-cluster, or combined training data. At the same time, we aim to assess the feasibility of using this zero-shot cross-lingual transfer in reducing any existing performance gap across varieties. So we perform zero-shot cross-lingual transfer from English to each variety for 6 tasks in total. We only leave out those tasks such as dialect identification that explicitly require in-cluster training data.
  5. 5.
    In-context learning (${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{S}_{\text{icl}}(v)}$): When evaluating large language models, we do not fine-tune them but instead rely on prompting and in-context learning. For this, we provide instructions and 5 examples in English as exemplars, followed by a prompt for predicting the test examples. Employing Mistral 7B Jiang et al. (2023), we assess the present effectiveness of a close-to-state-of-the-art LLM on language varieties. The task-specific example prompts are reported in Appendix I.
Table 2 summarizes the task-specific training procedures that we employ based on data availability. Note that, for MT, we perform zero-shot evaluation specifically in the translation direction, (standard variety to English) tested on (dialectal variety to English). But evaluation is a challenge because human-created reference translations into or out of non-standard varieties are usually very limited. Therefore, we adopt an evaluation protocol from previous work (Alam et al., 2023) that uses pseudo-references. Given $x$, a sentence in a variety and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{x}}$, the translation of $x$ into the standard variety, let $y$ be the output of the MT system on input $x$ and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{y}}$ be the output on input ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{x}}$. Then we measure the quality of $y$ (using, e.g., BLEU) compared against ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{y}}$ as a pseudo-reference.
3.3 Quantifying the Dialectal Gap
To quantify the performance disparity across various resource-specific settings, language clusters and varieties, we introduce a dialect performance gap metric ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{t}(u,v)}$: the relative decrease in performance of a system fine-tuned on variety $t$, tested on variety $v$ compared to a baseline variety $u$:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{t}(u,v)}=\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{S}_{t}(u)}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{S}_{t}(v)}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{S}_{t}(u)}}$
with a special case for in-variety fine-tuning:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{in-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}variety}}}(u,v)}=\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{S}_{u}(u)}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{S}_{v}(v)}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{S}_{u}(u)}}.$
For the baseline score, we use either the score on the standard variety for each cluster ($u={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}}$), or, in the zero-shot setting, the score on the language used for fine-tuning, namely English ($u=t=\text{eng}$). Rather than computing an absolute gap, we opt for a relative gap (i.e., dividing by the baseline score). We also indicate whether the training setting is zero-shot ($t=\text{eng}$) or fine-tuned on in-variety, in-cluster, or assembled data. Putting all these together, we compute the following three variations of dialectal inequality.
  1. 1.
    ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}(\text{eng},v)}$: We calculate this metric to get a comprehensive measurement of global disparity across all varieties in a resource-limited environment (zero-shot transfer from English).
  2. 2.
    ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},v)}$: Using this variation, we keep the setting fixed as zero-shot and calculate the gap between the representative variety and any other variety.
  3. 3.
    ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{t}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},v)}$: The two aforementioned metrics shed light on the extent of the variety performance gap in a resource-limited setting. To gain a more comprehensive perspective, we additionally compute another metric, this time utilizing the availability of resources. The computation approach remains as straightforward as before. We just use fine-tuning on a variety $t$ instead of zero-shot transfer from Standard English: $t=\text{in-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}variety}}$ for in-variety fine-tuning, $t={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}}$ for in-cluster fine-tuning, or $t$ is some set of varieties for combined fine-tuning.
For all three ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}}$ metrics, we compute them at the variety level and then average them at the cluster level:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{t}(u,C)}$ $=\frac{1}{|C|}\sum_{v\in C}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{t}(u,v)}$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{in-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}variety}}}(u,C)}$ $=\frac{1}{|C|}\sum_{v\in C}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{in-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}variety}}}(u,v)}.$
4 Results
num. num. avg.
Category Task cl. var. score Max-score cluster/variety Min-score cluster/variety
Structured prediction DEP parsing 16 40 64.3 sw. shifted romance/brazilian portuguese 94.4 tupi-guarani sg./mbyá guaraní (a:brazil) 9.0
POS tagging 17 51 72.1 norwegian/norwegian bokmål (m:written) 98.7 tupi-guarani sg./mbyá guaraní (a:brazil) 1.9
NER 27 85 70.1 eastern romance/romanian 94.2 anglic/jamaican creole english 0.0
Sequence classification NLI 15 38 64.2 anglic/english 83.4 sotho-tswana (s.30)/southern sotho 34.6
TC 15 38 77.7 sinitic/cmm. sinitic (o:traditional) 89.8 kurdish/central kurdish 19.4
DId 6 49 67.0 sinitic/mandarin chinese (a:taiwan, o:simp.) 98.6 sw. shifted romance/portuguese (m:written) 17.4
SA 1 9 80.3 arabic/tunisian arabic 94.6 arabic/south levantine arabic 58.9
Question Answering MRC 4 11 40.9 anglic/english 53.4 sotho-tswana (s.30)/southern sotho 29.0
EQA 5 24 74.2 arabic/arabic (a:saudi-arabia) 77.9 swahili/swahili (a:tanzania) 63.5
Generation MT-dialect 12 73 25.2 arabic/gulf arabic (a:riy) 43.1 common turkic/sakha 2.5
MT-region 2 41 33.0 high german/central alemannic (a:ur) 44.1 italian romance/italian (a:sardegna) 13.0
Table 3: Task specific result summary using Maximum Obtainable Score. The varieties with the minimum scores exhibit a noticeable lag in performance across various tasks when compared to the average task performance.
We, first of all, discuss results by highlighting the highest possible score per variety, aka the maximum obtainable evaluation scores regardless of evaluation method or training data. Next, we extend our discussion further by reporting the existing dialectal disparity across clusters and varieties.
4.1 Maximum Obtainable Scores
Here we provide key findings from our evaluation on each task. A task-specific summary is reported in Table 3. Detailed results comprising all tasks, models, language clusters and varieties are reported in Tables 10 to LABEL:zero-shot-mt in Appendix E.
Figure 3: Maximum scores (max. UAS) in Dep. Parsing task. Yellow-shaded region: Komi is the only cluster having no varieties seen during mBERT pertaining. Colored bars with diagonal stripes: the cluster representative variety. Low-resourced cluster varieties score lower compared to high-resource Germanic clusters.
Structured prediction
We present visualizations for the task-specific maximum scores. We show the one for Dependency Parsing in Fig. 3, where we observe that low-resource varieties from language clusters such as Tupi-Guarani (indigenous South American cluster), Saami and Komi (low-resource Uralic language clusters) have the lowest performance compared to Standard English and other closely related Germanic and Romance clusters. These low-resource varieties are also not included in the pretraining stage of our base language models (eg. mBERT). Furthermore, this trend is evident across all three structured prediction tasks. On the other hand, high-resource Indo-European languages such as Portuguese, French, and Norwegian usually perform better.
Sequence classification
For DId and SA, we generally collate different datasets for each language cluster and therefore, report the comparative classification results together. As a result, the locality level (e.g. city/region/country) also varies across clusters. For example, we report city-level DId results for Arabic and High German but country-level results for Portuguese, Spanish and English. In the case of SA, we have region/country-level results for Arabic varieties. For TC and NLI, we have the same set of clusters and varieties. However, we only report the zero-shot transfer performance from Standard English for NLI using the newly created translate-test NLI dataset.
For TC and NLI, we observe the largest in-cluster disparity in the Kurdish cluster, with Northern Kurdish outperforming all others. The Sotho varieties consistently perform significantly lower compared to other clusters. For all three sequence classification tasks, we generally find the Chinese cluster performing on par with high-resource Latin counterparts.
Question answering
We generally do not see large gaps in performance within varieties in each language cluster. In EQA zero-shot experiments, English and its varieties have the highest performance overall and Korean varieties score the lowest. Combined fine-tuning boosts performance on all language clusters except in English. It’s important to note that these EQA scores primarily indicate the model’s robustness to accent-level differences and transcription noise, rather than broad dialectal robustness. However, further investigation is needed to determine whether this robustness specifically applies to both accent-level differences and transcription noise, or to any character-level variation up to a certain threshold.
For the MRC task, the performance peaked at 53.4 for Standard English, while the lowest score was 29 for Southern Sotho. More detailed results are presented in Tables 19 and 15.
Machine translation
The performance gap here varies widely across and within language varieties. Performance is similar within the Swiss-German cluster, with higher performance (see Figure 5) across regions in Northern Switzerland, which is geographically closer to Germany. The performance gap for Norwegian dialects (Figure 10(b)) is surprising as we perform zero-shot transfer from Norwegian Nynorsk (a Western dialect) but obtain better performance on the Eastern dialect. Within Arabic, Riyadh is the highest performer while Sfax performs the worst. For the Bengali cluster, Jessore has the highest performance –this is not surprising since it is one of the dialects from which standard Bengali originated Alam et al. (2023). The Ethiopian variety of Tigrinya exhibits a higher performance than the Eritrean one, even though Tigrinya is more commonly spoken in Eritrea  [^5] https://en.wikipedia.org/wiki/Tigrinya_language. Amongst the clusters within the Basque cluster, Barkoxe and Maule have the lowest score while Azkaine scores the highest.
4.2 Dialectal Gap Across Language Clusters
(a) Dependency parsing
(b) Topic classification
(c) Extractive question answering
Figure 4: Dialectal gap visualization for language clusters utilizing zero-shot cross-lingual transfer from Standard English. In the x-axis, values far from zero have a larger performance gap from English whereas, in the y-axis, values far from zero have a larger within cluster gap. Ideally, we want both of them to be close to zero.
In Fig. 4, we plot the zero-shot dialectal gap for three tasks. In the $x$-axis, we report the aggregated cluster-level gap ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}(\text{eng},v)}$, compared against the fine-tuning variety (standard English) while in $y$-axis we report ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},v)}$, the gap compared against the representative variety of a cluster. In an ideal scenario, we would want both of these gap values to be close to zero. However, this is certainly not the case. The general observed trend is that the low-resource clusters have higher gaps of both ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}(\text{eng},C)}$ and $\mathcal{G}_{\text{eng}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},C)$, whereas high-resource Germanic and Sinitic language clusters consistently exhibit low dialectal gaps. That said, certain specific high-resource varieties, such as Standard German and its dialectal counterparts like Swiss German, showcase significant within-cluster dialectal gaps (Fig. 4(a)).
Task Gap metric Avg. val cluster (max) cluster (min)
DEP parsing ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}(\text{eng},C)}$ 34.0 arabic, 50.8 sw. shift. romance, 19.4
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},C)}$ 15.7 anglic, 34.2 sw. shift. romance, $-0.9$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{in-variety}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},C)}$ 26.4 arabic, 93.8 italian romance, 0.6
POS tagging ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}(\text{eng},C)}$ 27.4 arabic, 58.5 norwegian, 14.7
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},C)}$ 6.7 anglic, 20.9 ew. armenian, $-2.1$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{in-variety}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},C)}$ 6.2 arabic, 29.1 neva, $-0.5$
NER ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}(\text{eng},C)}$ 31.7 kurdish, 77.1 modern dutch, 12.3
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},C)}$ 22.3 kurdish, 78.2 hindustani, $-23.6$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},C)}$ $-28.6$ kurdish, 91.5 sorbian, -1162.7
TC ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}(\text{eng},C)}$ 22.4 kurdish, 74.2 sinitic, 0.4
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},C)}$ 12.5 kurdish, 60.6 norwegian, $-2.2$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},C)}$ $-1.9$ latvian, 21.0 kurdish, $-61.3$
Table 4: Comparative cluster-level dialectal gap across tasks. In general, the average disparity is larger for zero-shot transfer ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}(\text{eng},C)}$. However, when we move from zeroshot to finetune (i.e. $\mathcal{G}_{\text{eng}}\rightarrow\mathcal{G}_{\bar{v}/\text{in-variety}}$) and compute the distance from a cluster representative ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}}$, we observe increased dialectal disparity $|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},C)}|$.
We primarily report dialectal gaps using zero-shot transfer because the finetuning data available across task and cluster is very disproportionate. Often the in-cluster/variety data is not good enough in terms of data quality and quantity. For example, we have 37 varieties in 13 clusters for dependency parsing but out of these, only 20 varieties have data available for in-variety fine-tuning. This lacking becomes more apparent when we compare the statistics of two types of within-cluster dialectal gaps: zero-shot $\mathcal{G}_{\text{eng}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},C)$ against fine-tuning $\mathcal{G}_{\bar{v}/\text{in-variety}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},C)$ in Table 4. In general, the within-cluster dialectal disparity is smaller for zero-shot transfer (i.e. ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\text{eng}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},C)}\leq|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}_{\bar{v}/\text{in-variety}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{v}},C)}|$). Here, the in-cluster/variety fine-tuning results in a higher performance deviation primarily due to the inconsistent variety-specific finetuning data quality.
5 Discussion
High resource vs. low resource varieties
The highest-performing varieties are mostly standard high-resource languages and a few high-resource dialects (Norwegian dialects) whereas, the majority of the lowest-performing language variants are low-resourced varieties. This clear distinction of language varieties points towards the large existence of in-cluster dialectal gaps. Furthermore, this finding correlates with language script differences. We observe that 77.2% of top-10 varieties in terms of maximum obtainable score are written with Latin script. Another finding is the performance instability of low-resource varieties across tasks. For instance, Old Guarani performs better in DEP parsing whereas, Mbyá Guarani (Paraguay) surpasses it in POS tagging even though the dataset remains the same (i.e. UD). For more detailed comparisons, in Appendix Table 21, we report the top-10 highest and lowest-scoring varieties across different tasks.
zero-shot few-shot / FT
Task mBERT XLM-R mBERT XLM-R
DEP. Parsing 61.6 61.3 76.2 64.3
POS Tagging 69.5 69.7 89.8 89.1
NER 59.7 57.8 65.8 61.4
NLI 56.9 62.5
TC 72.3 71.4 73.1 68.9
MRC 39.4 40.3
EQA 53.9 51.9 69.2 67.2
SA 78.8 80.1
DId 65.8 59.3
win 4/6 2/6 6/8 2/8
Table 5: Base model comparison. We found mBERT was easier to fine-tune using the default hyperparameter setting thus, resulting in a higher winning rate.
Model hyperparameter tuning
In Table 5, we compare the baseline multilingual models. mBERT was comparatively easier to train than XLM-R using the default learning rates reported in the earlier task experiments. Often for tasks such as MRC, we needed to tune hyperparameters (e.g. learning rate, max. sequence length) in case of XLM-R. However, once we identify a hyperparameter configuration that converges in a zero-shot setting, we use the same to train all the language-variant training for that specific task. As a result, for some low-resource languages, XLM-R does not converge compared to mBERT. For example, XLM-R dependency parsing UAS score for Norwegian-NynorskLIA is 56.08 (zeroshot) and 8.25 (in-variety FT) whereas, we get 78.39 for in-variety mBERT fine-tuning. We suspect this hyperparameter tuning issue is one of the contributing factors toward a lower winning rate of XLM-R in few-shot / fine-tuning settings. However, this could be improved further with an extensive parameter grid search and settings specifically tailored for each language cluster.
Positive zero-shot transfer for Latin varieties
Low-resource varieties written in Latin script receive greater benefit in zero-shot because of effective transfer from high-resource Standard English. With the presence of In-cluster/variety finetuning data, we effectively diminish this script effect to some extent. For example in the Hindi cluster, Fiji performs better than its Latin non-standard counterparts with in-cluster finetuning for NER (LABEL:tab:ner_full). In summary, if the standard variety of a cluster is non-latin but high-resource, then the success rate of in-cluster/variety fine-tuning tends to be higher. However, where all dialects are low-resource, Latin script varieties utilizing zero-shot transfer, eventually surpass others in the performance hierarchy. As an example, we report the zero-shot NER instances where the low-resource varieties perform better than the representative ones in Appendix Table 22 (most of these use Latin script).
LLM evaluation via In-context learning
For SA and EQA tasks, we have in-context learning results using the Mistral7B LLM. Comparing the performance against zero-shot and fine-tuning using our encoder-based models, we find dialect performance of LLM is better than zero-shot transfer but falls behind the finetuned results. On top of that, data contamination Ahuja et al. (2023) during LLM evaluation is another existing issue while considering these few available dialectal resources. Creating translation-based comparable data might be a solution to perform a fair benchmarking of LLMs on low-resource varieties.
Misinterpreting evaluation metrics
We also report the cluster-level population-weighted average (i.e. demographic utility) which rewards a system more when it provides increased linguistic utility (eg. raw F1 score) for varieties, spoken by a larger population compared to varieties spoken by a smaller population. Alone, this metric could be misleading if we consider the fact that the performance gap among all varieties from a particular cluster should be minimal. On the other hand, solely looking into the linguistic utility average does not give a clear picture either (e.g. often overshadows the larger performance deviation of certain varieties having extreme scores). So for all clusters and tasks, we report the linguistic utility average as well as the demographic utility average, the minimum score of a cluster, and the standard deviation in Tables 23 to 30.
6 Conclusion
We propose DialectBench, the first-ever inclusive Dialectal NLP benchmark reporting performance evaluation and disparity across standard and non-standard varieties. This is one step towards the effort of bringing more and more language under the paradigm of NLP technology. We would like to further improve the benchmark, constructing high-quality comparable data and expanding the task coverage to speech-based NLP technologies.
Limitations
The data quality and quantity, variety coverage vary significantly across tasks because of data scarcity issues. We avoid full-scale LLM-evaluation consciously because of the uncertain data-contamination issue and their well-known lower performance threshold compared to smaller masked-language-modeling-based fine-tuned models. In addition, we focus on text-based NLP tasks for this current iteration. Moreover, we do not claim the representative varieties of each language clusters to be any kind of superior or standardized forms over the other varieties. These varieties are chosen to perform a well-informed comparison among the perceived well-resourced linguistic variety and its counterparts having lesser data availability. At the same time, the mutual intelligibility and phylogenetic similarity of the similar cluster varieties also vary across cluster and this was not selected in a numerically quantifiable manner.
Evaluation Limitations
To further improve the evaluation fairness of the current version of DialectBench, we need (i) Parallel corpus utilization to prepare task-specific data (ii) Translation-based task data generation to perform comparable analysis (iii) Quantifying the resource-supply and demand as well as population-coverage Song et al. (2023) to identify where a variety stands in the global landscape of linguistic utility. Here, we have accumulated data for diverse varieties across tasks that vary significantly in terms of quality, example count, and domain. However, to perform a perfectly fair comparison of dialectal inequality, we should consider high-quality comparable data (e.g. parallel corpus, similar varieties across tasks) which is not available at this point.
Continuity of DialectBench
Despite our best effort, this benchmark does not include every one of the already published task-specific dialectal datasets. So, our next steps on this project involve hosting the benchmark on the website that displays the current statistics of the datasets in DialectBench. We will also encourage researchers to add new and existing datasets for tasks, language clusters that might be currently missing alongside the respective baselines.
Space Limitations
Our study encompasses a large set of evaluation result tables, their corresponding visualizations and findings analysis. Due to space limitations, we have to move the detailed reports (Appendix E) and the rest of the visualizations (Appendix D) in the Appendix. To better assist, we include an Appendix Table of Content (Table 6) at the introductory section of Appendix (Appendix).
Ethics Statement
This work is a compilation of existent dialectal datasets across different tasks, including structured prediction and generative tasks. Our experiments do not particularly optimize for the best model performance of these tasks. Therefore we acknowledge that for some of the tasks, the baseline models might not be robust enough to handle dialectal text hence resulting in wrong predictions and generations. We believe that this underscores the need for building models robust to different language variations and future work should focus on this.
Acknowledgements
This material is based upon work supported by the US National Science Foundation under Grants No. IIS-2125466, IIS-2125948, CAREER Grant No. IIS2142739, as well as NSF Grants No. IIS-2203097 and IIS-2125201. We gratefully acknowledge support from Alfred P. Sloan Foundation Fellowship. This research is also supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #2022-22072200004. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
Appendix
In this supplementary material, we provide the following: (i) Relevant literature review (ii) Overall results and dataset details that we could not fit into the main body of the paper.
Section Desctiption
Appendix A Related Work
Appendix B Tasks of DialectBench
    Translate-Test NLI Dataset Statistics (LABEL:{tab:nli_data})
Appendix C Varieties and Clusters of DialectBench
    DialectBench Variety list (LABEL:lang_list)
    Language Clusters and Representative Varieties (LABEL:stdv)
Appendix D Result Visualizations
Regional maps with aggregated Machine Translation scores (Figs. 5 to 6)
    Map of Switzerland with aggregated BLEU scores of Swiss-German variety per region (Fig. 5)
    Map of Italy with aggregated BLEU scores of Italian variety per region (Fig. 6)
Task Specific Plot for Maximum Scores (Figs. 7 to 10)
    Parts-of-Speech Tagging (Fig. 7(a))
    Named entity recognition (Fig. 7(b))
    Topic classification (Fig. 7(c))
    Natural language inference (Fig. 8(a))
    Extractive question answering (Fig. 8(b))
    Multiple-choice machine reading comprehension (Fig. 8(c))
    Sentiment analysis (Fig. 9(a))
    Dialect identification (Fig. 9(b))
    Machine translation (MT-region) (Fig. 10(a))
    Machine translation (MT-dialect) (Fig. 10(b))
Dialectal Gap visualization utilizing zero-shot cross-lingual transfer from Standard English.
    Parts-of-Speech Tagging (Fig. 11(a))
    Named entity recognition (Fig. 11(b))
    Dialect identification (Fig. 11(c))
Appendix E Task Specific Evaluation Result Tables
    Dependency parsing (Table 10)
    Parts-of-Speech tagging (Table 11)
    Named entity recognition (LABEL:tab:ner_full)
    Natural language inference (Table 13)
    Extractive question answering (Tables 14 to 15)
    Dialect identification (Table 16)
    Topic classification (Table 17)
    Sentiment analysis (Table 18)
    Multiple-choice machine reading comprehension (Table 19)
    Machine translation (LABEL:zero-shot-mt)
Appendix F Highest performing and lowest performing varieties (LABEL:{tab:high-lang})
Appendix G Low-resource Variety performing better in zero-shot NER (Table 22)
Appendix H Cluster level Result Summaries with Demographic Utility and Standard Deviation report (Tables 23 to 31)
Appendix I In-Context Learning Details
Table 6: Table of contents for supplementary material.
Appendix A Related Work
The majority of multilingual benchmarks Hu et al. (2020); Ruder et al. (2023, 2021); Liang et al. (2020); Wilie et al. (2020); Park et al. (2021) have heavily focused on dominant language varieties. However, the performance disparity between standard languages and their dialectal counterparts has been studied Kantharuban et al. (2023); Ziems et al. (2022); Rios (2020); Ziems et al. (2023) and shown to be significantly large across different tasks. Still, the large-scale generalization of this finding comprising numerous task categories is yet to be done. There have been previous efforts to bridge this gap though. Some work has focused on curating dialectal datasets across several tasks within one language cluster, while others have focused on single tasks across many language clusters. For instance, ORCA Elmadany et al. (2023), ARLUE Abdul-Mageed et al. (2021) and ALUE Seelawi et al. (2021) are dedicated natural language understanding benchmarks focusing on Arabic varieties alone. Multi-VALUE Ziems et al. (2023) was developed for benchmarking NLP tasks in English varieties. Mirzakhalov et al. (2021a) is a suite of resources for benchmarking MT in several Turkic languages. There has also been a large body of work on dialect identification across several languages Jauhiainen et al. (2022); Hämäläinen et al. (2021). Recently CODET Alam et al. (2023) was released as a contrastive dialectal MT benchmark covering 882 different variations from nine different languages. To the best of our knowledge, we are the first to do a large scale aggregation of dialectal data across several language clusters and tasks.
Appendix B Tasks of DialectBench
DialectBench includes 10 NLP tasks falling into four broader categories: structured prediction, sentence classification, question answering, and text generation. In Table 1, we present statistics for the datasets for each task and briefly discuss each task below.
Dependency parsing
For the dependency parsing task, we include only those Universal Dependencies (UD) Zeman et al. (2021) languages that have dialectal data. Beyond the data available in UD 2.12, we incorporate two additional datasets for African-American English (AAVE) Twitter data Blodgett et al. (2018) and Singlish Wang et al. (2017). To make these two datasets compatible with the UD processing pipeline, we replace the original dependency labels with the labels corresponding to the official UD formalism.
Part of speech (POS) tagging
We use the same UD languages for POS tagging that we used for dependency parsing. At the same time, we use the POS data instances from Singlish Wang et al. (2017). Moreover, we include six Finnish dialects, four Arabic dialects and Occitan through the UPOS label standardized pipeline proposed by Blaschke et al. (2023).
Named entity recognition (NER)
We use data from the 176-language version of the Wikiann dataset processed by Rahimi et al. (2019). All these languages provide both training and test data. In addition, we include dialectal data from the original Wikiann dataset (282 languages) Pan et al. (2017) for evaluation. Moreover, we include three Norwegian dialects Johansen (2019) with train, test and validation datasets that use a slightly different set of NER tags (GEO, ORG, OTH, PER) compared to the one we use in Wikiann (LOC, ORG, PER). We leave these levels as it is and do not convert to the Wikiann tags.
Dialect identification
We include dialect identification experiments on Arabic, Greek, Portuguese, English, Spanish, and Swiss German dialectal datasets. In these datasets, we find large variations in the level of granularity with which dialects are classified. For instance, the MADAR corpus differentiates Arabic varieties at the city level Bouamor et al. (2018), whereas our English and Spanish datasets are labeled with country names Zampieri et al. (2023).
Sentiment classification
Here we include several different Arabic varieties. Like other dialectal datasets, these datasets do not follow one standard labeling process. However, all datasets contain two main sentiment types: positive and negative. A number of datasets contain additional labels such as “objective” or “neutral.” In our setting, we perform a further split of data to provide validation data for each dialect. However, we do not remove these extra labeled data for information preservation.
Topic classification
We use the SIB-200 dataset Adelani et al. (2023) for topic classification task SIB-200 was constructed from the FLORES-200 translation datasets. The authors annotated the English dataset of FLORES-200 with 6 topic labels and then further propagated these labels to the translated instances for all other languages. For our case of benchmarking dialectal segments, we use the dialectal and regional varieties from SIB-200.
Variety # Sentences
cluster Language code
anglic eng_Latn english 5010
arabic acm_Arab north mesopotamian arabic 5010
acq_Arab ta’izzi-adeni arabic 5010
aeb_Arab tunisian arabic 5010
ajp_Arab south levantine arabic 5010
apc_Arab levantine arabic (a:north) 5010
arb_Arab standard arabic 5010
ars_Arab najdi arabic 5010
ary_Arab moroccan arabic 5010
arz_Arab egyptian arabic 5010
common turkic azb_Arab south azerbaijani 5010
azj_Latn north azerbaijani 5010
tur_Latn central oghuz (m:spoken) 5010
gallo-italian lij_Latn ligurian 5010
lmo_Latn lombard 5010
vec_Latn venetian 5010
gallo-rhaetian fur_Latn friulian 5010
high german lim_Latn limburgan 5010
ltz_Latn luxemburgish 5010
italian romance ita_Latn italian 5010
scn_Latn sicilian 5010
kurdish ckb_Arab central kurdish 5010
kmr_Latn northern kurdish 5010
latvian ltg_Latn east latvian 5010
lvs_Latn latvian 5010
modern dutch nld_Latn dutch 5010
norwegian nno_Latn norwegian nynorsk (m:written) 5010
nob_Latn norwegian bokmål (m:written) 5010
sardo-corsican srd_Latn sardinian 5010
sinitic yue_Hant cantonese 5010
zho_Hans classical-middle-modern sinitic (o:simplified) 5010
zho_Hant classical-middle-modern sinitic (o:traditional) 5010
sotho-tswana (s.30) nso_Latn northern sotho 5010
sot_Latn southern sotho 5010
southwestern shifted romance glg_Latn galician 5010
oci_Latn occitan 5010
por_Latn portuguese (a:european) 5010
spa_Latn spanish 5010
Table 7: Data statistics for newly created translate-test natural language inference (NLI) dataset. We prepare this translate-test NLI dataset by translating XNLI Conneau et al. (2018) english evaluation dataset.
Natural language inference
For the natural language inference task, there is no existing dataset with varieties. So we use the existing English test set of XNLI Conneau et al. (2018) and construct a multilingual dialect-focused translated evaluation dataset. We use a state-of-the-art machine translation model (NLLB-200 3B) to translate the English test set to 12 language clusters encompassing 40 varieties (Complete data statistics are reported in Table 7). After that, we perform zero-shot cross-lingual transfer from the English finetuned NLI model. We refer to this setting as translate-test.
Multiple-choice machine reading comprehension
This task aims to evaluate the capability of multiple-choice question answering given a context passage. The question could be answered by understanding the context passage while the right answer is given at one of the multiple choices. We use the Chinese, Sotho and Arabic clusters data from the recently released Belebele MRC dataset Bandarkar et al. (2023). This is an evaluation-only dataset.
Extractive question answering
This task predicts the answer span given a question and context passage. We use the SD-QA dialectal question-answering dataset Faisal et al. (2021). SD-QA is an evaluation dataset built on top of TyDiQA Clark et al. (2020), another well-known typologically diverse question-answering dataset. SD-QA contains the spoken utterances and transcription of the original TyDiQA question from speakers of English, Bengali, Arabic, Korean, and Swahili. We only use the textual part that contains the transcription of the dialectal spoken question matching the original TyDiQA question text. Note that since the transcriptions of the questions are obtained through automatic speech recognition, they may include both dialectal variations and noise due to ASR transcription errors.
Machine translation
We evaluate variety translation using the CODET benchmark Alam et al. (2023) and the TIL MT corpus Mirzakhalov et al. (2021b). CODET contains a contrastive dataset of 882 different varieties from nine different languages. We evaluate dialects here at city level for all languages except Italian and Swiss-German and aggregate dialects at the region level for them. The TIL corpus contains parallel translations across 22 Turkic languages, but in our evaluations we only include 8 turkic languages (Turkic, Sakha, Kazakh, Karakalpak, Bashkir, Azerbaijani, Kyrgyz) that have parallel English translations.
Appendix C Varieties and Clusters of DialectBench
C.1 DialectBench Variety list
Table 8: Varieties represented in DialectBench.
Language cluster Variety name Glottocode mBERT seen UDP POS NER SDQA RCMC NLI TC SC DI MT
albanian albanian alba1267 - - - - - - - -
gheg albanian gheg1238 - - - - - - - - -
anglic african american vernacular english afri1276 - - - - - - - - - -
australian english aust1314 - - - - - - - - - -
english stan1293 - - -
english (a:scotland) stan1293 - - - - - - - - - -
english (a:uk) stan1293 - - - - - - - - - -
english (o:controlled) stan1293 - - - - - - - - - -
indian english (a:north) indi1255 - - - - - - - - - -
indian english (a:south) indi1255 - - - - - - - - - -
irish english iris1255 - - - - - - - - - -
jamaican creole english jama1262 - - - - - - - - - -
kenyan english keny1281 - - - - - - - - - -
new zealand english newz1240 - - - - - - - - - -
nigerian english nige1260 - - - - - - - - - -
north american english nort3314 - - - - - - - - - -
old english (ca. 450-1100) olde1238 - - - - - - - - - -
philippine english phil1246 - - - - - - - - - -
singlish sing1272 - - - - - - - - -
southeast american english sout3300 - - - - - - - - - -
southern african english sout3331 - - - - - - - - - -
arabic aleppo alep1241 - - - - - - - - -
algerian arabic alge1239 - - - - - - -
arabian peninsula arabic (a:yemen) arab1393 - - - - - - - - -
arabic (a:bahrain) - - - - - - - - - - -
arabic (a:jordan) - - - - - - - - - -
arabic (a:saudi-arabia) - - - - - - - - - -
egyptian arabic egyp1253 - - - -
egyptian arabic (a:alx) egyp1253 - - - - - - - - -
egyptian arabic (a:asw) egyp1253 - - - - - - - - -
egyptian arabic (a:cai) egyp1253 - - - - - - - - -
egyptian arabic (a:kha) egyp1253 - - - - - - - - -
fez. meknes fezm1238 - - - - - - - - -
gilit mesopotamian arabic meso1252 - - - - - - - - -
gulf arabic gulf1241 - - - - - - - - - -
gulf arabic (a:doh) gulf1241 - - - - - - - - -
gulf arabic (a:jed) gulf1241 - - - - - - - - -
gulf arabic (a:mus) gulf1241 - - - - - - - - -
gulf arabic (a:riy) gulf1241 - - - - - - - - -
levantine arabic nort3139 - - - - - - - - - -
levantine arabic (a:north) nort3139 - - - - - - - -
levantine arabic (a:north-dam) nort3139 - - - - - - - - -
libyan arabic liby1240 - - - - - - - - - - -
libyan arabic (a:ben) liby1240 - - - - - - - - -
moroccan arabic moro1292 - - - - - -
najdi arabic najd1235 - - - - - - - -
north african arabic nort3191 - - - - - - - - -
north mesopotamian arabic nort3142 - - - - - - - -
north mesopotamian arabic (a:bas) nort3142 - - - - - - - - -
north mesopotamian arabic (a:mos) nort3142 - - - - - - - - -
rabat-casablanca arabic raba1252 - - - - - - - - -
sfax sfax1238 - - - - - - - - -
south levantine arabic sout3123 - - - - - -
south levantine arabic (a:south-amm) sout3123 - - - - - - - - -
south levantine arabic (a:south-jer) sout3123 - - - - - - - - -
south levantine arabic (a:south-sal) sout3123 - - - - - - - - -
standard arabic stan1318 - -
sunni beiruti arabic sunn1238 - - - - - - - - -
ta’izzi-adeni arabic taiz1242 - - - - - - - - -
tripolitanian arabic trip1239 - - - - - - - - -
tunisian arabic tuni1259 - - - - - - -
tunisian arabic (a:tun) tuni1259 - - - - - - - - -
tunisian arabic (r:casual) tuni1259 - - - - - - - - - -
basque basque (a:amenduze) basq1248 - - - - - - - - -
basque (a:azkaine) basq1248 - - - - - - - - - -
basque (a:baigorri) basq1248 - - - - - - - - - -
basque (a:barkoxe) basq1248 - - - - - - - - - -
basque (a:donibane) basq1248 - - - - - - - - - -
basque (a:garruze) basq1248 - - - - - - - - - -
basque (a:iholdi) basq1248 - - - - - - - - - -
basque (a:jatsu) basq1248 - - - - - - - - - -
basque (a:jutsi) basq1248 - - - - - - - - - -
basque (a:larzabale) basq1248 - - - - - - - - - -
basque (a:luhuso) basq1248 - - - - - - - - - -
basque (a:sara) basq1248 - - - - - - - - - -
basque (a:senpere) basq1248 - - - - - - - - - -
basque (a:suhuskune) basq1248 - - - - - - - - - -
basque (a:uharte) basq1248 - - - - - - - - - -
navarro-labourdin basque (a:behorlegi) basq1249 - - - - - - - - - -
navarro-labourdin basque (a:bidarrai) basq1249 - - - - - - - - - -
navarro-labourdin basque (a:helete) basq1249 - - - - - - - - - -
navarro-labourdin basque (a:mugerre) basq1249 - - - - - - - - - -
navarro-labourdin basque (a:urruna) basq1249 - - - - - - - - - -
souletin (a:maule) basq1250 - - - - - - - - - -
bengali vanga (a:barisal) vang1242 - - - - - - - - - -
vanga (a:dhaka) vang1242 - - - - - - - -
vanga (a:jessore) vang1242 - - - - - - - - - -
vanga (a:khulna) vang1242 - - - - - - - - - -
vanga (a:kushtia) vang1242 - - - - - - - - - -
vanga (a:west bengal) vang1242 - - - - - - - - -
circassian adyghe adyg1241 - - - - - - - - - -
kabardian kaba1278 - - - - - - - - - -
common turkic bashkir bash1264 - - - - - - - - -
central oghuz azer1255 - - - - - - - - - -
central oghuz (m:spoken) azer1255 - - - - - - - - -
crimean tatar crim1257 - - - - - - - - - -
kara-kalpak kara1467 - - - - - - - - - -
kazakh kaza1248 - - - - - - - - -
kirghiz kirg1245 - - - - - - - - -
north azerbaijani nort2697 - - - - - - -
sakha yaku1245 - - - - - - - - - -
south azerbaijani sout2697 - - - - - - -
turkish nucl1301 - - - - - - - -
uzbek uzbe1247 - - - - - - - - -
eastern romance aromanian arom1237 - - - - - - - - - -
moldavian mold1248 - - - - - - - - - -
romanian roma1327 - - - - - - - - -
eastern-western armenian eastern armenian nucl1235 - - - - - - - -
western armenian homs1234 - - - - - - - - -
farsic dari dari1249 - - - - - - - - - -
frisian ems-weser frisian sate1242 - - - - - - - - - -
northern frisian nort2626 - - - - - - - - - -
western frisian west2354 - - - - - - - - -
gallo-italian emiliano-romagnolo emil1243 - - - - - - - - - -
ligurian ligu1248 - - - - - -
lombard lomb1257 - - - - - - -
piemontese piem1238 - - - - - - - - -
venetian vene1258 - - - - - - - -
gallo-rhaetian anglo-norman angl1258 - - - - - - - - - -
arpitan fran1260 - - - - - - - - - -
french stan1290 - - - - - - -
french (a:paris) stan1290 - - - - - - - - -
friulian friu1240 - - - - - - - -
old french (842-ca. 1400) oldf1239 - - - - - - - - -
romansh roma1326 - - - - - - - - - -
walloon wall1255 - - - - - - - - - -
greater panjabic eastern panjabi panj1256 - - - - - - - - -
western panjabi west2386 - - - - - - - - -
greek apulian greek apul1237 - - - - - - - - - -
cretan cret1244 - - - - - - - - - -
cypriot greek cypr1249 - - - - - - - - - - -
cypriot greek (r:casual, m:written, i:fb) cypr1249 - - - - - - - - - -
cypriot greek (r:casual, m:written, i:other) cypr1249 - - - - - - - - - -
cypriot greek (r:casual, m:written, i:twitter) cypr1249 - - - - - - - - - -
modern greek mode1248 - - - - - - - - -
modern greek (r:casual, m:written, i:fb) mode1248 - - - - - - - - - -
modern greek (r:casual, m:written, i:other) mode1248 - - - - - - - - - -
modern greek (r:casual, m:written, i:twitter) mode1248 - - - - - - - - - -
pontic pont1253 - - - - - - - - - -
high german bavarian bava1246 - - - - - - - - -
central alemannic swis1247 - - - - - - - - - -
central alemannic (a:ag) swis1247 - - - - - - - - - -
central alemannic (a:ai) swis1247 - - - - - - - - - -
central alemannic (a:ar) swis1247 - - - - - - - - - -
central alemannic (a:be) swis1247 - - - - - - - - -
central alemannic (a:bl) swis1247 - - - - - - - - - -
central alemannic (a:bs) swis1247 - - - - - - - - -
central alemannic (a:fr) swis1247 - - - - - - - - - -
central alemannic (a:gl) swis1247 - - - - - - - - - -
central alemannic (a:gr) swis1247 - - - - - - - - - -
central alemannic (a:lu) swis1247 - - - - - - - - -
central alemannic (a:nw) swis1247 - - - - - - - - - -
central alemannic (a:ow) swis1247 - - - - - - - - - -
central alemannic (a:sg) swis1247 - - - - - - - - - -
central alemannic (a:sh) swis1247 - - - - - - - - - -
central alemannic (a:so) swis1247 - - - - - - - - - -
central alemannic (a:sz) swis1247 - - - - - - - - - -
central alemannic (a:tg) swis1247 - - - - - - - - - -
central alemannic (a:ur) swis1247 - - - - - - - - - -
central alemannic (a:vs) swis1247 - - - - - - - - - -
central alemannic (a:zg) swis1247 - - - - - - - - - -
central alemannic (a:zh) swis1247 - - - - - - -
central bavarian cent1967 - - - - - - - - - -
german stan1295 - - - - - - -
kölsch kols1241 - - - - - - - - - -
limburgan limb1263 - - - - - - - -
luxemburgish luxe1243 - - - - - - -
pennsylvania german penn1240 - - - - - - - - - -
pfaelzisch-lothringisch pala1330 - - - - - - - - - -
upper saxon uppe1465 - - - - - - - - - -
hindustani fiji hindi fiji1242 - - - - - - - - - -
hindi hind1269 - - - - - - - - -
inuit alaskan inupiaq inup1234 - - - - - - - - - -
kalaallisut kala1399 - - - - - - - - - -
italian romance continental southern italian neap1235 - - - - - - - -
italian ital1282 - - - - -
italian (a:abruzzo) ital1282 - - - - - - - - - -
italian (a:basilicata) ital1282 - - - - - - - - - -
italian (a:calabria) ital1282 - - - - - - - - - -
italian (a:campania) ital1282 - - - - - - - - - -
italian (a:emilia-romagna) ital1282 - - - - - - - - - -
italian (a:friuli-venezia giulia) ital1282 - - - - - - - - - -
italian (a:lazio) ital1282 - - - - - - - - - -
italian (a:liguria) ital1282 - - - - - - - - - -
italian (a:lombardia) ital1282 - - - - - - - - - -
italian (a:marche) ital1282 - - - - - - - - - -
italian (a:molise) ital1282 - - - - - - - - - -
italian (a:piemonte) ital1282 - - - - - - - - - -
italian (a:puglia) ital1282 - - - - - - - - - -
italian (a:sardegna) ital1282 - - - - - - - - - -
italian (a:sicilia) ital1282 - - - - - - - - - -
italian (a:toscana) ital1282 - - - - - - - - - -
italian (a:trentino-alto adige/südtirol) ital1282 - - - - - - - - - -
italian (a:umbria) ital1282 - - - - - - - - - -
italian (a:unknown) ital1282 - - - - - - - - - -
italian (a:veneto) ital1282 - - - - - - - - - -
italian (r:casual, m:written, i:tweet) ital1282 - - - - - - - -
italian (r:formal, m:written, i:essay) ital1282 - - - - - - - -
italian romance (a:apulia, m:spoken, i:tarantino) - - - - - - - - - - -
sicilian sici1248 - - - - - - -
komi komi komi1267 - - - - - - - - - -
komi-permyak komi1269 - - - - - - - -
komi-zyrian (m:spoken) komi1268 - - - - - - - - -
komi-zyrian (m:written) komi1268 - - - - - - - - -
korean korean (a:south-eastern, m:spoken) kore1280 - - - - - - - - - -
seoul (m:spoken) seou1239 - - - - - - - - -
kurdish central kurdish cent1972 - - - - - - - -
kurdish kurd1259 - - - - - - - - - -
northern kurdish nort2641 - - - - - - - - -
sine’i sine1239 - - - - - - - - - -
sorani sora1257 - - - - - - - - - -
latvian east latvian east2282 - - - - - - - -
latvian latv1249 - - - - - - -
mari eastern mari east2328 - - - - - - - - - -
western mari west2392 - - - - - - - - - -
modern dutch dutch dutc1256 - - - - - - -
western flemish vlaa1240 - - - - - - - - - -
zeeuws zeeu1238 - - - - - - - - - -
neva central and north pohjanmaa (a:ostrobothnian) cent1985 - - - - - - - - - -
estonian esto1258 - - - - - - - - -
finnish finn1318 - - - - - - - - -
häme (a:tavastian) hame1240 - - - - - - - - - -
neva (a:south-west trans) - - - - - - - - - - -
savo (a:savonian) savo1254 - - - - - - - - - -
southeastern finnish (a:south-east) sout1743 - - - - - - - - - -
southwestern finnish (a:south-west) sout2677 - - - - - - - - - -
norwegian norwegian norw1258 - - - - - - - - - -
norwegian (a:eastern) norw1258 - - - - - - - - - -
norwegian (a:setesdal) norw1258 - - - - - - - - - -
norwegian (a:southwestern) norw1258 - - - - - - - - - -
norwegian (m:written, i:samnorsk) - - - - - - - - - - -
norwegian bokmål (m:written) norw1259 - - - - -
norwegian nynorsk (m:written) norw1262 - - - - -
norwegian nynorsk (m:written, i:old) norw1262 - - - - - - - - -
saami north saami nort2671 - - - - - - - - -
skolt saami skol1241 - - - - - - - - -
sabellic umbrian umbr1253 - - - - - - - - -
sardo-corsican corsican cors1241 - - - - - - - - - -
sardinian sard1257 - - - - - - - -
serbian-croatian-bosnian bosnian standard bosn1245 - - - - - - - - -
croatian standard croa1245 - - - - - - - - -
serbian standard serb1264 - - - - - - - - -
serbian-croatian-bosnian sout1528 - - - - - - - - -
sinitic cantonese cant1236 - - - - - - - -
classical chinese lite1248 - - - - - - - -
classical-middle-modern sinitic (a:hongkong, o:traditional) clas1255 - - - - - - - - -
classical-middle-modern sinitic (o:simplified) clas1255 - - - - -
classical-middle-modern sinitic (o:traditional) clas1255 - - - - - - -
hakka chinese hakk1236 - - - - - - - - - -
mandarin chinese mand1415 - - - - - - - - - -
mandarin chinese (a:mainland, o:simplified) mand1415 - - - - - - - - - -
mandarin chinese (a:mainland, o:traditional, i:synthetic) mand1415 - - - - - - - - - -
mandarin chinese (a:taiwan, o:simplified) mand1415 - - - - - - - - - -
mandarin chinese (a:taiwan, o:traditional, i:synthetic) mand1415 - - - - - - - - - -
min nan chinese minn1241 - - - - - - - - - -
wu chinese wuch1236 - - - - - - - - - -
sorbian lower sorbian lowe1385 - - - - - - - - - -
upper sorbian uppe1395 - - - - - - - - - -
sotho-tswana (s.30) northern sotho nort3233 - - - - - - -
southern sotho sout2807 - - - - - - -
southwestern shifted brazilian portuguese braz1246 - - - - - - - -
romance extremaduran extr1243 - - - - - - - - - -
galician gali1258 - - - - - - -
latin american spanish amer1254 - - - - - - - - - -
mirandese mira1251 - - - - - - - - - -
occitan occi1239 - - - - -
portuguese (a:european) port1283 - - - - - -
portuguese (i:mix) port1283 - - - - - - - - -
portuguese (m:written) port1283 - - - - - - -
spanish stan1288 - - - - - -
spanish (a:europe) stan1288 - - - - - - - - - -
swahili swahili swah1253 - - - - - - - - - -
swahili (a:kenya) swah1253 - - - - - - - - - -
swahili (a:tanzania) swah1253 - - - - - - - - - -
tigrinya tigrinya (a:eritrea) tigr1271 - - - - - - - - - -
tigrinya (a:ethiopia) tigr1271 - - - - - - - - - -
tupi-guarani subgroup i.a mbyá guaraní (a:brazil) mbya1239 - - - - - - - - -
mbyá guaraní (a:paraguay) mbya1239 - - - - - - - - -
old guarani oldp1258 - - - - - - - - -
west low german west low german west2357 - - - - - - -
yoruba yoruba (a:central nigeria) yoru1245 - - - - - - - - -
Table 8: Varieties represented in DialectBench.
C.2 Language Clusters and Representative Varieties
Table 9: Language clusters and their standard varieties.
Task Cluster Name Cluster Representative
DEP. Parsing albanian albanian
arabic standard arabic
eastern-western armenian western armenian
sinitic classical-middle-modern sinitic (o:simplified)
anglic english
gallo-rhaetian french
high german german
tupi-guarani subgroup i.a old guarani
italian romance italian (r:formal, m:written, i:essay)
komi komi-zyrian (m:spoken)
norwegian norwegian bokmål (m:written)
southwestern shifted romance brazilian portuguese
saami north saami
POS Tagging albanian albanian
anglic english
arabic standard arabic
eastern-western armenian western armenian
gallo-rhaetian french
high german german
italian romance italian (r:formal, m:written, i:essay)
komi komi-zyrian (m:spoken)
neva finnish
norwegian norwegian bokmål (m:written)
saami north saami
sinitic classical-middle-modern sinitic (o:simplified)
southwestern shifted romance brazilian portuguese
tupi-guarani subgroup i.a mbyá guaraní (a:paraguay)
NER anglic english
arabic standard arabic
circassian adyghe
common turkic turkish
eastern romance romanian
frisian western frisian
gallo-italian piemontese
gallo-rhaetian french
greater panjabic western panjabi
greek modern greek
high german german
hindustani hindi
inuit alaskan inupiaq
italian romance italian
komi komi
kurdish kurdish
latvian latvian
mari eastern mari
modern dutch dutch
norwegian norwegian bokmål (m:written)
sardo-corsican sardinian
serbian-croatian-bosnian croatian standard
sinitic mandarin chinese
sorbian lower sorbian
sotho-tswana (s.30) southern sotho
southwestern shifted romance spanish
EQA anglic southeast american english
arabic arabic (a:saudi-arabia)
bengali vanga (a:dhaka)
korean seoul (m:spoken)
swahili swahili (a:kenya)
MRC anglic english
arabic standard arabic
sinitic classical-middle-modern sinitic (o:simplified)
sotho-tswana (s.30) northern sotho
TC anglic english
arabic standard arabic
common turkic north azerbaijani
gallo-italian venetian
high german luxemburgish
italian romance italian
kurdish northern kurdish
latvian latvian
norwegian norwegian bokmål (m:written)
sinitic classical-middle-modern sinitic (o:simplified)
sotho-tswana (s.30) northern sotho
southwestern shifted romance spanish
SC arabic standard arabic
MT-Dialect arabic gulf arabic (a:riy)
basque basque (a:azkaine)
bengali vanga (a:dhaka)
high german central bavarian
kurdish sorani
norwegian norwegian (a:eastern)
tigrinya tigrinya (a:ethiopia)
common turkic turkish
greek cretan
MT-Region italian romance italian (a:umbria)
high german central alemannic (a:zh)
DId anglic english (a:uk)
arabic standard arabic
greek modern greek (r:casual, m:written, i:twitter)
high german central alemannic (a:zh)
sinitic mandarin chinese (a:mainland, o:simplified)
southwestern shifted romance brazilian portuguese
NLI anglic english
arabic standard arabic
common turkic north azerbaijani
gallo-italian venetian
high german luxemburgish
italian romance italian
kurdish northern kurdish
latvian latvian
norwegian norwegian bokmål (m:written)
sinitic classical-middle-modern sinitic (o:simplified)
sotho-tswana (s.30) northern sotho
southwestern shifted romance spanish
Table 9: Language clusters and their clusters representatives.
Appendix D Result Visualizations
D.1 Regional maps with aggregated Machine Translation scores
Figure 5: Map of Switzerland with aggregated BLEU scores of Swiss-German variety per region
Figure 6: Map of Italy with aggregated BLEU scores of Italian variety per region
D.2 Task Specific Plot for Maximum Scores
(a) Parts of Speech Tagging
(b) Named Entity Recognition
(c) Topic Classification
Figure 7: Task specific plot of maximum obtainable score for all varieties. The yellow-shaded region represents language clusters having no varieties seen during mBERT pertaining. The bars with colored stripes represent the standard variety of a cluster
(a) Natural Language Inference
(b) Extractive Question Answering
(c) Multiple Choice Machine Reading Comprehension
Figure 8: Task specific plot of maximum obtainable Linguistic Utility for all varieties. The yellow-shaded region represents language clusters having no varieties seen during mBERT pertaining. The bars with colored stripes represent the standard variety of a cluster. The dialect with the Rawlsian score in each cluster is that with the leftmost bar.
(a) Sentiment Analysis
(b) Dialect Identification
Figure 9: Task specific plot of maximum obtainable Linguistic Utility for all varieties. The yellow-shaded region represents language clusters having no varieties seen during mBERT pertaining. The bars with colored stripes represent the standard variety of a cluster. The dialect with the Rawlsian score in each cluster is that with the leftmost bar.
(a) Machine Translation (Region level aggregation)
(b) Machine Translation (Variety level)
Figure 10: Task specific plot of maximum obtainable Linguistic Utility for all varieties. The yellow-shaded region represents language clusters having no varieties seen during mBERT pertaining. The bars with colored stripes represent the standard variety of a cluster. The dialect with the Rawlsian score in each cluster is that with the leftmost bar.
D.3 Dialectal Gap visualizations (zeroshot)
(a) Parts-of-speech (POS) tagging
(b) Named entity recognition (NER)
(c) Natural language inference (NLI)
Figure 11: Dialectal Gap visualization for language clusters utilizing zero-shot cross-lingual transfer from Standard English. The x-axis is for cluster gap while comparing against Standard English variety. Values far from zero have a larger performance gap from English. The y-axis is for aggregated cluster gap while comparing against standard cluster variety and values far from zero have larger within cluster gap. Ideally, we want both of them to be close to zero.
Appendix E Evaluation results
cluster variety UD-code Zeroshot (mBERT) Zeroshot (XLM-R) Finetune (mBERT) Finetune (XLM-R)
albanian albanian UD_Albanian-TSA 81.78 83.08 - -
gheg albanian UD_Gheg-GPS 38.14 43.50 - -
anglic english UD_English-EWT 91.55 91.10 91.55 91.1
singlish singlish 69.88 68.55 82.36 11.0
african american vernacular english TwitterAAE 50.53 52.53 - -
arabic standard arabic UD_Arabic-PADT 50.87 55.91 88.47 2.48
south levantine arabic UD_South_Levantine_Arabic-MADAR 49.94 45.75 - -
north african arabic UD_Maghrebi_Arabic_French-Arabizi 34.33 28.04 5.46 54.7
eastern-western armenian western armenian UD_Western_Armenian-ArmTDP 54.63 59.71 88.84 89.29
eastern armenian UD_Armenian-ArmTDP 53.27 62.63 85.55 87.01
gallo-italian ligurian UD_Ligurian-GLT 50.22 43.78 13.06 9.1
gallo-rhaetian french UD_French-ParTUT 80.87 82.17 93.93 92.09
french (a:paris) UD_French-ParisStories 57.64 61.36 77.54 77.54
old french (842-ca. 1400) UD_Old_French-SRCMF 56.12 46.52 91.51 89.78
high german german UD_German-LIT 72.65 76.50 - -
central alemannic (a:zh) UD_Swiss_German-UZH 36.77 34.70 - -
italian romance italian UD_Italian-PUD 78.99 78.58 - -
italian (r:formal, m:written, i:essay) UD_Italian-MarkIT 76.83 79.38 86.34 83.81
italian (r:casual, m:written, i:tweet) UD_Italian-PoSTWITA 61.88 62.13 85.82 86.32
continental southern italian UD_Neapolitan-RB 30.00 50.00 - -
komi komi-zyrian (m:spoken) UD_Komi_Zyrian-IKDP 26.89 32.14 - -
komi-permyak UD_Komi_Permyak-UH 26.12 30.91 - -
komi-zyrian (m:written) UD_Komi_Zyrian-Lattice 21.01 27.55 - -
norwegian norwegian bokmål (m:written) UD_Norwegian-Bokmaal 79.55 82.95 93.57 93.49
norwegian nynorsk (m:written) UD_Norwegian-Nynorsk 76.45 76.76 93.15 93.14
norwegian nynorsk (m:written, i:old) UD_Norwegian-NynorskLIA 56.58 56.08 78.39 8.25
saami north saami UD_North_Sami-Giella 23.58 16.87 67.56 5.96
skolt saami UD_Skolt_Sami-Giellagas 19.41 28.28 - -
sabellic umbrian UD_Umbrian-IKUVINA 33.21 28.75 - -
sinitic classical-middle-modern sinitic (a:hongkong, o:traditional) UD_Chinese-HK 58.97 61.78 - -
classical-middle-modern sinitic (o:simplified) UD_Chinese-GSDSimp 53.35 52.21 87.85 87.37
classical chinese UD_Classical_Chinese-Kyoto 46.72 35.14 19.65 18.11
southwestern shifted romance portuguese (a:european) UD_Portuguese-PUD 76.67 77.28 - -
portuguese (i:mix) UD_Portuguese-Bosque 75.89 77.96 92.64 92.15
brazilian portuguese UD_Portuguese-GSD 73.30 74.21 94.37 93.75
portuguese (m:written) UD_Portuguese-CINTIL 69.34 72.94 83.35 84.31
tupi-guarani subgroup i.a old guarani UD_Guarani-OldTuDeT 22.58 29.03 - -
mbyá guaraní (a:paraguay) UD_Mbya_Guarani-Thomas 13.51 11.15 - -
mbyá guaraní (a:brazil) UD_Mbya_Guarani-Dooley 8.95 4.23 - -
west low german west low german UD_Low_Saxon-LSDC 40.75 37.38 - -
Table 10: Dependency parsing evaluation report comprising zeroshot score and in-language finetuning. We report UAS as evaluation score. Zeroshot scores are evaluated using model finetuned on Standerd English. If training data is not available, we skip those languages (mentioned as ’-’) for in-language finetuning.
cluster variety Dataset-code dataset Zeroshot (mBERT) Zeroshot (XLM-R) Finetune (mBERT) Finetune (XLM-R)
albanian albanian UD_Albanian-TSA ud 75.80 84.41 - -
gheg albanian UD_Gheg-GPS ud 48.96 55.84 - -
anglic english UD_English-EWT ud 96.41 97.16 96.41 97.16
singlish singlish ud 76.27 77.55 87.38 87.96
arabic south levantine arabic UD_South_Levantine_Arabic-MADAR ud 51.99 61.84 - -
standard arabic UD_Arabic-PADT ud 39.74 56.67 95.72 96.11
gulf arabic dar-glf noisy 38.84 50.12 - -
egyptian arabic dar-egy noisy 36.14 51.39 - -
levantine arabic dar-lev noisy 32.66 43.76 - -
north african arabic UD_Maghrebi_Arabic_French-Arabizi ud 28.30 26.01 67.89 59.29
eastern-western armenian eastern armenian UD_Armenian-ArmTDP ud 71.78 82.63 91.31 92.36
western armenian UD_Western_Armenian-ArmTDP ud 70.27 75.31 94.86 95.14
gallo-italian ligurian UD_Ligurian-GLT ud 58.90 52.78 13.16 5.09
gallo-rhaetian french UD_French-ParTUT ud 84.36 85.47 96.88 96.42
french (a:paris) UD_French-ParisStories ud 81.37 82.77 96.70 96.76
old french (842-ca. 1400) UD_Old_French-SRCMF ud 64.70 59.41 95.64 95.64
high german german UD_German-LIT ud 87.08 88.36 - -
central alemannic (a:zh) UD_Swiss_German-UZH ud 62.56 47.18 - -
italian romance italian UD_Italian-PUD ud 81.09 83.12 - -
italian (r:formal, m:written, i:essay) UD_Italian-MarkIT ud 80.00 81.87 93.35 92.70
italian (r:casual, m:written, i:tweet) UD_Italian-PoSTWITA ud 73.71 76.45 95.80 96.83
continental southern italian UD_Neapolitan-RB ud 30.00 57.14 - -
komi komi-zyrian (m:spoken) UD_Komi_Zyrian-IKDP ud 41.25 46.66 - -
komi-permyak UD_Komi_Permyak-UH ud 29.52 43.67 - -
komi-zyrian (m:written) UD_Komi_Zyrian-Lattice ud 20.40 35.12 - -
neva finnish UD_Finnish-TDT ud 81.29 86.21 96.05 97.76
estonian UD_Estonian-EDT ud 80.34 85.17 96.49 97.20
hame (a:tavastian) murre-HAAM noisy 55.63 70.08 - -
central and north pohjanmaa (a:ostrobothnian) murre-POH noisy 55.09 69.68 - -
savo (a:savonian) murre-SAV noisy 54.70 69.42 - -
southeastern finnish (a:south-east) murre-KAA noisy 51.68 68.71 - -
southwestern finnish (a:south-west) murre-LVA noisy 49.80 68.54 - -
neva (a:south-west trans) murre-LOU noisy 42.57 61.72 - -
norwegian norwegian bokmål (m:written) UD_Norwegian-Bokmaal ud 88.53 89.55 98.19 98.67
norwegian nynorsk (m:written) UD_Norwegian-Nynorsk ud 85.06 85.81 97.83 98.32
norwegian nynorsk (m:written, i:old) UD_Norwegian-NynorskLIA ud 73.25 79.29 95.47 95.72
saami north saami UD_North_Sami-Giella ud 35.92 32.13 78.89 71.50
skolt saami UD_Skolt_Sami-Giellagas ud 20.26 34.15 - -
sabellic umbrian UD_Umbrian-IKUVINA ud 11.90 5.44 - -
sinitic classical-middle-modern sinitic (a:hongkong, o:traditional) UD_Chinese-HK ud 68.99 35.49 - -
classical-middle-modern sinitic (o:simplified) UD_Chinese-GSDSimp ud 58.26 30.92 94.72 95.02
classical chinese UD_Classical_Chinese-Kyoto ud 35.80 20.85 89.62 89.49
southwestern shifted romance portuguese (a:european) UD_Portuguese-PUD ud 80.08 81.38 - -
brazilian portuguese UD_Portuguese-GSD ud 78.63 80.12 98.19 98.50
portuguese (i:mix) UD_Portuguese-Bosque ud 78.48 79.85 97.71 97.81
occitan ROci noisy 76.84 65.80 - -
portuguese (m:written) UD_Portuguese-CINTIL ud 76.19 78.76 97.67 97.87
tupi-guarani subgroup i.a mbyá guaraní (a:paraguay) UD_Mbya_Guarani-Thomas ud 27.89 28.77 - -
old guarani UD_Guarani-OldTuDeT ud 8.96 10.30 - -
mbyá guaraní (a:brazil) UD_Mbya_Guarani-Dooley ud 1.94 0.59 - -
west low german west low german UD_Low_Saxon-LSDC ud 69.65 54.93 - -
Table 11: Parts-of-speech evaluation report comprising zeroshot score and in-language finetuning. We report F1 as evaluation score. Zeroshot scores are evaluated using model finetuned on Standerd English. If training data is not available, we skip those languages (mentioned as ‘-’) for in-language finetuning.
Table 12: Named entity recognition (NER) evaluation report comprising zeroshot score and in-group finetuning. We report F1 as evaluation score. Zeroshot scores are evaluated using model finetuned on Standerd English. If training data is not available, we skip those languages (mentioned as ‘-’) for in-language finetuning.
cluster variety target-code source dataset support Zeroshot (mBERT) Zeroshot (XLM-R) Finetune (mBERT) Finetune (XLM-R)
anglic english (o:controlled) simple en wikiann 1000 89.07 86.03 89.07 86.03
english en en wikiann 10000 84.15 82.11 84.15 82.11
old english (ca. 450-1100) ang en wikiann 100 54.41 55.94 54.41 55.94
jamaican creole english jam en wikiann 0 0.00 0.00 0.00 0.00
arabic egyptian arabic arz ar wikiann 100 43.82 50.58 67.77 73.33
standard arabic ar ar wikiann 10000 41.12 41.76 89.10 87.85
circassian adyghe ady en wikiann 693 67.33 54.03 - -
kabardian kbd en wikiann 1482 47.51 34.79 - -
common turkic turkish tr tr wikiann 10000 73.56 75.71 92.90 91.80
north azerbaijani az az wikiann 1000 67.31 61.01 89.17 88.26
crimean tatar crh tr wikiann 100 47.81 40.57 57.99 52.67
south azerbaijani azb az wikiann 2567 31.67 11.16 30.22 22.14
eastern romance romanian ro ro wikiann 10000 74.62 70.82 94.17 93.47
aromanian roa-rup el wikiann 732 64.78 62.66 68.92 63.33
moldavian mo ro wikiann 345 63.31 56.53 60.60 64.56
frisian western frisian fy nl wikiann 1000 80.17 71.96 81.06 78.21
ems-weser frisian stq nl wikiann 1085 59.45 57.47 64.55 55.15
northern frisian frr nl wikiann 100 46.67 46.22 54.78 52.63
gallo-italian piemontese pms it wikiann 100 79.53 71.10 87.88 78.72
lombard lmo it wikiann 100 72.49 68.17 79.37 78.66
venetian vec it wikiann 100 62.71 55.20 73.73 67.51
ligurian lij it wikiann 100 45.36 34.75 56.51 47.79
emiliano-romagnolo eml it wikiann 100 33.55 33.23 42.80 45.45
gallo-rhaetian french fr fr wikiann 10000 79.16 76.52 90.96 89.00
anglo-norman nrm fr wikiann 1281 66.78 88.56 71.98 71.92
arpitan frp it wikiann 2358 63.30 63.67 68.38 71.13
romansh rm it wikiann 100 56.88 55.19 69.69 67.58
friulian fur it wikiann 100 51.41 50.75 64.12 56.30
walloon wa fr wikiann 100 46.33 41.27 45.19 42.50
greater panjabic western panjabi pnb pa wikiann 100 64.46 53.78 17.82 0.00
eastern panjabi pa pa wikiann 100 32.90 45.25 30.63 0.00
greek modern greek el el wikiann 10000 71.76 72.96 91.18 90.68
pontic pnt el wikiann 291 66.37 69.79 68.45 72.58
high german german de de wikiann 10000 79.08 75.67 89.97 87.80
central alemannic als de wikiann 100 75.36 65.15 84.84 79.02
luxemburgish lb nl wikiann 1000 71.61 49.22 79.22 58.71
limburgan li nl wikiann 100 63.03 63.72 78.03 73.15
bavarian bar de wikiann 100 56.62 55.96 76.36 68.84
kölsch ksh nl wikiann 100 54.80 39.42 62.50 48.51
pfaelzisch-lothringisch pfl de wikiann 1092 49.40 47.19 59.12 50.14
pennsylvania german pdc de wikiann 100 41.76 41.79 39.39 38.76
hindustani fiji hindi hif hi wikiann 715 81.29 79.67 74.28 85.15
hindi hi hi wikiann 1000 65.74 65.77 88.11 85.83
inuit alaskan inupiaq ik en wikiann 431 76.27 70.56 - -
kalaallisut kl en wikiann 1403 63.20 60.20 - -
italian romance italian it it wikiann 10000 81.44 78.35 92.21 90.22
italian romance (a:apulia, m:spoken, i:tarantino) roa-tara it wikiann 3811 62.97 66.02 60.34 64.45
sicilian scn it wikiann 100 60.26 54.46 72.25 60.87
continental southern italian nap it wikiann 100 57.35 54.81 61.48 58.33
komi komi kv en wikiann 2464 56.78 41.78 - -
komi-permyak koi en wikiann 2798 53.62 47.80 - -
kurdish kurdish ku ku wikiann 100 31.67 59.48 13.70 0.00
central kurdish ckb ku wikiann 1000 6.90 35.49 1.17 0.00
latvian latvian lv lv wikiann 10000 69.36 71.07 93.24 92.30
east latvian ltg lv wikiann 1036 48.27 48.86 50.10 49.95
mari eastern mari mhr mhr wikiann 100 46.67 40.94 32.70 0.00
western mari mrj mhr wikiann 6036 38.29 58.46 5.51 0.00
modern dutch dutch nl nl wikiann 10000 82.03 80.42 91.96 90.51
western flemish vls nl wikiann 100 73.36 72.08 77.66 79.85
zeeuws zea nl wikiann 100 65.98 66.20 79.55 76.09
norwegian norwegian nynorsk (m:written) nynorsk nynorsk norwegian_ner 1511 87.58 87.86 87.58 87.86
norwegian (m:written, i:samnorsk) samnorsk samnorsk norwegian_ner 3450 86.96 90.55 86.96 90.55
norwegian bokmål (m:written) bokmaal bokmaal norwegian_ner 1939 85.82 90.54 85.82 90.54
sardo-corsican sardinian sc it wikiann 917 67.32 65.26 80.87 81.17
corsican co it wikiann 100 56.65 56.41 70.59 66.06
serbian-croatian-bosnian croatian standard hr hr wikiann 10000 77.59 78.12 92.40 91.02
bosnian standard bs hr wikiann 1000 69.93 74.88 87.50 88.86
serbian standard sr hr wikiann 10000 64.38 60.71 63.68 65.86
serbian-croatian-bosnian sh hr wikiann 10000 38.92 69.14 85.05 85.43
sinitic wu chinese wuu zh wikiann 100 71.89 35.80 74.15 63.73
min nan chinese zh-min-nan zh wikiann 100 44.68 47.30 21.40 15.08
cantonese zh-yue zh wikiann 10000 43.73 26.55 76.19 71.97
mandarin chinese zh zh wikiann 10000 42.86 24.71 81.12 77.22
classical chinese zh-classical zh wikiann 100 28.03 16.78 67.28 62.39
hakka chinese hak zh wikiann 100 27.43 31.84 36.55 40.00
sorbian lower sorbian dsb hsb wikiann 862 71.04 68.80 3.03 0.00
upper sorbian hsb hsb wikiann 100 65.44 65.48 38.20 0.00
sotho-tswana (s.30) southern sotho st en wikiann 339 64.36 69.26 - -
northern sotho nso en wikiann 720 19.08 29.66 - -
southwestern shifted romance galician gl es wikiann 10000 81.98 80.27 87.99 86.14
spanish es es wikiann 10000 72.80 70.73 92.17 90.32
occitan oc it wikiann 100 72.00 67.58 78.50 75.35
mirandese mwl es wikiann 100 46.20 44.07 49.29 42.81
extremaduran ext es wikiann 100 44.83 38.33 61.82 40.00
west low german west low german nds de wikiann 100 80.29 66.44 79.40 70.99
Table 12: Named entity recognition (NER) evaluation report comprising zeroshot score and in-group finetuning. We report F1 as evaluation score. Zeroshot scores are evaluated using model finetuned on Standerd English. If training data is not available, we skip those languages (mentioned as ‘-’) for in-language finetuning.
cluster variety target-code src mBERT (acc) XLM-R (acc) mBERT (F1) XLM-R (F1)
anglic english eng_Latn eng_Latn 81.90 83.35 81.95 83.43
arabic standard arabic arb_Arab eng_Latn 65.57 73.83 65.57 73.85
najdi arabic ars_Arab eng_Latn 59.42 69.02 59.14 68.94
ta’izzi-adeni arabic acq_Arab eng_Latn 58.84 68.72 58.64 68.62
moroccan arabic ary_Arab eng_Latn 54.65 58.30 54.61 58.14
egyptian arabic arz_Arab eng_Latn 54.53 65.87 53.86 65.70
south levantine arabic ajp_Arab eng_Latn 54.09 64.13 53.42 63.81
north mesopotamian arabic acm_Arab eng_Latn 52.95 58.94 52.84 58.75
levantine arabic (a:north) apc_Arab eng_Latn 52.14 61.74 51.40 61.31
tunisian arabic aeb_Arab eng_Latn 47.54 50.18 47.42 50.20
common turkic north azerbaijani azj_Latn eng_Latn 59.76 73.17 59.20 73.17
central oghuz (m:spoken) tur_Latn eng_Latn 59.14 74.47 58.37 74.52
south azerbaijani azb_Arab eng_Latn 46.05 41.82 44.58 39.24
gallo-italian venetian vec_Latn eng_Latn 65.15 68.52 64.99 68.55
lombard lmo_Latn eng_Latn 59.38 56.39 59.34 56.16
ligurian lij_Latn eng_Latn 57.82 57.70 56.70 57.16
high german luxemburgish ltz_Latn eng_Latn 60.34 47.33 60.01 46.21
limburgan lim_Latn eng_Latn 50.80 59.88 50.31 59.75
italian romance italian ita_Latn eng_Latn 73.71 78.08 73.71 78.19
sicilian scn_Latn eng_Latn 62.69 56.17 62.66 55.82
kurdish central kurdish ckb_Arab eng_Latn 40.98 44.93 37.40 39.59
northern kurdish kmr_Latn eng_Latn 39.10 63.45 33.93 63.26
latvian latvian lvs_Latn eng_Latn 60.14 73.57 59.95 73.63
east latvian ltg_Latn eng_Latn 48.62 54.19 47.02 53.54
norwegian norwegian bokmål (m:written) nob_Latn eng_Latn 72.44 79.44 72.45 79.51
norwegian nynorsk (m:written) nno_Latn eng_Latn 68.08 71.16 68.10 71.06
sinitic classical-middle-modern sinitic (o:simplified) zho_Hans eng_Latn 68.52 72.61 68.54 72.57
classical-middle-modern sinitic (o:traditional) zho_Hant eng_Latn 61.72 64.89 61.48 64.49
cantonese yue_Hant eng_Latn 60.44 68.02 60.27 67.41
sotho-tswana (s.30) northern sotho nso_Latn eng_Latn 39.38 40.18 35.06 35.98
southern sotho sot_Latn eng_Latn 39.36 39.30 34.62 34.16
southwestern shifted romance spanish spa_Latn eng_Latn 75.13 79.00 75.15 79.09
portuguese (a:european) por_Latn eng_Latn 73.73 79.18 73.73 79.22
galician glg_Latn eng_Latn 73.39 78.48 73.39 78.55
occitan oci_Latn eng_Latn 68.48 63.09 68.47 62.96
Table 13: Natural language inference (NLI) evaluation report using zeroshot cross-lingual transfer from Standard English. We report F1 as evaluation score. NLI uses F1 as evaluation matric. We prepare a translate-train dataset to perform this evaluation.
cluster variety target-code count Finetune (mBERT) Finetune (XLM-R) Zeroshot (mBERT) Zeroshot (XLM-R)
anglic irish english english–irl 494 73.00 67.98 68.62 62.45
southeast american english english–usa 494 73.51 67.95 68.56 62.97
new zealand english english–nzl 494 73.62 68.50 68.21 63.05
southern african english english–zaf 494 73.50 68.06 68.13 63.19
english (a:scotland) english–gbr 494 73.30 67.57 67.90 62.49
australian english english–aus 494 73.23 68.04 67.71 62.33
nigerian english english–nga 494 72.53 66.56 67.39 62.37
philippine english english–phl 494 73.75 67.29 67.35 61.54
indian english (a:south) english–ind_s 494 70.96 66.04 65.43 61.72
kenyan english english–kenya 494 70.63 65.64 65.30 60.28
indian english (a:north) english–ind_n 494 70.28 65.32 64.46 60.84
arabic moroccan arabic arabic–mar 324 70.94 65.14 50.94 49.56
tunisian arabic arabic–tun 324 71.36 65.31 50.86 50.01
arabic (a:jordan) arabic–jor 324 70.60 65.77 50.81 49.56
arabic (a:bahrain) arabic–bhr 324 70.96 65.86 49.88 50.36
arabic (a:saudi-arabia) arabic–sau 324 70.96 65.58 49.29 49.63
egyptian arabic arabic–egy 324 70.06 65.34 48.72 48.71
algerian arabic arabic–dza 324 68.63 64.71 48.71 48.65
bengali bengali (a:west bengal) bengali–ind 107 67.64 70.89 28.06 38.24
bengali (a:dhaka) bengali–dhaka 107 68.98 66.84 27.17 37.32
korean seoul (m:spoken) korean–korn 60 9.60 28.08 6.89 20.24
korean (a:south-eastern, m:spoken) korean–kors 60 9.27 26.96 6.89 19.66
swahili swahili (a:kenya) swahili–kenya 1000 72.06 71.90 46.20 46.22
swahili (a:tanzania) swahili–tanzania 1000 71.08 70.08 45.11 46.00
Table 14: Extractive dialectal question answering evaluation on SD-QA development set. We report span F1 as evaluation score. Zeroshot scores are evaluated using model finetuned on Standerd English whereas, we use combined training set for supervised finetuning
cluster variety target-code count Finetune (mBERT) Finetune (XLM-R) Zeroshot (mBERT) Zeroshot (XLM-R) ICL Mistral
anglic english (a:scotland) english–gbr 440 76.38 70.34 71.82 63.15 70.18
southern african english english–zaf 440 76.66 71.18 71.49 63.87 71.14
new zealand english english–nzl 440 76.71 71.39 71.22 63.69 70.95
australian english english–aus 440 75.66 70.89 71.20 62.28 69.23
southeast american english english–usa 440 77.26 71.50 71.17 63.71 71.76
irish english english–irl 440 75.52 70.73 70.92 62.15 70.64
philippine english english–phl 440 76.37 70.64 70.47 62.22 70.11
nigerian english english–nga 440 73.61 68.33 69.10 61.27 68.10
indian english (a:north) english–ind_n 440 74.62 68.03 68.84 61.25 68.99
kenyan english english–kenya 440 72.59 66.68 68.72 58.64 64.91
indian english (a:south) english–ind_s 440 71.93 66.88 66.49 60.36 65.13
arabic arabic (a:bahrain) arabic–bhr 921 77.52 72.11 53.25 53.28 55.33
arabic (a:jordan) arabic–jor 921 77.35 71.29 52.72 53.72 54.93
arabic (a:saudi-arabia) arabic–sau 921 77.88 72.11 52.72 53.24 55.66
algerian arabic arabic–dza 921 77.85 72.34 52.56 53.52 55.18
tunisian arabic arabic–tun 921 76.72 71.64 52.28 52.94 55.03
moroccan arabic arabic–mar 921 76.73 71.57 51.86 52.17 53.92
egyptian arabic arabic–egy 921 76.53 70.75 51.80 51.99 54.66
bengali bengali (a:west bengal) bengali–ind 113 68.62 73.27 32.30 36.39 50.06
bengali (a:dhaka) bengali–dhaka 113 67.37 74.24 31.79 35.52 51.73
korean seoul (m:spoken) korean–korn 276 10.15 31.91 7.26 19.62 65.79
korean (a:south-eastern, m:spoken) korean–kors 276 9.92 31.01 7.22 20.08 64.63
swahili swahili (a:tanzania) swahili–tanzania 472 63.54 62.30 38.24 39.38 55.70
swahili (a:kenya) swahili–kenya 472 72.25 70.53 37.97 41.59 58.71
Table 15: Extractive dialectal question answering evaluation on SD-QA test set. We report span F1 as evaluation score. Zeroshot scores are evaluated using model finetuned on Standerd English whereas, we use combined training set for supervised finetuning
cluster variety dialect-code support precision (mBERT) precision (XLM-R) recall (mBERT) recall (XLM-R) F1 (mBERT) F1 (XLM-R)
anglic english (a:uk) english:en-gb 249 98.10 89.57 83.13 71.59 90.00 79.58
north american english english:en-us 349 93.27 88.14 83.38 82.09 88.05 85.01
arabic aleppo arabic:ale 200 59.50 58.50 59.50 48.35 59.50 52.94
algerian arabic arabic:alg 272 79.00 65.50 58.09 62.68 66.95 64.06
arabian peninsula arabic (a:yemen) arabic:san 177 60.50 55.50 68.36 56.63 64.19 56.06
egyptian arabic (a:alx) arabic:alx 192 70.50 74.50 73.44 66.82 71.94 70.45
egyptian arabic (a:asw) arabic:asw 221 56.00 52.00 50.68 45.02 53.21 48.26
egyptian arabic (a:cai) arabic:cai 130 35.50 36.50 54.62 72.28 43.03 48.50
egyptian arabic (a:kha) arabic:kha 244 63.50 55.50 52.05 44.05 57.21 49.12
fez. meknes arabic:fes 196 60.00 59.50 61.22 56.40 60.61 57.91
gilit mesopotamian arabic arabic:bag 203 57.50 47.50 56.65 49.48 57.07 48.47
gulf arabic (a:doh) arabic:doh 205 50.00 43.50 48.78 45.55 49.38 44.50
gulf arabic (a:jed) arabic:jed 196 58.00 46.00 59.18 40.89 58.59 43.29
gulf arabic (a:mus) arabic:mus 178 38.50 49.50 43.26 42.67 40.74 45.83
gulf arabic (a:riy) arabic:riy 311 62.00 56.50 39.87 37.92 48.53 45.38
levantine arabic (a:north-dam) arabic:dam 148 37.50 24.50 50.68 42.98 43.10 31.21
libyan arabic (a:ben) arabic:ben 238 56.50 47.50 47.48 52.78 51.60 50.00
north mesopotamian arabic (a:bas) arabic:bas 186 49.50 39.00 53.23 49.68 51.30 43.70
north mesopotamian arabic (a:mos) arabic:mos 188 71.50 70.00 76.06 69.31 73.71 69.65
rabat-casablanca arabic arabic:rab 153 50.00 40.00 65.36 60.61 56.66 48.19
sfax arabic:sfx 215 62.50 64.50 58.14 48.13 60.24 55.13
south levantine arabic (a:south-amm) arabic:amm 177 40.50 35.00 45.76 35.53 42.97 35.26
south levantine arabic (a:south-jer) arabic:jer 202 48.50 47.00 48.02 40.34 48.26 43.42
south levantine arabic (a:south-sal) arabic:sal 167 46.00 66.50 55.09 59.11 50.14 62.59
standard arabic arabic:msa 244 75.00 98.00 61.48 95.61 67.57 96.79
sunni beiruti arabic arabic:bei 192 58.00 52.50 60.42 68.18 59.18 59.32
tripolitanian arabic arabic:tri 201 66.00 60.00 65.67 60.30 65.84 60.15
tunisian arabic (a:tun) arabic:tun 164 52.50 37.00 64.02 56.49 57.69 44.71
greek cypriot greek (r:casual, m:written, i:other) greek:cg_other 81 74.14 84.48 53.09 56.32 61.87 67.59
cypriot greek (r:casual, m:written, i:twitter) greek:cg_twitter 36 51.11 44.44 63.89 68.97 56.79 54.05
modern greek (r:casual, m:written, i:twitter) greek:smg_twitter 94 89.83 100.00 56.38 53.15 69.28 69.41
high german central alemannic (a:be) swiss-dialects:be 389 72.89 51.58 71.21 62.42 72.04 56.48
central alemannic (a:bs) swiss-dialects:bs 340 73.50 57.83 75.88 61.14 74.67 59.44
central alemannic (a:lu) swiss-dialects:lu 335 72.91 65.13 75.52 59.47 74.19 62.17
central alemannic (a:zh) swiss-dialects:zh 359 78.84 73.33 75.77 63.73 77.27 68.19
sinitic mandarin chinese (a:mainland, o:simplified) mandarin_simplified:m 986 97.90 96.10 99.29 90.66 98.59 93.30
mandarin chinese (a:mainland, o:traditional, i:synthetic) mandarin_traditional:m 977 96.80 92.10 99.08 95.74 97.93 93.88
mandarin chinese (a:taiwan, o:simplified) mandarin_simplified:t 1014 99.30 90.10 97.93 95.85 98.61 92.89
mandarin chinese (a:taiwan, o:traditional, i:synthetic) mandarin_traditional:t 1023 99.10 95.90 96.87 92.39 97.97 94.11
southwestern shifted romance brazilian portuguese portuguese:pt-br 627 96.94 92.35 90.91 84.98 93.83 88.51
latin american spanish spanish:es-ar 207 81.06 9.25 88.89 91.30 84.79 16.80
portuguese (a:european) portuguese:pt-pt 349 91.45 83.64 70.49 63.92 79.61 72.46
portuguese (m:written) portuguese:pt 15 9.70 0.00 86.67 0.00 17.45 0.00
spanish spanish:es 290 74.21 74.53 81.38 47.69 77.63 58.16
spanish (a:europe) spanish:es-es 492 90.99 83.33 82.11 78.89 86.32 81.05
Table 16: Dialect Identification evaluation using language cluster specific datasets. We finetune a classificaiton model using either mBERT or XLM-R and then evaluate on the test data.
cluster variety target-code src mBERT (acc) XLM-R (acc) mBERT (F1) XLM-R (F1)
arabic standard arabic arb_Arab arb_Arab 85.25 83.96 86.71 82.27
ta’izzi-adeni arabic acq_Arab arb_Arab 84.96 82.05 86.44 81.98
najdi arabic ars_Arab arb_Arab 84.80 84.39 87.41 83.33
north mesopotamian arabic acm_Arab arb_Arab 82.97 80.95 84.77 80.36
south levantine arabic ajp_Arab arb_Arab 81.82 80.16 84.16 79.05
levantine arabic (a:north) apc_Arab arb_Arab 81.59 80.15 83.76 79.88
egyptian arabic arz_Arab arb_Arab 81.02 76.38 84.43 81.03
tunisian arabic aeb_Arab arb_Arab 79.45 72.88 83.97 77.33
moroccan arabic ary_Arab arb_Arab 73.87 79.14 78.76 78.55
common turkic north azerbaijani azj_Latn azj_Latn 80.46 79.87 82.00 79.55
central oghuz (m:spoken) tur_Latn azj_Latn 79.10 84.41 80.61 79.51
south azerbaijani azb_Arab azj_Latn 65.90 67.08 69.71 68.37
gallo-italian venetian vec_Latn ita_Latn 76.72 70.68 75.07 74.28
lombard lmo_Latn ita_Latn 69.92 59.90 70.65 64.56
ligurian lij_Latn lij_Latn 66.81 63.42 74.03 57.78
high german luxemburgish ltz_Latn nld_Latn 74.74 58.50 77.86 64.83
limburgan lim_Latn nld_Latn 71.09 65.83 71.12 65.73
italian romance italian ita_Latn ita_Latn 87.67 84.92 86.68 85.83
sicilian scn_Latn ita_Latn 75.22 59.71 72.70 59.47
kurdish northern kurdish kmr_Latn ckb_Arab 33.23 68.21 10.45 5.71
central kurdish ckb_Arab ckb_Arab 13.10 19.37 16.86 12.38
latvian latvian lvs_Latn lvs_Latn 76.35 83.75 80.63 82.80
east latvian ltg_Latn lvs_Latn 55.67 65.02 63.69 67.42
norwegian norwegian nynorsk (m:written) nno_Latn nob_Latn 85.66 79.94 89.20 79.06
norwegian bokmål (m:written) nob_Latn nob_Latn 83.81 82.90 83.82 84.14
sinitic classical-middle-modern sinitic (o:traditional) zho_Hant zho_Hans 89.82 86.80 89.02 86.39
cantonese yue_Hant zho_Hans 89.45 86.46 88.71 87.64
classical-middle-modern sinitic (o:simplified) zho_Hans zho_Hans 88.74 86.38 88.86 89.15
sotho-tswana (s.30) northern sotho nso_Latn nso_Latn 35.62 28.16 34.86 13.55
southern sotho sot_Latn nso_Latn 32.55 32.31 39.93 19.08
southwestern shifted romance portuguese (a:european) por_Latn spa_Latn 88.13 89.10 88.10 87.74
galician glg_Latn spa_Latn 86.99 89.00 86.93 87.83
spanish spa_Latn spa_Latn 86.74 85.93 84.87 86.55
occitan oci_Latn lij_Latn 84.12 74.80 78.53 62.56
Table 17: Topic classification evaluation using SIB-200 language data with dialectal presence. We report span F1 as evaluation score. Zeroshot scores are evaluated using model finetuned on Standerd English whereas, we use in-group training for supervised finetuning
dialect MBERT_Acc MBERT_F1 XLMR_Acc XLMR_F1 mistral7b
lang-group variety
arabic tunisian arabic aeb_arab 94.55 94.56 94.61 94.62 73.3
algerian arabic arq_arab 84.98 85.00 84.70 84.69 76.0
arabic (a:jordan) jor_arab 82.96 82.90 89.07 89.00 82.2
arabic (a:saudi-arabia) sau_arab 81.38 65.97 83.40 67.66 79.8
tunisian arabic (r:casual) aeb_latn 80.95 65.65 79.80 63.70 62.3
standard arabic arb_arab 80.63 70.01 83.96 72.91 65.7
moroccan arabic ary_arab 78.08 61.50 77.41 55.55 58.4
egyptian arabic arz_arab 67.03 40.00 69.03 47.89 50.0
south levantine arabic ar_lb 58.38 34.63 58.90 32.72 0.0
Table 18: Sentiment Analysis results. In addition to, using mBERT and XLM-R as the base models, we also perform in-context learning to evaluate the performance of large language models (Mistral-7B).
dialect acc (mBERT) acc (XLM-R) mBERT (F1) XLM-R (F1)
lang-group language
anglic english eng_Latn 52.22 53.56 51.97 53.44
arabic standard arabic arb_Arab 39.00 43.78 39.01 43.78
levantine arabic (a:north) apc_Arab 38.89 40.78 38.64 40.71
north mesopotamian arabic acm_Arab 38.11 41.33 37.99 41.35
moroccan arabic ary_Arab 37.00 37.67 36.94 37.61
egyptian arabic arz_Arab 36.22 38.00 36.21 37.98
najdi arabic ars_Arab 36.00 38.11 36.05 38.16
sinitic classical-middle-modern sinitic (o:simplified) zho_Hans 50.11 47.22 49.79 47.10
classical-middle-modern sinitic (o:traditional) zho_Hant 47.00 45.11 46.88 44.76
sotho-tswana (s.30) northern sotho nso_Latn 31.11 29.78 31.18 29.72
southern sotho sot_Latn 28.56 29.11 28.52 29.00
Table 19: Multiple-choice machine reading comprehension evaluation using Belebele dataset languages with dialectal presence. We report span F1 as evaluation score. We use combined finetuning using the aggregated training data provided with Belebele evaluationd data.
Table 20: Zero-shot results for Machine Translation. We evaluate NLLB_600m and NLLB_1_3bn by translating each dialectal variety to English. For all languages without reference translations in English, we evaluate dialectal translations using the standard language’s translation as the reference. For varieties with parallel data like Yoruba, Turkish, Farsi, Tigrinya and Bengali we use the English reference provided in the dataset.
language_group variety NLLB_600m-bleu NLLB_1_3bn-bleu
arabic gulf arabic (a:riy) 43.07 43.07
gulf arabic (a:mus) 40.32 40.32
egyptian arabic (a:alx) 38.90 38.90
egyptian arabic (a:cai) 38.71 38.71
egyptian arabic (a:kha) 38.10 38.10
south levantine arabic (a:south-sal) 37.79 37.79
fez. meknes 37.19 37.19
north mesopotamian arabic (a:mos) 36.47 36.47
arabian peninsula arabic (a:yemen) 35.93 35.93
gilit mesopotamian arabic 35.78 35.78
egyptian arabic (a:asw) 35.02 35.02
south levantine arabic (a:south-amm) 34.88 34.88
north mesopotamian arabic (a:bas) 34.41 34.41
gulf arabic (a:doh) 34.20 34.20
gulf arabic (a:jed) 34.10 34.10
levantine arabic (a:north-dam) 33.36 33.36
south levantine arabic (a:south-jer) 33.27 33.27
libyan arabic (a:ben) 32.45 32.45
aleppo 31.64 31.64
tripolitanian arabic 31.21 31.21
rabat-casablanca arabic 31.17 31.17
sunni beiruti arabic 31.07 31.07
algerian arabic 27.25 27.25
tunisian arabic (a:tun) 25.10 25.10
sfax 23.24 23.24
basque basque (a:azkaine) 25.61 22.58
navarro-labourdin basque (a:helete) 22.53 22.03
basque (a:garruze) 25.40 22.00
navarro-labourdin basque (a:behorlegi) 24.19 21.35
basque (a:luhuso) 22.58 20.95
basque (a:amenduze) 24.21 20.86
basque (a:senpere) 23.47 20.35
basque (a:sara) 23.04 20.25
basque (a:jutsi) 21.46 19.99
navarro-labourdin basque (a:mugerre) 22.00 19.71
basque (a:baigorri) 20.21 18.80
basque (a:uharte) 21.10 18.57
basque (a:suhuskune) 21.10 18.57
basque (a:donibane) 20.28 18.03
basque (a:larzabale) 20.37 17.88
navarro-labourdin basque (a:bidarrai) 17.94 16.90
navarro-labourdin basque (a:urruna) 18.50 16.25
basque (a:iholdi) 16.68 15.01
basque (a:jatsu) 17.01 14.56
basque (a:barkoxe) 10.93 10.82
souletin (a:maule) 11.55 10.36
bengali vanga (a:jessore) 20.71 21.44
vanga (a:khulna) 18.96 19.73
vanga (a:kushtia) 17.36 19.12
vanga (a:dhaka) 17.18 17.85
vanga (a:barisal) 11.33 12.68
common turkic uzbek 24.68 28.82
turkish 22.24 24.22
bashkir 20.88 23.78
central oghuz 18.82 22.65
kirghiz 18.68 22.53
kazakh 19.56 20.96
kara-kalpak 14.31 18.23
sakha 2.53 2.43
farsic dari 37.02 40.49
greek cretan 25.03 21.34
apulian greek 3.92 3.97
high german central bavarian 9.49 12.01
upper saxon 11.99 8.63
kurdish sorani 35.74 34.88
sine’i 22.56 22.60
norwegian norwegian (a:eastern) 24.00 27.49
norwegian (a:southwestern) 21.88 15.81
norwegian (a:setesdal) 2.24 4.23
southwestern shifted romance occitan 20.73 25.74
tigrinya tigrinya (a:ethiopia) 22.40 25.36
tigrinya (a:eritrea) 18.64 21.02
yoruba yoruba (a:central nigeria) 21.46 18.10
high german central alemannic (a:zh) 43.71 44.06
central alemannic (a:gl) 43.90 43.96
central alemannic (a:tg) 43.94 43.62
central alemannic (a:bs) 42.94 43.48
central alemannic (a:gr) 43.67 43.31
central alemannic (a:ag) 43.62 43.26
central alemannic (a:sh) 43.62 43.01
central alemannic (a:zg) 43.05 42.95
central alemannic (a:ai) 42.27 42.72
central alemannic (a:ar) 43.90 42.67
central alemannic (a:sz) 43.24 42.57
central alemannic (a:vs) 41.69 42.52
central alemannic (a:ur) 44.08 42.41
central alemannic (a:be) 42.51 42.36
central alemannic (a:lu) 42.86 42.22
central alemannic (a:nw) 43.04 41.89
central alemannic (a:bl) 43.11 41.68
central alemannic (a:fr) 42.13 41.05
central alemannic (a:ow) 42.29 40.89
central alemannic (a:sg) 41.11 40.55
central alemannic (a:so) 40.27 39.39
italian romance italian (a:umbria) 40.73 41.62
italian (a:lazio) 39.03 32.23
italian (a:veneto) 29.38 31.27
italian (a:toscana) 28.90 31.24
italian (a:marche) 26.29 30.22
italian (a:trentino-alto adige/südtirol) 23.40 25.56
italian (a:unknown) 22.25 24.54
italian (a:friuli-venezia giulia) 18.62 23.51
italian (a:sicilia) 22.91 23.00
italian (a:calabria) 20.11 20.48
italian (a:liguria) 17.98 20.22
italian (a:lombardia) 16.23 17.32
italian (a:campania) 17.03 17.03
italian (a:molise) 17.73 15.98
italian (a:puglia) 16.50 15.64
italian (a:piemonte) 14.48 15.54
italian (a:basilicata) 13.76 14.36
italian (a:emilia-romagna) 13.86 14.35
italian (a:sardegna) 11.67 12.98
italian (a:abruzzo) 13.64 11.02
Table 20: Machine Translation
Appendix F Highest performing and lowest performing varieties
Varieties with Highest Performance Varieties with Lowest Performance
Task (Dataset)
DEP parsing (UD) anglic/english* italian romance/italian (r:formal, m:written, i:essay)* tupi-guarani subgroup i.a/old guarani* arabic/north african arabic
albanian/albanian* southwestern shifted romance/portuguese (a:european) komi/komi-zyrian (m:written)* italian romance/continental southern italian
gallo-rhaetian/french* norwegian/norwegian nynorsk (m:written)* saami/skolt saami* komi/komi-zyrian (m:spoken)$\dagger$
norwegian/norwegian bokmål (m:written)* southwestern shifted romance/portuguese (i:mix)* tupi-guarani subgroup i.a/mbyá guaraní (a:paraguay) komi/$\dagger$
italian romance/italian* southwestern shifted romance/brazilian portuguese* tupi-guarani subgroup i.a/mbyá guaraní (a:brazil)* saami/north saami*$\dagger$
EQA (SD-QA-test) anglic/english (a:scotland) anglic/irish english swahili/swahili (a:kenya)* arabic/algerian $\dagger$
anglic/southern african english anglic/philippine english bengali/vanga (a:west bengal)$\dagger$ arabic/tunisian $\dagger$
anglic/new zealand english anglic/nigerian english bengali/vanga (a:dhaka)$\dagger$ arabic/moroccan $\dagger$
anglic/australian english anglic/indian english (a:north) korean/seoul (m:spoken)$\dagger$ arabic/egyptian arabic*$\dagger$
anglic/southeast american english* anglic/kenyan english korean/korean (a:south-eastern, m:spoken)$\dagger$ swahili/swahili (a:tanzania)*
TC (SIB-200) sinitic/classical-middle-modern sinitic (o:traditional)$\dagger$ italian romance/italian* latvian/east latvian* arabic/moroccan $\dagger$
anglic/english* southwestern shifted romance/galician* sotho-tswana (s.30)/northern sotho* high german/limburgan
sinitic/cantonese*$\dagger$ southwestern shifted romance/spanish* kurdish/northern $\dagger$ gallo-italian/lombard
sinitic/classical-middle-modern sinitic (o:simplified)$\dagger$ norwegian/norwegian nynorsk (m:written)* sotho-tswana (s.30)/southern sotho* gallo-italian/ligurian
southwestern shifted romance/portuguese (a:european) arabic/standard arabic*$\dagger$ kurdish/central $\dagger$ common turkic/south $\dagger$
MRC (Belebele) anglic/english* arabic/north mesopotamian $\dagger$ arabic/moroccan $\dagger$ sinitic/classical-middle-modern sinitic (o:simplified)$\dagger$
sinitic/classical-middle-modern sinitic (o:simplified)$\dagger$ arabic/moroccan $\dagger$ arabic/egyptian arabic*$\dagger$ sinitic/classical-middle-modern sinitic (o:traditional)$\dagger$
sinitic/classical-middle-modern sinitic (o:traditional)$\dagger$ arabic/egyptian arabic*$\dagger$ arabic/najdi $\dagger$ arabic/standard arabic*$\dagger$
arabic/standard arabic*$\dagger$ arabic/najdi $\dagger$ sotho-tswana (s.30)/northern sotho* arabic/levantine arabic (a:north)$\dagger$
arabic/levantine arabic (a:north)$\dagger$ sotho-tswana (s.30)/northern sotho* sotho-tswana (s.30)/southern sotho* arabic/north mesopotamian $\dagger$
NER (Wikiann) anglic/english (o:controlled)* modern dutch/dutch* sinitic/classical $\dagger$ mari/western $\dagger$
norwegian/norwegian nynorsk (m:written)* southwestern shifted romance/galician* sinitic/hakka chinese*$\dagger$ gallo-italian/emiliano-romagnolo
norwegian/norwegian (m:written, i:samnorsk) italian romance/italian* sotho-tswana (s.30)/northern sotho* greater panjabic/eastern panjabi*$\dagger$
norwegian/norwegian bokmål (m:written)* hindustani/fiji $\dagger$ kurdish/central $\dagger$ common turkic/south $\dagger$
anglic/english* frisian/western frisian*$\dagger$ anglic/jamaican creole english kurdish/kurdish*$\dagger$
NLI (XNLI-translate-test) anglic/english* norwegian/norwegian bokmål (m:written)* common turkic/south $\dagger$ arabic/north mesopotamian $\dagger$
southwestern shifted romance/spanish* sinitic/classical-middle-modern sinitic (o:simplified)$\dagger$ kurdish/central $\dagger$ arabic/levantine arabic (a:north)$\dagger$
southwestern shifted romance/portuguese (a:european) southwestern shifted romance/occitan sotho-tswana (s.30)/northern sotho* high german/limburgan
italian romance/italian* norwegian/norwegian nynorsk (m:written)* sotho-tswana (s.30)/southern sotho* arabic/tunisian $\dagger$
southwestern shifted romance/galician* arabic/standard arabic*$\dagger$ kurdish/northern $\dagger$ latvian/east latvian*
POS tagging (UD) anglic/english* gallo-rhaetian/french (a:paris) tupi-guarani subgroup i.a/mbyá guaraní (a:paraguay) sinitic/classical $\dagger$
norwegian/norwegian bokmål (m:written)* neva/finnish* komi/komi-zyrian (m:written)* arabic/levantine $\dagger$
high german/german* italian romance/italian* saami/skolt saami* italian romance/continental southern italian
norwegian/norwegian nynorsk (m:written)* neva/estonian* tupi-guarani subgroup i.a/old guarani* komi/$\dagger$
gallo-rhaetian/french* southwestern shifted romance/portuguese (a:european) tupi-guarani subgroup i.a/mbyá guaraní (a:brazil)* arabic/north african arabic
writing system (non-MT) Latin (77.2%) Latin (44.2%)
dialect (non-MT) Standard (7%) Standard (4.5%)
Table 21: Varieties with the highest and lowest performance (in terms of raw evaluation score) on various tasks in the zero-shot setting. On the left are the top 10. We find that most of these varieties are high-resource standard varieties, and only a few high-resource non-standard varieties. At right are the bottom 10, which are mainly low-resource, non-standard varieties. We use the following notations for variety type and writing system quantification: * marks standard varieties (for some clusters, here we consider multiple varieties as standard because of the substantial resource presence of both of the varieties; e.g. portuguese/european and portuguese/mix) and $\dagger$ notes mix and non-Latin writing system.
Appendix G Low-resource variety performing better in zero-shot NER
Cluster High-resource  variety Script Low-resource variety Script
Hindustani Hindi Devanagari Fiji hindi Latin
Greater panjabic Eastern panjabi Devanagari Western panjabi Arabic
Sotho-tswana (s.30) Northern sotho Latin Southern sotho Latin
Table 22: NER cases where low-resource varieties perform better in zero-shot. The script plays a significant role here as most of the cases, high-resource scripts such as Latin, and Arabic could uplift the performance of a low-resource variety over a high-resource non-Latin script.
Appendix H Cluster level Result Summaries with Demographic Utility and Standard Deviation report
Table 23: Language clusters with their linguistic utility, demographic utility, variety with minimum linguistic utility and standard deviation. Task: DEP Parsing
clusters linguistic utility demographic utility variety (minimum ling. u.) standard deviation
albanian 63.3 69.1 (gheg albanian, 43.5) 28.0
anglic 75.5 90.7 (african american vernacular english, 52.5) 20.4
arabic 64.4 83.3 (south levantine arabic, 49.9) 21.0
eastern-western armenian 88.2 87.5 (eastern armenian, 87.0) 1.6
gallo-italian 50.2 50.2 (ligurian, 50.2) -
gallo-rhaetian 87.7 93.4 (french (a:paris), 77.5) 8.8
high german 56.6 76.5 (central alemannic (a:zh), 36.8) 28.1
italian romance 75.4 76.5 (continental southern italian, 50.0) 17.3
komi 30.2 30.1 (komi-zyrian (m:written), 27.6) 2.4
norwegian 88.4 - (norwegian nynorsk (m:written, i:old), 78.4) 8.6
saami 47.9 67.0 (skolt saami, 28.3) 27.8
sabellic 33.2 - (umbrian, 33.2) -
sinitic 65.4 - (classical chinese, 46.7) 20.8
southwestern shifted romance 87.2 91.6 (portuguese (a:european), 77.3) 7.9
tupi-guarani subgroup i.a 17.2 22.0 (mbyá guaraní (a:brazil), 9.0) 10.5
west low german 40.8 40.8 (west low german, 40.8) -
Table 24: Language clusters with their linguistic utility, demographic utility, variety with minimum linguistic utility and standard deviation. Task: POS Tagging
clusters linguistic utility demographic utility variety (minimum ling. u.) standard deviation
albanian 70.1 74.3 (gheg albanian, 55.8) 20.2
anglic 92.6 97.1 (singlish, 88.0) 6.5
arabic 61.9 81.4 (levantine arabic, 43.8) 18.9
eastern-western armenian 93.8 93.0 (eastern armenian, 92.4) 2.0
gallo-italian 58.9 58.9 (ligurian, 58.9) -
gallo-rhaetian 96.4 96.9 (old french (842-ca. 1400), 95.6) 0.7
high german 75.5 88.4 (central alemannic (a:zh), 62.6) 18.2
italian romance 82.6 80.9 (continental southern italian, 57.1) 17.9
komi 41.8 41.6 (komi-zyrian (m:written), 35.1) 6.0
neva 75.4 97.7 (neva (a:south-west trans), 61.7) 13.9
norwegian 97.6 - (norwegian nynorsk (m:written, i:old), 95.7) 1.6
saami 56.5 78.3 (skolt saami, 34.1) 31.6
sabellic 11.9 - (umbrian, 11.9) -
sinitic 84.5 - (classical-middle-modern sinitic (a:hongkong, o:traditional), 69.0) 13.7
southwestern shifted romance 90.5 95.7 (occitan, 76.8) 10.5
tupi-guarani subgroup i.a 13.7 18.5 (mbyá guaraní (a:brazil), 1.9) 13.7
west low german 69.7 69.7 (west low german, 69.7) -
Table 25: Language clusters with their linguistic utility, demographic utility, variety with minimum linguistic utility and standard deviation. Task: NER
clusters linguistic utility demographic utility variety (minimum ling. u.) standard deviation
anglic 57.3 83.9 (jamaican creole english, 0.0) 40.9
arabic 81.2 86.0 (egyptian arabic, 73.3) 11.1
circassian 57.4 54.1 (kabardian, 47.5) 14.0
common turkic 67.9 85.1 (south azerbaijani, 31.7) 28.8
eastern romance 75.9 79.3 (moldavian, 64.6) 16.0
frisian 66.8 80.4 (northern frisian, 54.8) 13.3
gallo-italian 68.6 68.0 (emiliano-romagnolo, 45.5) 17.3
gallo-rhaetian 71.8 90.3 (walloon, 46.3) 16.5
greater panjabic 54.9 53.2 (eastern panjabi, 45.2) 13.6
greek 81.9 90.1 (pontic, 72.6) 13.2
high german 71.5 88.0 (pennsylvania german, 41.8) 15.9
hindustani 86.6 88.1 (fiji hindi, 85.2) 2.1
inuit 69.7 63.5 (kalaallisut, 63.2) 9.2
italian romance 73.0 88.4 (continental southern italian, 61.5) 13.6
komi 55.2 56.6 (komi-permyak, 53.6) 2.2
kurdish 47.5 54.8 (central kurdish, 35.5) 17.0
latvian 71.7 88.7 (east latvian, 50.1) 30.5
mari 52.6 47.4 (eastern mari, 46.7) 8.3
modern dutch 83.8 91.2 (zeeuws, 79.5) 7.1
norwegian 89.7 - (norwegian nynorsk (m:written), 87.9) 1.5
sardo-corsican 75.9 79.8 (corsican, 70.6) 7.5
serbian-croatian-bosnian 83.1 81.0 (serbian standard, 65.9) 11.9
sinitic 64.3 78.3 (hakka chinese, 40.0) 16.8
sorbian 68.3 67.4 (upper sorbian, 65.5) 3.9
sotho-tswana (s.30) 49.5 41.1 (northern sotho, 29.7) 28.0
southwestern shifted romance 74.0 92.1 (mirandese, 49.3) 18.1
west low german 80.3 80.3 (west low german, 80.3) -
Table 26: Language clusters with their linguistic utility, demographic utility, variety with minimum linguistic utility and standard deviation. Task: NLI
clusters linguistic utility demographic utility variety (minimum ling. u.) standard deviation
anglic 83.4 83.4 (english, 83.4) -
arabic 63.3 70.0 (tunisian arabic, 50.2) 7.1
common turkic 64.1 62.3 (south azerbaijani, 44.6) 16.9
gallo-italian 61.7 63.6 (ligurian, 57.2) 6.0
gallo-rhaetian 54.6 54.6 (friulian, 54.6) -
high german 59.9 59.8 (limburgan, 59.7) 0.2
italian romance 70.4 77.1 (sicilian, 62.7) 11.0
kurdish 51.4 55.5 (central kurdish, 39.6) 16.7
latvian 63.6 71.5 (east latvian, 53.5) 14.2
modern dutch 76.5 76.5 (dutch, 76.5) -
norwegian 75.3 - (norwegian nynorsk (m:written), 71.1) 6.0
sardo-corsican 58.3 58.3 (sardinian, 58.3) -
sinitic 68.2 67.4 (classical-middle-modern sinitic (o:traditional), 64.5) 4.1
sotho-tswana (s.30) 35.3 35.6 (southern sotho, 34.6) 1.0
southwestern shifted romance 76.3 79.1 (occitan, 68.5) 5.2
Table 27: Language clusters with their linguistic utility, demographic utility, variety with minimum linguistic utility and standard deviation. Task: TC
clusters linguistic utility demographic utility variety (minimum ling. u.) standard deviation
anglic 89.7 89.7 (english, 89.7) -
arabic 84.5 85.7 (moroccan arabic, 79.1) 2.4
common turkic 78.7 77.3 (south azerbaijani, 69.7) 7.9
gallo-italian 73.8 73.7 (lombard, 70.6) 3.0
gallo-rhaetian 68.8 68.8 (friulian, 68.8) -
high german 74.5 72.7 (limburgan, 71.1) 4.8
italian romance 81.4 86.8 (sicilian, 75.2) 8.8
kurdish 43.8 52.3 (central kurdish, 19.4) 34.5
latvian 75.6 82.0 (east latvian, 67.4) 11.5
modern dutch 89.6 89.6 (dutch, 89.6) -
norwegian 86.7 - (norwegian bokmål (m:written), 84.1) 3.6
sardo-corsican 71.0 71.0 (sardinian, 71.0) -
sinitic 89.5 89.5 (classical-middle-modern sinitic (o:simplified), 89.2) 0.3
sotho-tswana (s.30) 37.8 36.9 (northern sotho, 35.6) 3.1
southwestern shifted romance 87.2 86.9 (occitan, 84.1) 2.3
Table 28: Language clusters with their linguistic utility, demographic utility, variety with minimum linguistic utility and standard deviation. Task: DId
clusters linguistic utility demographic utility variety (minimum ling. u.) standard deviation
anglic 89.0 88.4 (north american english, 88.0) 1.4
arabic 58.1 89.0 (south levantine arabic (a:south-amm), 43.0) 11.3
greek 64.6 - (cypriot greek (r:casual, m:written, i:twitter), 56.8) 6.8
high german 74.5 - (central alemannic (a:be), 72.0) 2.1
sinitic 98.3 98.3 (mandarin chinese (a:mainland, o:traditional, i:synthetic), 97.9) 0.4
southwestern shifted romance 73.3 82.7 (portuguese (m:written), 17.4) 27.9
Table 29: Language clusters with their linguistic utility, demographic utility, variety with minimum linguistic utility and standard deviation. Task: SA
clusters linguistic utility demographic utility variety (minimum ling. u.) standard deviation
arabic 80.3 81.4 (levantine/south, 58.9) 10.7
Table 30: Language clusters with their linguistic utility, demographic utility, variety with minimum linguistic utility and standard deviation. Task: MRC
clusters linguistic utility demographic utility variety (minimum ling. u.) standard deviation
anglic 53.4 53.4 (english, 53.4) -
arabic 39.9 42.1 (moroccan arabic, 37.6) 2.4
sinitic 48.3 - (classical-middle-modern sinitic (o:traditional), 46.9) 2.1
sotho-tswana (s.30) 30.1 30.5 (southern sotho, 29.0) 1.5
Table 31: Language clusters with their linguistic utility, demographic utility, variety with minimum linguistic utility and standard deviation. Task: EQA
clusters linguistic utility demographic utility variety (minimum ling. u.) standard deviation
anglic 75.2 75.4 (indian english (a:south), 71.9) 1.8
arabic 77.2 77.0 (egyptian arabic, 76.5) 0.6
bengali 73.8 - (vanga (a:west bengal), 73.3) 0.7
korean 65.2 64.6 (korean (a:south-eastern, m:spoken), 64.6) 0.8
swahili 67.9 64.1 (swahili (a:tanzania), 63.5) 6.2
Appendix I In-Context Learning Details
I.1 Prompts
We adapt the prompts from Super-NaturalInstructions Wang et al. (2022) for our in-context learning experiments.
Sentiment Analysis.
For sentiment analysis we provide 4 few-shot examples in the prompt. The prompt template is given below:
In this task, you are given a piece of text. Your task is to classify the sentiment of the text based on its content.
Sentence: <Sentence Example 1>
Label: <Label for Example 1, Positive, negative, neutral>
$\cdots$
Sentence: <Sentence Example $k$>
Label: <Label for Example $k$, Positive, negative, neutral>
Sentence: <Test Example Input>
Label:
Extractive Question Answering.
We provide 2 few-shot examples i.e. $k=2$ due to the long-form nature of text for this task.
This task is about writing a correct answer for the reading comprehension task. Based on the information provided in a given passage, you should identify the shortest continuous text span from the passage that serves as an answer to the given question. Avoid answers that are incorrect or provides incomplete justification for the question.
Passage: <Passage for Example 1>
Question: <Question for Example 1>
Answer: <Answer for Example 1>
$\cdots$
Passage: <Passage for Example $k$>
Question: <Question for Example $k$>
Answer: <Answer for Example $k$>
Passage: <Passage for Test Example>
Question: <Question for Test Example>
Answer: